### v2

# Regularity of Einstein $5$-manifolds via $4$-dimensional gap theorems

Abstract.
We refine the regularity of noncollapsed limits of $5$-dimensional manifolds with bounded Ricci curvature. In particular, for noncollapsed limits of Einstein $5$-manifolds, we prove that
  • tangent cones are unique of the form $\mathbb{R}\times\mathbb{R}^{4}/\Gamma$ on the top stratum, hence outside a countable set of points; this follows from a new isolation result for cones of the form $\mathbb{R}\times\mathbb{R}^{4}/\Gamma$ among all tangent cones,
  • the singular set is entirely contained in a countable union of Lipschitz curves and points,
  • away from a nowhere dense subset, these Lipschitz curves consist of smooth geodesics,
  • the interior of any geodesic is removable: limits of Einstein manifolds are real-analytic orbifolds with singularities along geodesic and bounded curvature away from their extreme points, and
  • if an asymptotically Ricci-flat $5$-manifold with Euclidean volume growth has one tangent cone at infinity that splits off a line, then it is the unique tangent cone at infinity.
These results prompt the question of the orbifold regularity of noncollapsed limits of Einstein manifolds off a codimension $5$ set in arbitrary dimension.
The proofs rely on a new result of independent interest: all spherical and hyperbolic $4$-orbifolds are isolated among Einstein $4$-orbifolds in the Gromov-Hausdorff sense. This yields various gap theorems for Einstein $4$-orbifolds, which do not extend to higher dimensions. The proofs of these gap theorems require a careful analysis of singular metrics and families of metrics that degenerate.
1. Introduction
An Einstein manifold $(M,g)$ satisfies the fundamental curvature condition
${\rm Ric}(g)=\Lambda g,\quad\Lambda\in\mathbb{R}.$ (1.1)
Such metrics are distinguished not only by having a curvature constant but also as stationary points of significant geometric functionals. Another curvature condition considered in this article is that of bounded Ricci curvature, which we normalize in dimension $n$ as
$(n-1)g\leqslant{\rm Ric}\leqslant(n-1)g.$
In dimensions greater than or equal to four, compact manifolds with uniform diameter bounds and noncollapsing conditions, sequences of Einstein metrics may develop singularities. Understanding these singularities and their formation has been a central question since the 1970’s. By the end of the 1980’s, noncollapsed limits of $4$-manifolds were roughly understood. More recently, the $4$-dimensional picture has been further clarified and a basic structure is emerging in higher-dimensions.
In the present article, we refine our understanding of limits in dimension $4$ and deduce several gap theorems for spherical and hyperbolic orbifolds, only valid in this dimension. This leads to substantial applications to the regularity of limits of Einstein $5$-manifolds, towards an orbifold regularity theorem away from a conjectually codimension $5$ singular set.
1.1. Background
1.1.1. Noncollapsed degeneration of Einstein $4$-manifolds
A central objective in four-dimensional differential geometry is to understand the structure and topology of the moduli space of Einstein metrics on a given compact manifold $M^{4}$, i.e. the set of Einstein metrics modulo the action of the diffeomorphism group, denoted $\mathbf{E}(M^{4})$. When equipped with the Gromov-Hausdorff metric $d_{GH}$, the completion of $\mathbf{E}(M^{4})$ includes Einstein orbifolds, as proven in [And89, BKN89, Tia90]. The need for a common background manifold $M^{4}$ was weakened to a noncollapsing assumption in [CN15]. The structure of this completion has recently been improved substantially in [Biq13, Biq16, Ozu22a, Ozu22b, Ozu20]. It has been shown in particular that the moduli space of Einstein metrics $GH$-close to a given Einstein $4$-orbifold $(M_{o},g_{o})$ was embedded in finite-dimensional set obtained by a gluing construction. Its dimension is controlled by the noncollapsedness.
It has long been an open question, explicitly posed in [And94, And10], whether Einstein orbifolds can be approximated or resolved by sequences of smooth Einstein metrics. Such approximations have been obtained by gluing-perturbation, see e.g. [Biq13, Don12, LS94]. On the other hand, in [Ozu21b], it was proven that spherical and hyperbolic orbifolds with the simplest singularities, $\mathbb{R}^{4}/\pm$, could never be limits of smooth Einstein $4$-manifolds in the Gromov-Hausdorff sense. In the present article, we show that all hyperbolic and spherical orbifolds are $GH$-isolated, even within the class of Einstein $4$-orbifolds. This is not a mere extension of [Ozu21b] and requires completely new ideas in order to deal with the deeply technical case of multiscale trees of singularities and singular metrics. The resulting isolation theorem implies a series of new surprising quantitative gap phenomena for Einstein $4$-orbifolds. These results are genuinely $4$-dimensional: they have no analogue in higher dimensions $n\geqslant 5$, even with isolated singularities. This is the first instance of gap theorems that deals with orbifold degenerations, while previous ones such as [HM25] relied on topological or geometric assumptions that prevented said degenerations.
1.1.2. Regularity of noncollapsed Einstein $n$-manifolds for $n\geqslant 5$
The higher-dimensional Einstein manifolds have been extensively studied over the last four decades. The modern compactness theory originates with the Cheeger-Gromov program of pointed $C^{\infty}$ convergence under curvature control. A foundational leap was the structure theory for spaces with lower Ricci curvature bounds [CC96, Col97, CC97, CC00a, CC00b], which introduced the stratification of the singular set and established the regularity of the limit space.
Uniqueness of tangent cones. One of the key problems in the regularity theory is the uniqueness of the tangent cones. For a noncompact manifold with nonnegative Ricci curvature and Euclidean volume growth, it was proved in [CC96] that the tangent cone at infinity is a metric cone. A priori, however, different rescaling sequences may lead to different tangent cones. A long-standing well-known question asks whether the cross-section of the tangent cone at infinity for noncollapsed Ricci-flat manifolds depends on the rescaling sequence of the blow-downs. It is known that uniqueness may fail if assuming merely the lower bound on the Ricci curvature [Per97, CC96, CoN13] and for collapsed Ricci-flat metrics [Hat17].
Such uniqueness for noncollapsed Ricci-flat manifolds was first established by Cheeger-Tian [CT94] under the assumption that the cross-sections of all tangent cones are smooth and integrable. Two decades later, Colding-Minicozzi [CM14] proved uniqueness assuming only that a single tangent cone is smooth, using the Łojasiewicz-Simon inequality. See also the related developments [KSz24, YZ25]. By contrast, when the cone has a nonsmooth cross-section – most notably for cones that split off a line $\mathbb{R}\times C(Z)$ – the theory remains far less developed. In the Kähler-Einstein setting, Donaldson-Sun [DS17] established uniqueness of tangent cones by means of algebro-geometric methods.
As our first result in higher dimensions, we prove that in dimension five, if one tangent cone splits off a line, then it is the unique tangent cone. Our argument does not rely on the Łojasiewicz-Simon inequality as in [CM14], nor on algebro-geometric techniques as in [DS17]. To the best of the authors’ knowledge, this is the first uniqueness result for a general Ricci-flat cone with a nonsmooth cross-section. As a consequence, we obtain a clearer description of the asymptotic geometry and new curvature decay estimates for Ricci-flat $5$-manifolds.
Analogous uniqueness questions arise in other geometric PDEs, including minimal surface theory [AA81, Sim83, Whi83, Sim94, Sze20b, CL23, FTW25], mean curvature flow [Sch14, CIM15, CM15, CS21, CM23, LZ24, BL25], and Ricci flow [SW15, CMZ24, LW24, CM25, FL25a], where uniqueness has been obtained under various smoothness, integrability, and structural assumptions on the cross-sections. We refer to the literature for further developments.
Regularity of limit spaces and singular sets. Through a series of works [And89, BKN89, Tia90, CN15], it was shown that limits of Einstein $4$-manifolds with bounded Ricci curvature and uniform lower volume bounds are Riemannian orbifolds with locally finitely many singularities. A natural question is whether such structural results extend to higher dimensions. Key recent milestones include the convexity of the regular set [CoN12], the resolution of the codimension-four conjecture [CN15] and the rectifiability of the singular set [CJN21, JN21]. Here, rectifiability means that, up to an $\mathcal{H}^{n-4}$-measure zero subset, $\mathcal{S}$ is contained in a countable union of bi-Lipschitz images of subsets of $\mathbb{R}^{n-4}$.
Several issues remained largely open, notably the treatment of the exceptional $\mathcal{H}^{n-4}$-measure zero subset and the possible emergence of an orbifold structure in higher dimensions. In dimension $5$ we prove:
  1. (1)
    Refined rectifiability and convexity. We prove the uniqueness of the tangent cone on the entire top stratum: if one tangent cone is $\mathds{R}\times\mathds{R}^{4}/\Gamma$, then it is the unique tangent cone. The uniqueness yields a refined stratification of the singular set $\mathcal{S}$ as a finite disjoint union of sets $\mathcal{S}^{0}$ and $\mathcal{S}^{1}_{\Gamma}$ for finitely many $\Gamma\subset O(4)$, where $\mathcal{S}^{0}$ consists of countably many points whose tangent cones do not split, and $\mathcal{S}^{1}_{\Gamma}$ consists of points whose tangent cone is $\mathbb{R}\times\mathbb{R}^{4}/\Gamma$. We show that each $\mathcal{S}^{1}_{\Gamma}$ is locally contained in a bi-Lipschitz image of a subset of $\mathbb{R}$, thereby establishing a bi-Lipschitz manifold structure for the singular set without discarding any $\mathcal{H}^{1}$-measure zero subset. Moreover, we prove that each connected component of $\mathcal{S}^{1}_{\Gamma}$ is either a single point or a geodesic, and hence locally convex.
  2. (2)
    Orbifold regularity. We extend the orbifold regularity to dimension $5$ and formulate a codimension-five orbifold conjecture for Einstein limits $X$. In particular, the singular curves in $\mathcal{S}^{1}_{\Gamma}$ are geodesic and possess a real-analytic orbifold structures with bounded curvature. Together with the geodesics in the regular set, these exhaust all geodesics in the limit space. This leads us to introduce the orbifold-regular-singular decomposition $X=\mathcal{R}_{\mathrm{orb}}\cup\mathcal{S}_{\mathrm{orb}}$. We prove that the orbifold-regular set $\mathcal{R}_{\mathrm{orb}}$ is an open convex 5-dimensional orbifold, and the orbifold-singular set $\mathcal{S}_{\mathrm{orb}}$ consists precisely of the extreme points of $X$, namely those points that do not lie on any geodesic joining two other points. We further show that $\mathcal{S}_{\mathrm{orb}}$ is contained in a nowhere dense subset of a countable collection of bi-Lipschitz curves. Since Einstein metrics are real-analytic, we further conjecture – by unique continuation – that $\mathcal{S}_{\mathrm{orb}}$ is countable. This will be discussed in detail in Section 1.3.
Many of the results above require only bounded Ricci curvature, and less central statements hold in all dimensions $n\geqslant 5$. See [Col96a, Col96b, CCT02, CT06, Col12, CN13, Bam17, Bam18, Bam20] for more developments in this direction.
1.2. Main results
1.2.1. Gap theorems for Einstein $4$-manifolds
We first show the isolation of all hyperbolic or spherical $4$-orbifolds, i.e. spaces with isolated flat conical singularities and sectional curvatures constant equal to $\pm 1$.
Theorem 1.1 (Isolation theorem).
For any compact hyperbolic or spherical $4$-orbifold $(M_{o},g_{o})$ with isolated singularities, there exists $\delta_{0}>0$ such that if an Einstein orbifold $(M,g)$ with ${\rm Ric}(g)=\pm 3g$ satisfies
$d_{GH}\left((M,g),(M_{o},g_{o})\right)<\delta_{0},$
then $(M,g)$ is isometric to $(M_{o},g_{o})$.
This result extends that of [Ozu21b] that was only dealing with smooth metrics close to orbifolds with simplest singularities $\mathbb{R}^{4}/\mathbb{Z}_{2}$. The present article deals with much more challenging general case which allows formation of trees of singularities which substantially complicate the problem and require new ideas.
Remark 1.2.
There are Ricci-flat $4$-manifolds approaching flat orbifolds [LS94] on $K3$, but not on arbitrary topologies [BK17, Ozu21a].
Theorem 1.1 yields gap theorems for Einstein $4$-orbifolds with no extension to higher dimensions.
Corollary 1.3 (Gap theorems).
For all $v>0$, there exists $\varepsilon_{0}=\varepsilon_{0}(v)>0$ such that if an Einstein $4$-orbifold $(M,g)$ with ${\rm Ric}(g)=3g$ and ${\rm Vol}(M,g)>v$ satisfies one of the following:
  1. (1)
    (Almost Bonnet-Myers theorem) ${\rm diam}(M,g)>\pi-\varepsilon_{0}$,
  2. (2)
    (Almost Obata-Lichnerowicz theorem) the first eigenvalue $\lambda_{1}$ of the Laplacian satisfies $\lambda_{1}<4+\varepsilon_{0}$,
  3. (3)
    (Almost conformally flat) $\|W_{g}\|_{L^{1}}<\varepsilon_{0}$,
  4. (4)
    (Almost Lévy-Gromov theorem) there exists some $t\in(17^{-17},1-17^{-17})$ with $\big|I_{(M,g)}(t)-I_{\mathbb{S}^{4}}(t)\big|\leqslant\varepsilon_{0}$, where $I$ denotes the isoperimetric profile $I_{(M,g)}(t):=\inf\big\{\mathrm{Per}(\Omega):\Omega\subset M,{\rm Vol}_{g}(\Omega)=t\,{\rm Vol}_{g}(M)\big\},$
  5. (5)
    (Almost Bishop-Gromov theorem) $(M,g)$ has a singularity modeled on $\mathbb{R}^{4}/\Gamma$ and ${\rm Vol}(M)>\frac{|\mathbb{S}^{4}|}{|\Gamma|}-\varepsilon_{0}$,
then $(M,g)$ is a spherical $4$-orbifold. In all cases but (3), $(M,g)$ is isometric to a spherical suspension over $\mathbb{S}^{3}/\Gamma$ for some finite $\Gamma\subset O(4)$.
Remark 1.4.
From [CTZ12], spherical 4-orbifolds are spherical suspensions over $\mathbb{S}^{3}/\Gamma$ and their $\mathbb{Z}_{2}$-quotient.
Remark 1.5.
Each gap theorem above strengthens well-known rigidity theorems for spherical suspensions. For (1) use the maximal-diameter theorem of Cheng [Che75], for (2) use the Lichnerowicz-Obata theorem [Lic58, Oba62], for (4) use the Lévy-Gromov estimate [Gro81] (see a survey [Ant25]), for (5) use Bishop-Gromov based at the singular point [BC64, Gro81]. In the RCD setting, equality yields a spherical suspension; see for instance [Ket15, CaM17]. Note also that the $L^{1}$ smallness in (3) is implied by the $L^{p}$ smallness with any $p\geqslant 1$ and that gap theorems for smooth Einstein $n$-manifolds with small $L^{\frac{n}{2}}$ norm of their Weyl tensors are proven in [Sin92, She90, IS02].
Remark 1.6.
Direct generalizations of Corollary 1.3 fail in dimension $5$ and higher: in [Böh98], Böhm constructed a sequence of Einstein $5$-metrics $(S^{5},g_{i})$ with ${\rm Ric}(g_{i})=4g_{i}$ and volume uniformly bounded below such that $(S^{5},g_{i})_{i\in\mathbb{N}}$ converges in the Gromov-Hausdorff sense to a spherical suspension where the diameter and the first eigenvalue of the $(S^{5},g_{i})$ are arbitrarily close to $\pi$ and $5$ respectively, and supposedly, since the singularities forming along the sequence are conical, we should have: for $1\leqslant p<\frac{5}{2}$, $\|W\|_{L^{p}}\to 0$. See Question 1.31 for a reasonable higher-dimensional generalization.
Remark 1.7.
Imposing loose bounds on $\rm Ric$} is not sufficient to obtain (almost) rigidity: by [Ozu22b, Remark 51], there exist smooth metrics $(\mathbb{S}^{2}\times\mathbb{S}^{2},g_{t})_{t\in(0,1)}$ with $3g_{t}\leqslant{\rm Ric}(g_{t})\leqslant 6.1g_{t}$ converging to a spherical orbifold. The metrics $g_{t}$ satisfy ${\rm diam}(\mathbb{S}^{2}\times\mathbb{S}^{2},g_{t})\to\pi$, one can show that the first eigenvalue tends to $4$, and for any $p\in[1,+\infty)$ the $L^{p}$-norm of the Weyl curvature converges to zero as $t\to 0$.
Remark 1.8.
The constants in our gap theorems of Corollary 1.3 are not explicit since they are proven by contradiction. Gap theorems with explicit constant, such as [Gur00] are exceptionally rare; see [AES19].
1.2.2. Tangent cones of Einstein $5$-manifolds
Our main new tool to study the regularity of noncollapsed limits of Einstein $5$-manifolds is the following result.
Theorem 1.9 (Isolation of $1$-symmetric Ricci-flat cones).
For all $v>0$, there exists $\varepsilon_{0}=\varepsilon_{0}(v)>0$ such that the following holds.
Let $(C(Y),d_{C(Y)},y)$ and $(C(Z),d_{C(Z)},z)$ be two metric cones that are limits of sequences of pointed 5-manifolds $(M_{i},g_{i},p_{i})$ and $(N_{i},\tilde{g}_{i},q_{i})$ both satisfying $|{\rm Ric}_{i}|\to 0$ and that the balls centered at $p_{i},q_{i}$ of radius $1$ have volume bounded below by $v>0$. Suppose that $(C(Z),d_{C(Z)},z)$ isometrically splits off a line and
$d_{GH}(Y,Z)<\varepsilon_{0}.$
Then the cones are isometric: $(C(Y),d_{C(Y)},y)\cong(C(Z),d_{C(Z)},z)\cong(\mathds{R}\times C(\mathbb{S}^{3}/\Gamma),g_{0},0)$ for $\Gamma\subset O(4)$ finite.
Remark 1.10.
Once again, the higher-dimensional analogous statement cannot hold by considering cones over the examples of [Böh98].
In particular, if these two cones arise as tangent cones at some point, this quantitative isolation directly implies the uniqueness of tangent cones splitting off a line.
Theorem 1.11 (Uniqueness of tangent cones).
Let $(M_{i},g_{i},p_{i})_{i\in\mathbb{N}}$ be a sequence of pointed $5$-manifolds with $|{\rm Ric}|<4$ and ${\rm Vol}(B_{1}(p_{i}))>v>0$ $d_{GH}$-converging to $(X,d,p)$. Assume that $(\mathbb{R}\times C(Y),d_{\mathbb{R}\times C(Y)})$ is a tangent cone at $x\in X$, then it is the unique tangent cone at $x$. In particular, except for countably many points, the tangent cone is uniquely isometric to $\mathds{R}\times C(\mathbb{S}^{3}/\Gamma)$ where $\Gamma\subset O(4)$ acts freely.
Moreover, we obtain quantitative regularity on a wedge region $\mathcal{W}(x)$ around any point $x\in X$ whose tangent cone splits off a line. The wedge region, defined in (4.3), consists of a neighborhood of $x$ with a scaling-invariant tubular neighborhood of the splitting axis removed.
Corollary 1.12.
If one tangent cone at $x\in X$ splits, then there exists some wedge region $\mathcal{W}(x)$ around $x$ and constant $\rho>0$ such that the harmonic radius (see for instance [And90] for a definition), $r_{h}(y)$, at the point $y\in\mathcal{W}(x)$ has a lower bound $r_{h}(y)\geq\rho\cdot d(x,y)$. If we further assume that the sequence $M_{i}$ is Einstein, then we have the curvature estimate $|{\rm Rm}|(y)=o\big(d(x,y)^{-2}\big)$ in $\mathcal{W}(x)$.
We next describe the asymptotic geometry of complete, noncollapsed, asymptotically Ricci-flat $5$-manifolds with Euclidean volume growth. When a tangent cone at infinity splits off a line, then it is unique.
Theorem 1.13.
Let $(M,g,p)$ be a 5-manifold with Euclidean volume growth and satisfy $\sup_{B_{2R}(p)\setminus B_{R}(p)}|{\rm Ric}|=o(R^{-2})$. If one of the tangent cones at infinity splits, then this is the unique tangent cone at infinity.
Finally, we obtain curvature control near infinity.
Corollary 1.14.
Let $(M,g,p)$ be a 5-manifold with Euclidean volume growth and satisfy $\sup_{B_{2R}(p)\setminus B_{R}(p)}|{\rm Ric}|=o(R^{-2})$. If one of the tangent cones at infinity splits, then there exists some wedge region around infinity $\mathcal{W}(\infty)$ and constant $\rho>0$ such that the harmonic radius at the point $y\in\mathcal{W}(\infty)$ has a lower bound $r_{h}(y)\geq\rho\cdot d(p,y)$. If we further assume that the manifold $M$ is Ricci-flat, then we have the curvature estimate $|{\rm Rm}|(y)=o\big(d(p,y)^{-2}\big)$ in $\mathcal{W}(\infty)$.
1.2.3. Singular set of noncollapsed limits of Einstein $5$-manifolds, and orbifold structure
We refine the structure of the singular set $\mathcal{S}$, and the regularity of the Gromov-Hausdorff limit space $X$. By the uniqueness of $1$-symmetric tangent cones 1.11, we can further stratify the singular set $\mathcal{S}(X)$ as follows:
Theorem 1.15.
There exists some $N_{0}=N_{0}(v)$ depending on the noncollapsing constant $v$ such that
$\mathcal{S}(X)=\mathcal{S}^{0}\sqcup\bigsqcup_{2\leqslant|\Gamma|\leqslant N_{0}}\mathcal{S}^{1}_{\Gamma},$ (1.2)
where $\mathcal{S}_{\Gamma}^{1}$ is the set of points whose unique tangent cone is $\mathds{R}\times\mathds{R}^{4}/\Gamma$, $\Gamma\subset O(4)$ and $\mathcal{S}^{0}$ is a set of countably many points where no tangent cone splits.
Our first result shows that locally each $\mathcal{S}^{1}_{\Gamma}$ is contained in the bi-Lipschitz image of some subset of $\mathds{R}$. In particular, this refines the classical rectifiability statement of [CJN21, JN21], eliminating the need to discard an $\mathcal{H}^{1}$-measure zero subset:
Theorem 1.16 (Singular set contained in Lipschitz manifolds).
Let $(M_{i},g_{i},p_{i})_{i\in\mathbb{N}}$ be a sequence of pointed $5$-manifolds with $|{\rm Ric}|<4$ and ${\rm Vol}(B_{1}(p_{i}))>v>0$. Let $(X,d,p)$ be a Gromov-Hausdorff limit of the sequence. For any $x\in\mathcal{S}^{1}_{\Gamma}$, there exists some $r>0$ and some function $\phi$ such that $\phi:B_{r}(x)\cap\mathcal{S}^{1}_{\Gamma}\to\mathds{R}$ is bi-Lipschitz. In particular, the singular set $\mathcal{S}$ satisfies: There exist countably many points $q_{j}$ and measurable subsets $Z_{i}\subset\mathcal{S}^{1}_{\Gamma_{i}}$ with bi-Lipschitz maps $\phi_{i}:Z_{i}\to\mathds{R}$ such that
$\mathcal{S}=\bigcup_{i=1}^{\infty}Z_{i}\,\,\cup\,\,\bigcup_{j=0}^{\infty}\{q_{j}\}.$ (1.3)
Remark 1.17.
This resolves, in dimension $5$ and under a two-sided Ricci bound, the analogue of [Nab20, Conjecture 2.15]. The original conjecture concerns lower Ricci bounds.
The points $\{q_{j}\}$ in fact coincide with the set of 0-symmetric singularities $\mathcal{S}^{0}$. Theorem 1.16 says that all of the $1$-symmetric singular points lie on Lipschitz curves. It turns out that we understand the singularities completely when they exhaust such a curve in $\mathcal{S}^{1}_{\Gamma}$: they are orbifold singularities along geodesics and the curvature stays bounded. In other words: the curve of singularities in $\mathcal{S}^{1}_{\Gamma}$ is removable. More generally, given any curve inside the singular set $\mathcal{S}$, we have the following structure theorem:
Theorem 1.18 (Structure of curves of singularities).
Let $(M_{i},g_{i},p_{i})$ be a sequence of pointed Einstein $5$-manifolds with $|{\rm Ric}|<4$ and ${\rm Vol}(B_{1}(p_{i}))>v>0$. Let $(X,d,p)$ be a Gromov-Hausdorff limit of the sequence. Assume that there exists a unit-speed curve $\gamma:I=(-\varepsilon,\varepsilon)\to X$ contained in the singular set, $\gamma(I)\subset\mathcal{S}(X)$, for some $\varepsilon>0$. Then, there exists a relatively closed nowhere dense subset $\mathcal{I}\subset I$ such that $I=\mathcal{I}\sqcup\bigsqcup_{j=1}^{\infty}(s_{j},s_{j+1})$ satisfies:
  • on each $(s_{j},s_{j+1})$, there exists a finite group $\Gamma_{j}\subset O(4)$ such that at all of the points in $\gamma(s_{j},s_{j+1})$, the tangent cone is unique and isometric to $\mathbb{R}\times\mathbb{R}^{4}/\Gamma_{j}$,
  • there exist an open sets $U_{j}\subset X$ with $\gamma\big((s_{j},s_{j+1})\big)\subset U_{j}$ such that on $U_{j}\setminus\gamma$ the Einstein metric extends real-analytically across $\gamma$ as a real-analytic $5$-orbifold metric with singularities $\mathbb{R}\times\mathbb{R}^{4}/\Gamma_{j}$ along $\gamma$. In particular, the curvature stays bounded.
  • $\gamma_{|(s_{j},s_{j+1})}$ is a unit-speed geodesic of $X$.
If $\gamma(I)\subset\mathcal{S}^{1}_{\Gamma}$, then $\mathcal{I}=\emptyset$, $\gamma$ is a geodesic and the metric extends as a real-analytic orbifold across it.
Corollary 1.19.
Each connected component $\Omega$ of $\mathcal{S}^{1}_{\Gamma}$ is either a single point or a geodesic, and thus locally convex. That is, for every $x\in\Omega$ there exists $r>0$ such that for any $y\in\Omega\cap B_{r}(x)$, every minimizing geodesic $\gamma$ joining $x$ and $y$ is contained in $\Omega$.
Remark 1.20.
In general, the subset $\mathcal{I}\subset I$ in Theorem 1.18 is non-empty and one cannot expect convexity of the entire singular set $\mathcal{S}$. For example, let $C(Z)$ be a Ricci-flat cone that does not split, where $Z$ is an Einstein $4$-orbifold with at least two singular points. Then the singular set is not locally convex around the vertex.
Theorem 1.18 characterizes any curve of singularities up to a nowhere dense subset: they are open geodesic segments. We now turn to the converse and completely characterize neighborhoods of open geodesics of our limit spaces.
By establishing the sharp Hölder continuity of tangent cones along geodesics in a seminal work [CoN12], Colding-Naber, and later generalized by Deng [Den25], classified the geodesics (meaning locally minimizing, in the metric sense) with one regular point: the whole interior part is regular. Combining Theorems 1.11, 1.18 and their estimates, we can now fully classify the geodesics by one of their interior points:
Theorem 1.21 (Classification of geodesics).
Let $\sigma:(-\varepsilon,\varepsilon)\to X$ be a geodesic in a noncollapsed Einstein $5$-dimensional limit space $(X,d)$ with $|{\rm Ric}_{M_{i}}|\leqslant 4$. Then either $\sigma\big((-\varepsilon,\varepsilon)\big)$ is regular, or the entire $\sigma\big((-\varepsilon,\varepsilon)\big)\subset\mathcal{S}^{1}_{\Gamma}$ for some $\Gamma\subset O(4)$ and the interior of the curve is removable in the sense of Theorem 1.18.
This implies that the entire interior of a geodesic has isometric tangent cones $\mathds{R}\times\mathds{R}^{4}/\Gamma$ for some fixed $\Gamma\subset O(4)$ (with $\Gamma=\{id\}$ corresponding to regular points). In particular, $\mathcal{S}^{0}$ cannot lie on the interior of any geodesic. This yields a dimension-five Einstein-manifold version of the uniqueness of tangent cones along geodesics in the spirit of the conjecture of Burago proved by Petrunin [BGP92, Pet98] for Alexandrov spaces.
Finally we discuss the orbifold regularity of the limit space $X$. Theorem 1.18 says the “interior” of each $\mathcal{S}^{1}_{\Gamma}$ is locally a geodesic along which the metric extends real-analytically. Then it is natural to introduce the orbifold-regular-singular decomposition as follows:
$\mathcal{R}_{\mathrm{orb}}=\mathcal{R}\,\sqcup\bigsqcup_{{2\leqslant|\Gamma|\leqslant N_{0}}}\mathring{\mathcal{S}}^{1}_{\Gamma}\quad\text{ and }\quad\mathcal{S}_{\mathrm{orb}}=X\setminus\mathcal{R}_{\mathrm{orb}},$ (1.4)
where $\mathcal{R}$ is the regular set, and $\mathring{\mathcal{S}}^{1}_{\Gamma}\equiv\{x\in\mathcal{S}^{1}_{\Gamma}~|$~there exists some curve $\sigma:(-\varepsilon,\varepsilon)\to\mathcal{S}^{1}_{\Gamma}$ with $\sigma(0)=x\}.$
Combining the refined rectifiability Theorem 1.16, the orbifold structure Theorem 1.18 and the classification of the geodesics Theorem 1.21, we can prove the following partial orbifold-regularity of $X$:
Corollary 1.22.
The orbifold-regular subset $\mathcal{R}_{\mathrm{orb}}$ and the orbifold-singular subset $\mathcal{S}_{\mathrm{orb}}$ defined in (1.4) satisfy the following:
  • (Orbifold structure) $\mathcal{R}_{\mathrm{orb}}$ is a real-analytic Einstein $5$-orbifold.
  • (Convexity) $\mathcal{R}_{\mathrm{orb}}$ is convex: any geodesic with endpoints in $\mathcal{R}_{\mathrm{orb}}$ is contained in $\mathcal{R}_{\mathrm{orb}}$.
  • (Extreme points) $\mathcal{S}_{\mathrm{orb}}$ is exactly the set of extreme points of $(X,d)$, i.e. the set of points that do not lie on a geodesic between two other points. In particular, the curvature of $(X,d)$ is bounded around every non-extreme point (i.e. points in $X\setminus\mathcal{S}_{\mathrm{orb}}=\mathcal{R}_{\mathrm{orb}}$).
  • (Nowhere denseness) There exist countably many points $\{q_{j}\}$, measurable subsets $Z_{i}\subset X$ and bi-Lipschitz maps $\phi_{i}:Z_{i}\to\mathds{R}$ with $\phi_{i}(Z_{i})$ nowhere dense in $\mathds{R}$, such that $\mathcal{S}_{\mathrm{orb}}=\bigcup_{i}Z_{i}\cup\bigcup_{j}\{q_{j}\}$.
In particular, we show that the orbifold-regular set $\mathcal{R}_{\mathrm{orb}}$ is homeomorphic to a real-analytic variety, thus confirming a conjecture of Naber [Nab14, Conjecture 6.4] away from extreme points in dimension $5$.
1.2.4. Singular set of higher dimensional noncollapsed limits
In higher dimensions we can distinguish the (n4)(n-4)-stratum from the (n5)(n-5)-stratum, and this separation already forces a uniqueness phenomenon.
Theorem 1.23.
Let $(M_{i},g_{i},p_{i})_{i\in\mathbb{N}}$ be a sequence of pointed $n$-manifolds with $|{\rm Ric}_{g_{i}}|\leqslant n-1$ and ${\rm Vol}(B_{1}(p_{i}))\geqslant v>0$, and let $(X,d,p)$ be a Gromov-Hausdorff limit. Fix $x\in X$. Assume that every tangent cone at $x$ splits off an $\mathbb{R}^{n-5}$-factor and that at least one tangent cone at $x$ splits off $\mathbb{R}^{n-4}$. Then the tangent cone at $x$ is unique isometric to $\mathbb{R}^{n-4}\times\mathbb{R}^{4}/\Gamma$ for some $\Gamma\subset O(4)$.
Corollary 1.24.
Let $(X,d)$ be a noncollapsed limit of $6$-manifolds with $|{\rm Ric}|\leqslant 5$. Let $p=\gamma(0)$ be a point lying on an open unit-speed geodesic $\gamma:(-\varepsilon,\varepsilon)\to X$ for some $\varepsilon>0$. Assume that one tangent cone at $p$ splits off $\mathbb{R}^{n-4}$. Then, there exists $\Gamma\subset O(4)$ such that for any $s\in(-\varepsilon,\varepsilon)$, the tangent cone at $\gamma(s)$ is unique and isometric to $\mathbb{R}^{2}\times\mathbb{R}^{4}/\Gamma$.
1.3. Questions and conjectures
Motivated by the new $5$-dimensional results of the present articles, we propose a program towards an orbifold regularity theory off a codimension $5$ set. This is analogous to the energy minimizing harmonic maps: while in this context, the maps are controllably smooth away from a codimension $2$ set, it was proven in [SU82] that the limits are in fact smooth away from a codimension $3$ set.
1.3.1. Dimension 5
Einstein metrics being real-analytic, one might conjecture by unique continuation that $\mathcal{S}_{\mathrm{orb}}$ cannot have accumulation points.
Question 1.25.
Are there noncollapsed limits of Einstein $5$-manifolds with isolated singular points in $\mathcal{S}^{1}$? If singular points in $\mathcal{S}^{1}$ accumulate at a point, do they all lie on a common singular curve?
Remark 1.26.
Motivated by the Calabi-Yau singular models [HN, Sze19], it is reasonable to expect that singularities of the form $\mathbb{R}^{2k}\times\mathbb{R}^{4}/\Gamma$ in Einstein limits can be isolated; the same may plausibly hold for $\mathbb{R}\times\mathbb{R}^{4}/\Gamma$. The question is whether they can accumulate without forming curves. See also [Sze21] in the context of minimal surfaces.
Ruling out the accumulation of singular points, possibly by using the real-analyticity of the Einstein equation, that do not lie on curves of $\mathcal{S}$ would essentially prove the following conjecture.
Conjecture 1.27 (Codimension-$5$ conjecture in dimension $5$).
Let $(X,d)$ be a Gromov-Hausdorff limit of noncollapsed Einstein $5$-manifolds. Then, the set $\mathcal{S}_{\mathrm{orb}}=X\setminus\mathcal{R}_{\mathrm{orb}}$ is countable. In particular, away from a countable set of points, $(X,d)$ has the structure of a smooth Einstein $5$-orbifold with singularities of the form $\mathbb{R}\times\mathbb{R}^{4}/\Gamma$ along countably many geodesics.
Remark 1.28.
In dimension $4$, while the control of the degeneration cannot be better than codimension $4$, the limit spaces are real-analytic orbifolds without singularities up to taking local finite covers. For $6$-dimensional Kähler-Einstein metrics, Section 5 of [DS14] proves that the link of any tangent cone is a 5D Sasaki-Einstein orbifold, whose singular set is finitely many periodic Reeb orbits, and the local quotient by the Reeb action is a Kähler-Einstein 4-orbifold.
Motivated by the Calabi-Yau models [Li19, CR23, Sze19, Sze20a], with tangent cones of the form $\mathbb{R}^{2k}\times\mathbb{R}^{4}/\Gamma$, a natural question is the following one.
Question 1.29.
Are there complete $5$-dimensional Ricci-flat metrics whose asymptotic cones split a line at infinity while the interior metrics do not split a line?
1.3.2. Higher dimensions
In higher dimensions, we have less evidence, but based on our $5$-dimensional results, we ask natural questions of isolation and gap theorems.
Question 1.30 (Higher-dimensional isolation of spherical orbifolds and gap theorem).
Given $(M_{o}^{n},g_{o})$ a spherical orbifold with codimension $4$ singularities. Is there a sequence of smooth Einstein $n$-metrics converging to $(M_{o}^{n},g_{o})$ in the Gromov-Hausdorff sense?
Question 1.31.
For every $n\geq 4,v>0$, does there exist $\varepsilon_{0}=\varepsilon_{0}(n,v)>0$ such that if $(M_{o}^{n},g_{o})$ with ${\rm Vol}(g_{o})>v$ and ${\rm Ric}(g_{o})=(n-1)g_{o}$ satisfies $\lambda_{n-3}<n+\varepsilon_{0}$ with $\lambda_{k}$ the $k^{th}$ eigenvalue of the Laplace-Beltrami operator?
Example 1.32.
The simplest example of such an orbifold is $(M_{o}^{5},g_{o})$ the quotient of $\mathbb{S}^{5}\subset\mathbb{R}^{4}\times\mathbb{R}^{2}$ by the involution $(-\operatorname{Id},\operatorname{Id})$ with a circle of $\mathbb{R}\times\mathbb{R}^{4}/\mathbb{Z}_{2}$ singularities. The orbifold $(M_{o}^{5},g_{o})$ is a $5$-dimensional example, but it is also $\mathbb{S}^{1}$-invariant.
Remark 1.33.
A natural tentative would be to glue in a rescaled cylinder $\mathbb{S}^{1}$ times the Eguchi-Hanson metric and perturb in an $\mathbb{S}^{1}$-invariant way. This is however impossible by our isolation theorem 1.1 since the $\mathbb{S}^{1}$-invariance reduces the problem to that of desingularizing a spherical $4$-orbifold by Einstein $4$-orbifolds.
Question 1.34 (Codimension-$5$ conjecture).
Let $(X,d)$ be a Gromov-Hausdorff limit of noncollapsed Einstein $n$-manifolds. Then there exists a closed set $\mathcal{S}_{\mathrm{orb}}\subset X$ of Hausdorff codimension at least $5$ such that $\mathcal{R}_{\mathrm{orb}}:=X\setminus\mathcal{S}_{\mathrm{orb}}$ is a smooth Einstein $n$-orbifold. Moreover, the singular set in $\mathcal{R}_{\mathrm{orb}}$ is a countable union of totally geodesic codimension $4$ submanifolds, and each point on this set admits a neighborhood modeled on $\mathbb{R}^{n-4}\times\mathbb{R}^{4}/\Gamma$ for some finite $\Gamma\subset O(4)$.
Remark 1.35.
The full singular locus of $X$ is still expected to have (optimal) codimension $4$, but outside the codimension-$5$ “bad set” $\mathcal{S}_{\mathrm{orb}}$ all singularities are removable in the sense that the metric extends as a real-analytic Einstein orbifold metric with bounded curvature across the codimension-$4$ strata. See, the analytic removability and regularity mechanisms in [SU82, BKN89, And89].
Towards a proof in all dimensions. The present paper essentially proves the iteration from n1=4n-1=4 to $n=5$ of a plausible, long-term strategy in order to answer Question 1.34 affirmatively by an iteration on the dimension. Assume that noncollapsed limits of (n1)(n-1)-dimensional Einstein metrics have codimension $5$ regularity in the sense of Question 1.34.
  1. (1)
    Isolation of higher-dimensional spherical orbifolds. Show that (n1)(n-1)-dimensional spherical suspensions over $\mathbb{S}^{3}/\Gamma$ are quantitatively GH-isolated among noncollapsed limits of (n1)(n-1)-dimensional Einstein metrics with ${\rm Ric}=(n-2)$.
  2. (2)
    Isolation of (n4)(n-4)-symmetric cones. Deduce that cones whose links are spherical suspensions over $\mathbb{S}^{3}/\Gamma$ (equivalently, local models $\mathbb{R}^{n-4}\times\mathbb{R}^{4}/\Gamma$) are quantitatively isolated among Ricci-limit cones.
  3. (3)
    Singular set contained in Lipschitz submanifolds. Show that $\mathcal{S}^{n-4}$ is contained in a union of Lipschitz codimension $4$ submanifolds.
  4. (4)
    Removable codimension $4$ singularities. Show that subsets of $\mathcal{S}^{n-4}$ that exhaust (n4)(n-4)-dimensional subsets of the Lipschitz maps must be removable.
  5. (5)
    Codimension $5$ singularities. Show that the remaining singularities in $\mathcal{S}^{n-4}$ form a subset of Hausdorff dimension (n5)(n-5).
1.4. Strategy and Organization
The article is divided into two parts, the first one concerns gap theorems for Einstein orbifolds, and the second the regularity of noncollapsed Einstein $5$-manifolds.
1.4.1. Isolation theorem for Einstein $4$-manifolds
We first outline the proof of the isolation theorem for spherical and hyperbolic Einstein $4$-orbifolds, Theorem 1.1.
Setup and refined approximation. Assume towards a contradiction that there exists a sequence of Einstein manifolds $(M_{i},g_{i})_{i}$ converging in the Gromov-Hausdorff (GH) sense to a spherical or hyperbolic Einstein $4$-orbifold $(M_{o},g_{o})$. By the GH-to-smooth theory of [Ozu22a, Ozu22b], each $g_{i}$ is smoothly approximated by a naïve gluing metric $g_{i}^{D}$ capturing all bubble trees of ALE Ricci-flat pieces. A key new input is a refined obstruction space adapted to these trees. We first glue and perturb every bubble tree to a Ricci-flat modulo obstructions configuration, and then upgrade $g_{i}^{D}$ to an Einstein modulo obstructions approximation $g_{i}^{A}$ with an explicit refined obstruction $w_{i}^{A}$ supported on the bubble region.
Quantitative control. In appropriate, divergence-free gauges, the linearization of the Einstein (mod. obstructions) map is invertible between specific weighted Hölder spaces, hence, by inverse function theorem,
$\|g_{i}^{A}-g_{i}\|+\|w_{i}^{A}\|\;\lesssim\;\|{\rm Ric}(g_{i}^{A})-\Lambda g_{i}^{A}+w_{i}^{A}\|,$ (1.5)
where $\Lambda\in\{\pm 3\}$ is the Einstein constant of $(M_{o},g_{o})$. Thus, to reach a contradiction, it suffices to prove that $\|w_{i}^{A}\|$ cannot be dominated by the right-hand side.
Decomposition of the obstruction. A precise analysis of the refined obstruction shows a canonical splitting
$w_{i}^{A}=u_{i}^{B}+u_{i}^{o},$
where $u_{i}^{B}$ is the obstruction to perturbing the bubble trees toward Ricci-flat metrics, and $u_{i}^{o}$ is the obstruction created when matching the trees to the core orbifold $(M_{o},g_{o})$. When $(M_{o},g_{o})$ is spherical or hyperbolic, then $u_{i}^{o}$ is nonvanishing. The main conceptual step of the proof then consists in proving that if $u_{i}^{B}$ compensates $u_{i}^{o}$, then it contains a much larger piece orthogonal to $u_{i}^{o}$. In either case this contradicts (1.5):
$\|w_{i}^{A}\|\ \gg\ \|{\rm Ric}(g_{i}^{A})-\Lambda g_{i}^{A}+w_{i}^{A}\|.$
Conclusion. No Einstein orbifold can be GH-close to $(M_{o},g_{o})$ unless it is isometric to it.
1.4.2. Regularity of Einstein $5$-manifolds
We then focus on the regularity of Einstein $5$-manifolds.
Uniqueness of tangent cones. In the 5-dimensional limit space, we first establish an isolation theorem for $1$-symmetric cones, building on the gap theorems for Einstein $4$-orbifolds. This, in particular, yields the uniqueness of $1$-symmetric tangent cones. Combined with $\varepsilon$-regularity, we then obtain quantitative curvature estimates on wedge regions around such points, and analogous results hold for tangent cones at infinity and their asymptotic curvature behavior. These results are proved in Section 4.2.
Refined rectifiability. The isolation theorem implies that $1$-symmetry at a given scale at some point $x\in\mathcal{S}^{1}_{\Gamma}$ propagates to uniform symmetry across all smaller scales. Combined with the neck decomposition developed in [JN21, CJN21], this uniform symmetry allows one to identify the local $1$-stratum $S^{1}_{\Gamma}$ near $x$ with the center $\mathcal{C}_{0}$ of an associated neck region. A bi-Lipschitz coordinate is then constructed by using a splitting function to order points along $\mathcal{C}_{0}$ and pushing forward the Ahlfors regular measure associated to the neck region onto a subset of $\mathds{R}$ via the ordering. A countable selection of such parametrizations yields a covering of $\mathcal{S}^{1}$ by bi-Lipschitz images of subsets of $\mathbb{R}$, thereby establishing a refined rectifiability of the singular set without discarding any exceptional subsets. Moreover, we show that along each curve with constant volume density, the tangent cones remain identical. The proof is given in Section 4.3.
Orbifold structure. Along a Lipschitz curve of singularity type $\mathds{R}\times\mathds{R}^{4}/\Gamma$, GH-close model charts to $\mathbb{R}\times\mathbb{R}^{4}/\Gamma$ can be glued horizontally along the direction parallel to the singular curve. Controlling the gluing error uniformly is subtle: it requires compatibility of the local $\Gamma$-actions on adjacent charts. Crucially, the transition maps are constrained to be approximated only by the reflection or translation along the singular axis $\mathds{R}$ and rotation preserving the $\Gamma$-action on $\mathds{R}^{4}/\Gamma$ – a property that fails in flat $\mathds{R}^{5}$. These glued charts then yield $C^{0}$ orbifold-type coordinates on punctured neighborhoods.
To control the curvature in these necks, we apply Hardy-type inequalities and a Moser iteration for subsolutions of the following equation
$\Delta u\gtrsim-|y|^{-2}u$ (1.6)
with $|y|$ the distance to the singular curve. Equation (1.6) is satisfied by $u=|{\rm Rm}|^{1-\alpha}$ for some $\alpha>0$ by 1) $\varepsilon$-regularity and our isolation theorem for $\mathbb{R}\times\mathbb{R}^{4}/\Gamma$ cones, 2) the differential equation satisfied by curvature on Einstein metrics, 3) an improved Kato inequality. This gives uniform pointwise curvature bounds. Once curvature is uniformly bounded, we upgrade the coordinate regularity to $C^{1,\alpha}$, and then to analyticity via the Einstein equations. This establishes the analytic orbifold structure along the singular curves. Thanks to these coordinates, the singular curve, which is the fixed-point set of the $\Gamma$-action, is a geodesic. The details are presented in Section 5.
1.5. Acknowledgements
The first author is grateful to Tobias Colding for suggesting the uniqueness problem and for his patient guidance and many insightful discussions. The first author also thanks Wenshuai Jiang for his support and helpful conversations. The authors would like to thank Max Hallgren and Wenshuai Jiang for very valuable comments that helped us improve a step of the proof of Theorem 1.16 in Section 4.3.
During this project, Yiqi Huang was partially supported by a Simons Dissertation Fellowship and by NSF Grant DMS–2405393, and Tristan Ozuch was partially supported by the National Science Foundation under grant DMS–2405328.
Part I Isolation and gap theorems for Einstein $4$-orbifolds
2. Preliminaries on the degeneration of Einstein $4$-manifolds
2.1. Einstein orbifolds and their infinitesimal deformations
For $\Gamma$ a finite subgroup of $O(4)$ acting freely on $\mathbb{S}^{3}$, let us denote $(\mathbb{R}^{4}/\penalty 50\Gamma,e)$ the flat cone obtained by the quotient by the action of $\Gamma$, and $r_{e}:=d_{e}(.,0)$.
Definition 2.1 ($4$-orbifold (with isolated singularities)).
We will say that a metric space $(M_{o},g_{o})$ is an orbifold of dimension $4$ if there exists $\varepsilon_{0}>0$ and a finite number of points $(p_{j})_{j}$ of $M_{o}$ called singular such that we have the following properties:
  1. (1)
[…]
  2. (2)
    for each singular point $p_{j}$ of $M_{o}$, there exists a neighborhood of $p_{j}$, $U_{j}\subset M_{o}$, a finite subgroup acting freely on $\mathbb{S}^{3}$, $\Gamma_{j}\subset O(4)$, and a diffeomorphism $\Phi_{j}:B_{e}(0,\varepsilon_{0})\subset\mathbb{R}^{4}/\penalty 50\Gamma_{j}\to U_{j}\subset M_{o}$ for which, the pull-back of $\Phi_{j}^{*}g_{o}$ on the covering $\mathbb{R}^{4}$ is smooth.
Definition 2.2 (The function $r_{o}$ on an orbifold).
Given $S_{o}\subset M_{o}$ a subset of the singularities, we define $r_{o}$, a function on $M_{o}$ smooth away from $S_{o}$ satisfying $r_{o}:=(\Phi_{j})_{*}r_{e}$ on each $U_{j}$ with $U_{j}\cap S_{o}\neq\emptyset$, and such that on $M_{o}\backslash U_{j}$, we have $\varepsilon_{0}\leqslant r_{o}<1$ (the different choices will be equivalent for our applications). We will denote, for $0<\varepsilon\leqslant\varepsilon_{0}$,
$M_{o}(\varepsilon):=\{r_{o}>\varepsilon\}=M_{o}\backslash\Big(\bigcup_{j}\Phi_{j}\Big(\overline{B_{e}(0,\varepsilon)}\Big)\Big).$
We also define $\mathbf{O}(g_{o})$ the space of traceless and divergence-free infinitesimal Einstein deformations of $g_{o}$.
2.2. Ricci-flat ALE spaces and their deformations
Let us now turn to ALE Ricci-flat metrics.
Definition 2.3 (ALE orbifold (with isolated singularities)).
An ALE orbifold of dimension $4$, $(N,g_{b})$ is a metric space for which there exists $\varepsilon_{0}>0$, singular points $(p_{j})_{j}$ and a compact $K\subset N$ for which we have:
  1. (1)
    $(N,g_{b})$ is an orbifold of dimension $4$,
  2. (2)
    there exists a compact subset $K\subset N$ and a diffeomorphism $\Psi_{\infty}:(\mathbb{R}^{4}/\penalty 50\Gamma_{\infty})\backslash\overline{B_{e}(0,\varepsilon_{0}^{-1})}\to N\backslash K$, for $\Gamma_{\infty}\subset O(4)$ such that r_e^l—^l(Ψ_^* g_b - e)—_eC_l r_e^-4.
Definition 2.4 (The function $r_{b}$ on an ALE orbifold).
[…]
For $0<\varepsilon\leqslant\varepsilon_{0}$, we will denote
[…]
We define $\mathbf{O}(g_{b})$ to be the set of infinitesimal Ricci-flat deformations which are transverse and traceless.
2.3. Naïve desingularizations
We recall the definition of a naïve desingularization of an orbifold, [Ozu22a].
2.3.1. Gluing of ALE spaces to orbifold singularities
Let $0<2\varepsilon<\varepsilon_{0}$ be a fixed constant, $t>0$ a scale of gluing, $(M_{o},g_{o})$ an orbifold and $\Phi:B_{e}(0,\varepsilon_{0})\subset\mathbb{R}^{4}/\penalty 50\Gamma\to U$ a local chart of Definition 2.1 around a singular point $p\in M_{o}$. Let also $(N,g_{b})$ be an ALE orbifold asymptotic to $\mathbb{R}^{4}/\penalty 50\Gamma$, and $\Psi_{\infty}:(\mathbb{R}^{4}/\penalty 50\Gamma)\backslash B_{e}(0,\varepsilon_{0}^{-1})\to N\backslash K$ a chart at infinity of Definition 2.3.
For $s>0$, define $\phi_{s}:x\in\mathbb{R}^{4}/\penalty 50\Gamma\to sx\in\mathbb{R}^{4}/\penalty 50\Gamma$. For $t<\varepsilon_{0}^{4}$, we define $M_{o}\#N$ as $N$ glued to $M_{o}$ thanks to the diffeomorphism
$\Phi\circ\phi_{\sqrt{t}}\circ\Psi^{-1}:\Psi(A_{e}(\varepsilon_{0}^{-1},\varepsilon_{0}t^{-1/2}))\to\Phi(A_{e}(\varepsilon_{0}^{-1}\sqrt{t},\varepsilon_{0})).$
Consider moreover $\chi:\mathbb{R}^{+}\to\mathbb{R}^{+}$, a $C^{\infty}$ cut-off function supported on $[0,2]$ and equal to $1$ on $[0,1]$.
Definition 2.5 (Naïve gluing of an ALE space to an orbifold).
[…]
$g_{o}\#_{p,t}g_{b}:=t\chi(t^{-\frac{1}{4}}r_{e})\phi_{t^{-1/2}}^{*}(\Psi_{\infty}^{*}g_{b})+\Big(1-\chi(t^{-\frac{1}{4}}r_{e})\Big)\Phi^{*}g_{o}.$ (2.1)
Given two tensors $h_{o}$ and $h_{b}$ on $M_{o}$ and $N$, we similarly define the naïve gluing:
$h_{o}\#_{p,t}h_{b}:=t\chi(t^{-\frac{1}{4}}r_{e})\phi_{t^{-1/2}}^{*}\Psi_{\infty}^{*}h_{b}+\Big(1-\chi(t^{-\frac{1}{4}}r_{e})\Big)\Phi^{*}h_{o}$ (2.2)
In general, one has to desingularize by trees of Ricci-flat ALE orbifolds. Consider $(M_{o},g_{o})$ an Einstein orbifold (the index $o$ stands for orbifold) and $S_{o}$ a subset of its singular points, and $(N_{j},g_{b_{j}})_{j}$ (the index $b_{j}$ stands for $j$-th bubble) a family of Ricci-flat ALE spaces asymptotic at infinity to $\mathbb{R}^{4}/\penalty 50\Gamma_{j}$ and $(S_{b_{j}})_{j}$ subsets of their singular points. Now, assume that there is a one-to-one map $p:j\mapsto p_{j}\in S_{o}\cup\bigcup_{k}S_{b_{k}}$, where the singularity at $p_{j}$ is $\mathbb{R}^{4}/\penalty 50\Gamma_{j}$. We will call $D:=\big((M_{o},g_{o},S_{o}),(N_{j},g_{b_{j}},S_{b_{j}})_{j},p\big)$ a desingularization pattern.
Definition 2.6 (Naïve (partial) desingularization by a tree of singularities).
Let $0<2\varepsilon<\varepsilon_{0}$, $D$ be a desingularization pattern for $(M_{o},g_{o})$, and let $0<t_{j}<\varepsilon^{4}$ be relative gluing scales. The metric $g^{D}_{t}$ is then the result of the following finite iteration:
  1. (1)
    start with a deepest bubble $(N_{j},g_{b_{j}})$, that is, $j$ such that $S_{b_{j}}=\emptyset$,
  2. (2)
[…]
  3. (3)
    choose another deepest bubble and do the same.
For $t=(t_{j})_{j}$, if $N_{j}$ is glued to $p_{j}\in N_{j_{1}}$, and $N_{j_{1}}$ is glued to $p_{j_{1}}\in N_{j_{2}}$, …, $N_{j_{k-1}}$ is glued to $p_{j_{k-1}}\in N_{j_{k}}$, which is glued to $M_{o}$, we define the absolute scale $T_{j}:=t_{j}t_{j_{1}}t_{j_{2}}...t_{j_{k}}$. This way, on each $N_{j}(\varepsilon)$, $g^{D}_{t}|_{N_{j}(\varepsilon)}=T_{j}g_{b_{j}}$.
Remark 2.7.
The above definition considers partial desingularizations by considering subsets $S_{o}$ and $S_{b_{j}}$ of all singularities. The resulting metric is only smooth if $S_{o}$ and the $S_{b_{j}}$ are all of the singularities.
Definition 2.8 (Tree of singularities forming at a given singular point).
Let $p_{1}\in S_{o}$ be a singular point of $M_{o}$ (resp. $p_{1}\in S_{b_{j}}$), and consider $J_{p_{1}}$ the set of indices $j_{0}$ such that in the above desingularization pattern, there is a path from $j_{0}$ to $p_{1}$, that is a set $j_{1},\dots,j_{m}$ such that $p(j_{k})\in S_{b_{j_{k+1}}}$ for $0\leqslant k\leqslant m$, and $p(j_{m})=p_{1}$.
We call $\mathcal{B}_{p_{1}}=\left((N_{j},g_{b_{j}},S_{b_{j}},t_{j})_{j},p_{1}\right)$ the tree of singularities at $p_{1}$. It yields a topology $N_{J_{p_{1}}}$ given by the above gluing and admits an ALE metric $g^{B}_{t}$ equal to $t^{-1}_{p^{-1}(p_{1})}g^{D}_{t}$ where both metrics are defined.
We additionally call $(N_{j_{m}},g_{b_{j_{m}}})$ the core orbifold of the tree. It plays the role of $(M_{o},g_{o})$ in all of the previous constructions.
Let $(M,g^{D})$ a naïve desingularization of an Einstein orbifold $(M_{o},g_{o})$ by a tree of ALE Ricci-flat orbifolds $(N_{j},g_{b_{j}})_{j}$ with relative scales $t_{j}>0$. The manifold $M$ is covered as $M=M_{o}^{t}\cup\bigcup_{j}N_{j}^{t}$, where
$M_{o}^{t}:=M_{o}\backslash\Big(\bigcup_{j}\Phi_{j}(B_{e}(0,t_{j}^{\frac{1}{4}}))\Big),$ (2.3)
where $t_{j}>0$ is the relative gluing scale of $N_{k}$ at the singular point $p_{k}\in M_{o}$, and where
$N_{j}^{t}:=\Big(N_{j}\backslash\Psi_{\infty}\Big((\mathbb{R}^{4}/\penalty 50\Gamma_{\infty})\backslash B_{e}(0,2t_{j}^{-\frac{1}{4}})\Big)\Big)\backslash\Big(\bigcup_{k}\Psi_{k}(B_{e}(0,t_{k}^{\frac{1}{4}})\Big).$ (2.4)
On $M_{o}^{16t}\subset M_{o}^{t}$, we have $g^{D}=g_{o}$ and on each $N_{j}^{16t}\subset N_{j}^{t}$, we have $g^{D}=T_{j}g_{b_{j}}$. By Definition 2.5, on the intersection $N_{j}^{t}\cap M_{o}^{t}$ we then have $\sqrt{T_{j}}r_{b_{j}}=r_{o}$, and on the intersection $N_{j}^{t}\cap N_{k}^{t}$, we have $\sqrt{T_{j}}r_{b_{j}}=\sqrt{T_{k}}r_{b_{k}}$.
Definition 2.9 (Function $r_{D}$ on a naïve desingularization).
On a naïve desingularization $(M,g^{D})$, we define a smooth function $r_{D}$ in the following way:
  1. (1)
    $r_{D}=r_{o}$ on $M_{o}^{t}$,
  2. (2)
    $r_{D}=\sqrt{T_{j}}r_{b_{j}}$ on each $N_{j}^{t}$.
As in [Ozu22b, Definitions 1.11, 1.12], we decompose our orbifold $M$ into particular regions and define associated cut-off functions.
Definition 2.10 (Neck regions, $\mathcal{A}_{k}(t,\varepsilon)$).
Let $(N_{k},g_{b_{k}})$ be a Ricci-flat ALE orbifold of the above tree of singularities. We define $\mathcal{A}_{k}(t,\varepsilon)$ as the connected region with $\varepsilon^{-1}\sqrt{T_{k}}<\sqrt{T_{k}}r_{b_{k}}=r_{D}<\varepsilon t_{k}^{-\frac{1}{2}}\sqrt{T_{k}}$ with a nonempty intersection with $N_{k}^{t}$.
Definition 2.11 (Cut-off functions $\chi_{M_{o}^{t}}$, $\chi_{N_{j}^{t}}$, $\chi_{\mathcal{A}_{k}(t,\varepsilon)}$ and $\chi_{B(p_{k},\varepsilon)}$).
We define the following cut-off functions thanks to the cut-off function $\chi$ used in Definition 2.5.
  • $\chi_{M_{o}^{t}}$, equal to $1$ on $M_{o}^{16t}$ and equal to $1-\chi(t_{k}^{-\frac{1}{4}}r_{o})$ on each annulus $\mathcal{A}_{k}(t,\varepsilon)$. It is supported on $M_{o}^{t}$.
  • $\chi_{N_{j}^{t}}$, equal to $1$ on $N_{j}^{16t}$ and equal to $1-\chi(t_{k}^{-\frac{1}{4}}r_{b_{j}})$ on each annulus $\mathcal{A}_{k}(t,\varepsilon)$ at its singular points and $\chi(t_{j}^{\frac{1}{4}}r_{b_{j}})$ in a neighborhood of infinity. It is supported on $N_{j}^{t}$.
  • $\chi_{\mathcal{A}_{k}(t,\varepsilon)}$ equal to $1$ on $\mathcal{A}_{k}(t,\frac{1}{2}\varepsilon)$, and equal to $\chi(\varepsilon^{-1}t_{k}^{\frac{1}{2}}r_{b_{k}})-\chi(\varepsilon r_{b_{k}})$. It is supported on $\mathcal{A}_{k}(t,\varepsilon)$.
  • $\chi_{B(p_{k},\varepsilon)}$ for $p_{k}\in M_{o}$ equal to $1$ on $r_{o}<\varepsilon$ equal to $\chi(\varepsilon^{-1}r_{o})$ around $p_{k}$. It is supported in supported in $r_{o}<2\varepsilon$ around $p_{k}$.
2.3.2. Refined approximate kernel of the linearization
One of the subtleties of our proof is that we need to carefully estimate the divergence of our tensors at several stages. This requires a refined notion of obstructions from that of [Ozu22b, Definition 4.5].
Extending elements of $\mathbf{O}(g_{o})$ or of $\mathbf{O}(g_{b_{j}})$ to deeper bubbles.
Definition 2.12 (Extension of elements of $\mathbf{O}(g_{o})$ or of $\mathbf{O}(g_{b_{j}})$).
  1. (1)
    Consider an element of $\mathbf{o}\in\mathbf{O}(g_{o})$ (resp. of $\mathbf{o}\in\mathbf{O}(g_{b_{j}})$).
  2. (2)
    For each $p_{[0]}\in S_{o}$ (resp. $p_{[0]}\in S_{b_{j}}$) consider the constant $2$-tensor $H_{[0]}:=\mathbf{o}(p_{[0]})$ on the cone.
  3. (3)
    Consider $j_{[0]}=p^{-1}(p_{[0]})$ and the Ricci-flat ALE metric $(N_{j_{[0]}},g_{b_{j_{[0]}}})$ and consider the unique symmetric $2$-tensor $\mathbf{h}_{j_{[0]}}$ asymptotic to $H_{[0]}$ (obtained as in [Ozu21a, Lemma 2.7]) and satisfying:
    $P_{g_{b_{j_{[0]}}}}\mathbf{h}_{j_{[0]}}=0,\qquad\delta_{g_{b_{j_{[0]}}}}\mathbf{h}_{j_{[0]}}=0,\qquad\operatorname{tr}_{g_{b_{j_{[0]}}}}\mathbf{h}_{j_{[0]}}=0.$ (2.5)
  4. (4)
    We then iterate: if for $k\geqslant 0$, $\mathbf{h}_{j_{[k]}}$ is defined, consider $p_{[k]}\in S_{b_{j_{[k]}}}$, and $j_{[k+1]}=p^{-1}(p_{[k]})$ and the constant $2$-tensor $H_{[k]}:=\mathbf{h}_{j_{[k]}}(p_{[k]})$. We then define $\mathbf{h}_{j_{[k+1]}}$ satisfying
    $P_{g_{b_{j_{[k+1]}}}}\mathbf{h}_{j_{[k+1]}}=0,\qquad\delta_{g_{b_{j_{[k+1]}}}}\mathbf{h}_{j_{[k+1]}}=0,\qquad\operatorname{tr}_{g_{b_{j_{[k+1]}}}}\mathbf{h}_{j_{[k+1]}}=0.$ (2.6)
We therefore have tensors $\mathbf{h}_{k}$ defined on each $(N_{k},g_{b_{k}})$ for $k\in J_{p(j)}$.
Projecting elements of $\mathbf{O}(g_{o})$ or of $\mathbf{O}(g_{b_{j}})$ to the naïve desingularization $(M,g^{D})$.
We define the projection of an element $\mathbf{o}_{o}\in\mathbf{O}(g_{o})$ on the naïve desingularization:
$\mathrm{pr}(\mathbf{o}_{o}):=\tilde{\mathbf{o}}_{o,t}:=\chi_{M_{o}^{t}}\mathbf{o}_{o}+\sum_{k}\chi_{N_{k}^{t}}\mathbf{h}_{k},$
where the $\mathbf{h}_{k}$ are defined in Definition 2.12. Similarly, for $\mathbf{o}_{j}\in\mathbf{O}(g_{b_{j}})$,
$\mathrm{pr}(\mathbf{o}_{j}):=\tilde{\mathbf{o}}_{j,t}:=\chi_{N_{j}^{t}}\mathbf{o}_{j}+\sum_{k\in J_{p(j)}}\chi_{N_{k}^{t}}\mathbf{h}_{k}.$
Definition 2.13 (Space of refined obstructions).
Let $(M,g^{D}_{t})$ be a naïve desingularization of a (compact or ALE) Einstein orbifold $(M_{o},g_{o})$. On $M$, we will denote
$\tilde{\mathbf{O}}(g^{D}):=\Big\{\tilde{\mathbf{o}}_{o,t}+\sum_{j}\tilde{\mathbf{o}}_{j,t},\;\mathbf{o}_{o}\in\mathbf{O}(g_{o}),\;\mathbf{o}_{j}\in\mathbf{O}(g_{b_{j}})\Big\}.$
We have also defined a one-to-one map $\mathrm{pr}:\mathbf{O}(g_{o})\sqcup\bigsqcup_{j}\mathbf{O}(g_{b_{j}})\to\tilde{\mathbf{O}}(g^{D})$.
The reason to refine our obstructions is to improve how divergence-free they are. This also makes them better approximations of infinitesimal Einstein deformations, i.e. their image by the linearization operator $P_{g^{D}}$, defined in (2.4.1), are smaller than in [Ozu22b, Lemma 4.5]. In particular, the proofs of [Ozu22b] work with the above Definition 2.13 for $\tilde{\mathbf{O}}(g^{D})$.
Lemma 2.14.
Whether $g_{o}$ is compact or ALE, we have the following controls on the elements of $\tilde{\mathbf{o}}\in\tilde{\mathbf{O}}(g^{D})$: for $C>0$ depending on $g_{o}$ and $g_{b_{j}}$, with $t_{\max}=\max_{j}t_{j}$ and with the function spaces of Section A,
$\|\delta_{g^{D}}\tilde{\mathbf{o}}\|_{r_{D}^{-1}C^{1,\alpha}_{\beta}}<Ct_{\max}^{\frac{2-\beta}{4}}\|\tilde{\mathbf{o}}\|_{C^{2,\alpha}_{\beta,*}},\quad\|\delta_{g^{D}}\tilde{\mathbf{o}}\|_{r_{D}^{-3}C^{1,\alpha}_{\beta}}<Ct_{\max}^{\frac{2-\beta}{4}}\|\tilde{\mathbf{o}}\|_{L^{2}(g^{B})}\quad\text{and }\quad\|P_{g^{D}}\tilde{\mathbf{o}}\|_{r_{D}^{-2}C^{0,\alpha}_{\beta}}<Ct_{\max}^{\frac{2-\beta}{4}}\|\tilde{\mathbf{o}}\|_{C^{2,\alpha}_{\beta,*}}.$ (2.7)
Proof.
First note that on $M^{16t}_{o}$ and each $N_{j}^{16t}$, one has $\delta_{g^{D}}\tilde{\mathbf{o}}=0$ and $P_{g^{D}}\tilde{\mathbf{o}}=0$. There remains to understand the element in the annular regions $M^{t}_{o}\backslash M^{16t}_{o}$ and $N^{t}_{j}\backslash N^{16t}_{j}$.
In such regions, where $\frac{1}{2}\sqrt{T_{k}}t_{k}^{-\frac{1}{4}}<r_{D}<2\sqrt{T_{k}}t_{k}^{-\frac{1}{4}}$ one has
$|\nabla^{l}_{g_{e}}(g^{D}-g_{e})|\leqslant Cr_{D}^{-l}t_{k}^{\frac{1}{2}}$ (2.8)
for $l\leqslant 4$ and $C>0$. At the same time, by construction, for a $g_{e}$-parallel symmetric $2$-tensor $H$, we have
$|\nabla_{e}^{l}(\tilde{\mathbf{o}}_{o}-H)|\leqslant Cr_{D}^{-l}t_{k}^{\frac{1}{2}},$ (2.9)
for $l\leqslant 4$ and $C>0$ on these regions $M^{t}_{o}\backslash M^{16t}_{o}$ and $N^{t}_{j}\backslash N^{16t}_{j}$ where $\frac{1}{2}\sqrt{T_{k}}t_{k}^{-\frac{1}{4}}<r_{D}<2\sqrt{T_{k}}t_{k}^{-\frac{1}{4}}$.
Combining (2.8) and (2.9), we find for $C>0$ independent of $\tilde{\mathbf{o}}$ and $l\leqslant 2$
$\nabla^{l}(\delta_{g^{D}}\tilde{\mathbf{o}})\leqslant Cr_{D}^{-1-l}t_{k}^{\frac{1}{2}},\qquad\text{ and }\qquad\nabla^{l}(P_{g^{D}}\tilde{\mathbf{o}})\leqslant Cr_{D}^{-2-l}t_{k}^{\frac{1}{2}},$
which yields the result. The control by the $L^{2}$ norm comes from the fact that the $L^{2}(g^{B})$ norm and the $r_{B}^{-2}C^{l,\alpha}_{\beta}(g^{B})$ for $l\geqslant 0$ norms are equivalent independently of $t$, see [Ozu22b, Remark 4.45]. ∎
2.4. Refined compactness of noncollapsed Einstein $4$-manifolds
Let us explain the results from [Ozu22a, Ozu22b, Ozu20] which we will need in the rest of the article. Consider for $t=(t_{j})_{j}$, a naïve gluing $g^{D}_{t}$ as in Definition 2.5 and its associated space of obstructions $\tilde{\mathbf{O}}(g^{D}_{t})$.
2.4.1. Einstein modulo obstruction perturbation of a naïve gluing
Define $\tilde{\delta}_{g}=\pi_{\tilde{K}_{o}^{\perp}}\delta_{g}$ is the $L^{2}(g^{D}_{t})$-projection of the divergence orthogonally to the set $\tilde{K}_{o}^{\perp}$ of cut-offs of the Killing vector fields of $(M_{o},g_{o})$, whose definition we skip, see [Ozu22b, Section 3.2]. The operator $\tilde{\delta}_{g}\delta_{g}^{*}$ is invertible in the relevant weighted space by [Ozu22b, Proposition 3.3]. We define $\pi_{g}:=1-\delta_{g}^{*}(\delta_{g}\delta_{g}^{*})^{-1}\tilde{\delta}_{g}$, the $L^{2}(dv_{g})$-projection on $\ker\tilde{\delta}_{g}$. Thanks to the Einstein operator $\mathrm{E}(g)={\rm Ric}(g)-\frac{\mathrm{R}(g)}{2}g$, we introduce the $g^{D}_{t}$-gauged Einstein operator
$\mathbf{\Phi}_{g^{D}_{t}}(g)=\mathrm{E}(g)+\delta_{g}^{*}\tilde{\delta}_{g^{D}_{t}}g.$
Theorem 2.15 ([Ozu22b, Theorem 4.6 and Remark 4.56]).
There exists $\varepsilon_{0}>0$ such that for any $0<t<\varepsilon_{0}^{4}$ and any $v\in\tilde{\mathbf{O}}(g^{D}_{t})$ with $|v|\leqslant\varepsilon_{0}$, there is a unique solution $(\hat{g},\hat{w})=(\hat{g}_{t,v},\hat{w}_{t,v})\in C^{2,\alpha}_{\beta,*}(g^{D}_{t})\times\tilde{\mathbf{O}}(g^{D}_{t})$ with $\|\hat{g}-g^{D}_{t}\|_{C^{2,\alpha}_{\beta,**}}{<\varepsilon_{0}}$ (the norm $C^{2,\alpha}_{\beta,**}$} is recalled in Definition A.4) and to the Einstein modulo obstructions equation close to $g^{D}_{t}$: satisfying $\hat{g}-(g^{D}_{t}+v)\perp_{L^{2}(g^{D}_{t})}\tilde{\mathbf{O}}(g^{D}_{t})$ and
$\mathbf{\Phi}_{g^{D}_{t}}(\hat{g})+\pi_{\hat{g}}\hat{w}=0.$ (2.10)
The resulting metric $\hat{g}$ is Einstein if and only if the obstruction $\hat{w}$ vanishes.
Remark 2.16.
By the Bianchi identity, taking $\tilde{\delta}_{\hat{g}}$ of (2.10) we find: $(\tilde{\delta}_{\hat{g}}\delta_{\hat{g}}^{*})\tilde{\delta}_{g^{D}_{t}}{\hat{g}}=0,$ hence $^{\hat{g}}$ automatically solves the gauge condition $\tilde{\delta}_{g^{D}_{t}}{\hat{g}}=0$ by the invertibility of $\delta_{\hat{g}}\delta_{\hat{g}}^{*}$. We could therefore replace $\mathbf{\Phi}_{g^{D}_{t}}({\hat{g}})+\pi_{\hat{g}}\hat{w}=0$ by $\mathrm{E}({\hat{g}})+\pi_{\hat{g}}\hat{w}=0$ and $\tilde{\delta}_{g^{D}_{t}}{\hat{g}}=0$.
Remark 2.17.
Here, we considered the gauged notion of Einstein modulo obstructions metric of [Biq13, Section 6] and [Ozu20, Remark 4.56], but different from [Ozu22b].
The linearization of $g\mapsto\mathbf{\Phi}_{g^{D}}(g)$ at $g^{D}$is
$\bar{P}_{{g^{D}}}(h):=\;d_{g^{D}}\mathbf{\Phi}_{g^{D}}(h)=$ $\frac{1}{2}\Big(\nabla^{*}_{g^{D}}\nabla_{g^{D}}h-2\mathring{\mathrm{R}}_{g^{D}}(h)+{\rm Ric}_{g^{D}}\circ h+h\circ{\rm Ric}_{g^{D}}-\mathrm{R}_{g^{D}}h+\langle{\rm Ric}_{g^{D}},h\rangle_{g^{D}}{g^{D}}\Big)$
$-2\delta_{g^{D}}^{*}\delta_{g^{D}}h+2\delta^{*}_{{g^{D}}}\tilde{\delta}_{g^{D}}h-(\delta_{g^{D}}\delta_{g^{D}}h){g^{D}}+(\Delta_{g^{D}}\textup{tr}_{g^{D}}h){g^{D}}-\nabla_{g^{D}}^{2}\textup{tr}_{g^{D}}h,$ (2.11)
in dimension $4$. If $g^{D}$ were an Einstein metric and $h$ a divergence-free and traceless symmetric $2$-tensor, then the operator $\bar{P}_{g^{D}}$ would reduce to $P_{g^{D}}:=\frac{1}{2}\nabla_{g^{D}}^{*}\nabla_{g^{D}}-\mathring{\mathrm{R}}_{g^{D}}.$ The proof of Theorem 2.15 relies on a Lyapunov-Schmidt reduction and an application of the implicit function theorem between Banach spaces $C^{2,\alpha}_{\beta,*}$ and $r_{D}^{-2}C^{\alpha}_{\beta}$. There is a quantitative uniqueness associated with Theorem 2.15 as it comes from an application of the implicit function theorem, see [Biq13, Section 9] and [Ozu22b, Section 4].
Corollary 2.18.
There exists $C>0$ independent on $t,v$ so that if an approximation $(g^{A}_{t,v},w^{A}_{t,v})\in C^{2,\alpha}_{\beta,*}(g^{D}_{t})\times\tilde{\mathbf{O}}(g^{D}_{t})$ satisfies: $\|g^{A}_{t,v}-g^{D}_{t}\|_{C^{2,\alpha}_{\beta,*}(g^{D}_{t})}<\frac{1}{C},$ and $g^{A}_{t,v}-g^{D}_{t}+v\perp_{L^{2}(g^{D}_{t})}\tilde{\mathbf{O}}(g^{D}_{t})$, then, one has:
$\|g^{A}_{t,v}-\hat{g}_{t,v}\|_{C^{2,\alpha}_{\beta,*}}+\|w^{A}_{t,v}-\hat{w}_{t,v}\|_{L^{2}(g^{D}_{t})}<C\big\|\mathbf{\Phi}_{g^{D}_{t}}(g^{A}_{t,v})+\pi_{g^{A}_{t,v}}w^{A}_{t,v}\big\|_{r_{D}^{-2}C^{\alpha}_{\beta}(g^{D}_{t})}.$ (2.12)
Remark 2.19.
Applied to $(g^{A}_{t,0},w^{A})=(g^{D}_{t},0)$, (2.12) implies that $\|g^{D}_{t}-\hat{g}_{t,0}\|_{C^{2,\alpha}_{\beta,*}}\leqslant Ct_{\max}^{\frac{2-\beta}{4}}$.
2.4.2. Einstein metric Gromov-Hausdorff close to an Einstein orbifold
Theorem 2.15 is not merely a gluing-perturbation theorem: the tedious norms involved are not chosen for their convenient analytical properties but from the Gromov-Hausdorff convergence of Einstein metrics. Any Einstein metric sufficiently close to an orbifold is reached by Theorem 2.15.
Theorem 2.20 ([Ozu22a, Ozu22b]).
Let $(M_{o},g_{o})$ be an Einstein orbifold. There exists $\delta>0$ such that if $(M,g)$ is an Einstein $4$-orbifold satisfying
$d_{GH}\big((M,g),(M_{o},g_{o})\big)<\delta,$
Then, there exists a naïve gluing $g^{D}_{t}$ and $v\in\tilde{\mathbf{O}}(g^{D}_{t})$ such that:
$\hat{g}_{t,v}=\phi^{*}g\qquad\text{and}\qquad\hat{w}_{t,v}=0.$
3. Gromov-Hausdorff isolation of spherical orbifolds
Let us assume that an Einstein metric $(M,g)$ is sufficiently close to an Einstein orbifold $(M_{o},g_{o})$. Then, with $\phi:M\to M$ from Theorem 2.20, since $(\phi^{*}g,0)$ is the only solution of the equation, one has by (2.12):
$\|g^{A}_{t,v}-\phi^{*}g\|_{C^{2,\alpha}_{\beta,*}}+\|w^{A}\|_{L^{2}(g^{D})}<C\big\|\mathbf{\Phi}_{g^{D}_{t}}(g^{A})+\pi_{g^{A}_{t,v}}w^{A}\big\|_{r_{D}^{-2}C^{\alpha}_{\beta}(g^{D}_{t})}.$ (3.1)
In particular, if $\|w^{A}\|_{L^{2}(g^{D})}$ is much larger than $C\big\|\mathbf{\Phi}_{g^{D}_{t}}(g^{A})+\pi_{g^{A}}w^{A}\big\|_{r_{D}^{-2}C^{\alpha}_{\beta}(g^{D}_{t})}$, we reach a contradiction as used in [Ozu22b, Section 5.3]. Our strategy is therefore to construct such an approximation $(g^{A},w^{A})=(g^{A}_{t,v},w^{A}_{t,v})$ and to estimate $w^{A}$ carefully. Constructing this approximation will be done following these steps:
  1. (1)
    For each singular point $p_{o}$ of $M_{o}$, consider the bubble tree composed of the Ricci-flat ALE metrics collapsing at $p_{o}$ and $g^{B}=g^{B}_{t}$ the glued metric so that $g^{B}=g^{D}_{t}$ as in Definition 2.8.
  2. (2)
    We start by perturbing $(g^{B},0)$ to a Ricci-flat modulo obstruction metric $(\bar{g}_{B},u_{B})$ and show that ${\rm Ric}(\bar{g}_{B})$ decays fast at infinity. This is the content of Proposition 3.1.
  3. (3)
    We then show in Corollary 3.2 that either $u_{B}$ is orthogonal to a specific rescaling obstruction $\mathbf{o}_{1}$ or it is much larger than its projection on $\mathbf{o}_{1}$.
  4. (4)
    In Section 3.2, we then construct our approximate metric $g^{A}$ by gluing the various Ricci-flat modulo obstructions gluings of bubble trees, $\bar{g}_{B}$, to a hyperbolic or spherical orbifold $(M_{o},g_{o})$ while keeping track of additional obstructions $u_{o}\approx\lambda_{1}\mathbf{o}_{1}$ with $\lambda_{1}\neq 0$ this gluing induces.
  5. (5)
    On the bubble tree region, the resulting approximate obstruction is $w^{A}=u_{B}+\lambda_{1}\mathbf{o}_{1}$. The end of the proof of Theorem 1.1 is in Section 3.3. The argument is: if the projection of $w^{A}=u_{B}+\lambda_{1}\mathbf{o}_{1}$ onto $\mathbf{o}_{1}$ were small enough not to contradict (3.1), then $u_{B}$ would in fact be much larger than $\lambda_{1}\mathbf{o}_{1}$ and the projection of $w^{A}$ on the orthogonal of $\mathbf{o}_{1}$ would contradict (3.1).
3.1. Gluing bubble trees and controls
In our context, $(M,g^{D}_{t})$ is composed of the core orbifold $(M_{o},g_{o})$ and of a bubble tree at each singular point. Let us pick one such singular point $p_{o}\in M_{o}$ and denote $g^{B}_{t}$ the associated bubble tree metric. The Ricci-flat ALE metrics in this tree are $(N_{j},g_{b_{j}})_{j\in J_{p_{o}}}$. We similarly keep track of $v^{B}\in\tilde{\mathbf{O}}(g^{B}_{t})$, the piece of $v\in\tilde{\mathbf{O}}(g^{D}_{t})$, supported in the bubble tree associated with $g^{B}_{t}$.
We first extend [Ozu22b, Theorem 4.6] and show that $g^{B}_{t}$ can be perturbed to a Ricci-flat modulo obstruction metric in suitable spaces. There is no need to project the divergence orthogonally to Killing vector fields. Indeed, since there are no Killing vector fields growing slower than linearly at infinities of Ricci-flat ALE metrics, the operator ${\delta}_{g^{B}_{t}}\delta_{g^{B}_{t}}^{*}$ (without tilda) is invertible between the relevant spaces.
Proposition 3.1.
For any $0<\alpha<1$, $0<\beta<2$ and $\varepsilon>0$, there exist $C>0$ and $\delta>0$ such that for any $t_{\max},\|v\|_{C^{2,\alpha}_{\beta,*}}<\delta$, there is a unique solution $(\bar{g}^{B}_{t,v},u^{B})$ to
  • $\|g^{B}_{t}-\bar{g}^{B}_{t,v}\|_{C^{2,\alpha}_{\beta,*}(g^{D}_{t})}<\varepsilon,$
  • $g^{B}_{t}+v^{B}-\bar{g}^{B}_{t,v^{B}}\perp_{L^{2}(g^{D}_{t})}\tilde{\mathbf{O}}(g^{\textbf{D}}_{t})$,
  • $\mathbf{\Phi}_{g^{B}_{t}}(\bar{g}^{B}_{t,v})+\pi_{\bar{g}^{B}_{t,v}}u^{B}=0$, or equivalently: $\mathrm{E}(\bar{g}^{B}_{t,v})+\pi_{\bar{g}^{B}_{t,v}}u^{B}=0,\text{ and}$ $\delta_{g^{B}_{t}}(\bar{g}^{B}_{t,v})=0.$
where we denote $\tilde{\mathbf{O}}(g^{B}_{t})$ the set of elements of $\tilde{\mathbf{O}}(g^{D}_{t})$ supported in the bubble tree region associated to $g^{B}_{t}$, and where $u^{B}\in\tilde{\mathbf{O}}(g^{B}_{t})$. We have the following estimates:
$\|g^{B}_{t}-\bar{g}^{B}_{t,v}\|_{C^{2,\alpha}_{\beta,*}(g^{D}_{t})}\leqslant C(t^{\frac{2-\beta}{4}}+\|v^{B}\|_{C^{2,\alpha}_{\beta,*}}^{2}),$ (3.2)
as well as $\mathrm{E}(\bar{g}_{B})=\mathcal{O}(r^{-6})=o(r^{-4})$ at infinity, hence in particular,
$\pi_{\bar{g}_{B}}u_{B}=\mathcal{O}(r^{-6})=o(r^{-4}).$ (3.3)
Proof.
For ease of notation, we will denote $g^{B}=g^{B}_{t}$, $\bar{g}_{B}=\bar{g}^{B}_{t,v}$. The proof is an application of the implicit function theorem between Banach spaces. For each element $v\in\tilde{\mathbf{O}}(g^{B})$ (in particular $v^{B}$), one searches for $(g(v),u(v))$ with $g(v)=g^{B}+v+h(v)$, such that
$\mathbf{\Phi}_{g^{B}}(g(v))+\pi_{g(v)}u(v)=0$
where $(h(v),u(v))\in(C^{2,\alpha}_{\beta,**}\cap\tilde{\mathbf{O}}(g^{B})^{\perp}\cap\ker{\delta_{g^{B}}})\times\tilde{\mathbf{O}}(g^{B})$ where $C^{2,\alpha}_{\beta,**}$ is defined in Section III.
The implicit function theorem is applied to the set of equations parametrized by $v$:
$\Psi_{g^{B}}(v,h,u)=\mathbf{\Phi}_{g^{B}}(g^{B}+v+h)+\pi_{g^{B}+v+h}u.$
$\Psi_{g^{B}}$ is a smooth map in its arguments between the Banach spaces:
$\left(\tilde{\mathbf{O}}(g^{B})\cap C^{2,\alpha}_{\beta,**}\right)\times\left(C^{2,\alpha}_{\beta,**}\cap\tilde{\mathbf{O}}(g^{B})^{\perp}\cap\ker{\delta_{g^{B}}}\right)\times\left(\tilde{\mathbf{O}}(g^{B})\cap r_{B}^{-2}C^{\alpha}_{\beta}\right)\to(1+r_{B})^{-4}C^{\alpha}_{\beta}.$
First note that as in the proof of Lemma 2.14, one has for some $C>0$ independent of $t$,
$\|\Psi_{g^{B}}(0,0,0)\|_{r_{D}^{-2}C^{\alpha}_{\beta}}\leqslant Ct_{\max}^{\frac{2-\beta}{4}}.$ (3.4)
Denote by $\pi_{\tilde{\mathbf{O}}(g^{B})^{\perp}}$ the $L^{2}(g^{B})$-orthogonal projection on $\tilde{\mathbf{O}}(g^{B})^{\perp}$. We know from Lemma A.10 that the linearization of $\pi_{\tilde{\mathbf{O}}(g^{B})^{\perp}}\Psi_{g^{B}}$ at $(0,0,0)$, is invertible in the direction $\{0\}\times(C^{2,\alpha}_{\beta,**}\cap\tilde{\mathbf{O}}(g^{B})^{\perp}\cap\ker{\delta_{g^{B}}})\times\{0\}$ with an inverse whose norm is independent of $t$. Consequently, the linearization of $\Psi_{g^{B}}$ is invertible in the last two components and the implicit function theorem applies. The resulting $(h(v),u(v))$ is controlled as:
$\|h(v)\|_{C^{2,\alpha}_{\beta,**}}+\|u(v)\|_{r_{B}^{-2}C^{\alpha}_{\beta}}\leqslant C\Big(t^{\frac{2-\beta}{4}}+\|v\|_{C^{2,\alpha}_{\beta,*}}^{2}\Big).$
By Remark A.7, $h(v)$ has no $r^{-2}$ term in its expansion and $\nabla^{l}h(v)=\mathcal{O}(r^{-4-l})$. Consequently, we find
$\mathrm{E}\big(\bar{g}^{B}_{t,v^{B}}\big)=\mathrm{E}\left(g^{B}_{t,v^{B}}+v^{B}+h(v^{B})\right)=\mathcal{O}(r^{-6}).$
At infinity, we assume that $g_{B}$ is in a set of coordinates where
$g_{B}=g_{e}+H^{4}+\mathcal{O}(r_{b}^{-5}),$
where $H^{4}$ is a homogeneous harmonic symmetric $2$-tensor with $|H^{4}|=\mathcal{O}(r^{-4})$ with $\mathrm{tr}_{e}H^{4}=0$, $\delta_{e}H^{4}=0$ and $H^{4}(\partial_{r_{b}},\partial_{r_{b}})=0$. The existence of such coordinates is proven in [DO24, Lemma 11.12]
Let us start by decomposing $\tilde{\mathbf{O}}(g^{B})$ the refined space of obstructions from Section 2.3.2. Denote $(N,g_{b})=(N_{j_{o}},g_{b_{j_{o}}})$, where $j_{o}=j_{p^{-1}(p_{o})}$, the core orbifold of the bubble tree as in Definition 2.8.
An infinitesimal Ricci-flat deformation of $g_{b}$, $\mathbf{o}_{1}\in\mathbf{O}(g_{b})$, plays a distinct role:
$\mathbf{o}_{1}:=\mathrm{Hess}_{g_{b}}u-\frac{1}{4}(\Delta_{g_{b}}u)g_{b},$
where $u$ is the unique solution of $\Delta_{g_{b}}u=8$, $u=r_{b}^{2}+o(1)$, see [BH23, Ozu20]. On $(N,g_{b})$, we decompose $L^{2}(g_{b})$-orthogonally
[…]
and denote $(\tilde{\mathbf{o}}_{k})_{k\geqslant 2}$ an $L^{2}(g^{B})$-orthogonal basis of $\tilde{\mathbf{O}}_{0}(g^{B})$. Then, there exist numbers $(\mu_{k})_{k}$ such that
$u_{B}=\mu_{1}\tilde{\mathbf{o}}_{1}+\sum_{k\geqslant 2}\mu_{k}\tilde{\mathbf{o}}_{k}.$ (3.5)
Corollary 3.2.
We have the estimate:
$\mu_{1}\|\tilde{\mathbf{o}}_{1}\|_{L^{2}(g_{b})}+\sum_{k\geqslant 2}\varepsilon_{k}(t,v^{B})|\mu_{k}|=0,$ (3.6)
where $\varepsilon_{k}(t,v^{B})\to 0$ as $(t,v^{B})\to 0$.
Remark 3.3.
This corollary says that if $\mu_{1}\neq 0$, then at least one of the other $\mu_{k}$ is much larger than $\mu_{1}$. This is an extension to trees of singularities of [Ozu21b, Proposition 1.12], which relied on the real-analyticity of the moduli space of Ricci-flat modulo obstructions metrics close to a single smooth Ricci-flat ALE metric. There is no such result for trees of Ricci-flat ALE metrics.
Proof.
The proof relies on an application of [DO24, Lemma 11.16] together with (3.5) which says that if $\bar{g}^{B}$ is asymptotic to $\mathbb{R}^{4}/\Gamma$, then $\tilde{\mathbf{o}}_{1}$ is the only component of $u_{B}$ inducing some $\frac{dr_{b}^{2}}{r_{b}^{4}}$ at infinity:
$u_{B}(\partial_{r_{b}},\partial_{r_{b}})$ $=\mu_{1}\tilde{\mathbf{o}}_{1}(\partial_{r_{b}},\partial_{r_{b}})+\sum_{k\geqslant 2}\mu_{k}\tilde{\mathbf{o}}_{k}(\partial_{r_{b}},\partial_{r_{b}})$ (3.7)
$=C\mu_{1}\frac{|\Gamma|}{|\mathbb{S}^{3}|}\|\tilde{\mathbf{o}}_{1}\|^{2}_{L^{2}(g_{b})}r_{b}^{-4}+o(r_{b}^{-4}).$
If we had the same expansion for $\pi_{\bar{g}_{B}}u_{B}(\partial_{r_{b}},\partial_{r_{b}})$, we would conclude instantly from (3.3) that $\mu_{1}=0$. The rest of the proof consists in controlling the $O(r_{b}^{-4})$ terms of $\pi_{\bar{g}_{B}}u_{B}-u_{B}$.
From (3.5), we have $\pi_{\bar{g}_{B}}u_{B}=\mu_{1}\pi_{\bar{g}_{B}}\tilde{\mathbf{o}}_{1}+\sum_{k\geqslant 2}\mu_{k}\pi_{\bar{g}_{B}}\tilde{\mathbf{o}}_{k}$. Let us control $\pi_{\bar{g}_{B}}\tilde{\mathbf{o}}_{k}-\tilde{\mathbf{o}}_{k}=-\delta^{*}_{\bar{g}_{B}}(\delta_{\bar{g}_{B}}\delta^{*}_{\bar{g}_{B}})^{-1}\delta_{\bar{g}_{B}}\tilde{\mathbf{o}}_{k}$.
Each of the $\tilde{\mathbf{o}}_{k}$ is an element of $C^{k,\alpha}_{\beta,**}$ and $\delta_{\bar{g}_{B}}\tilde{\mathbf{o}}_{k}\in(1+r_{B})^{-4}r_{B}^{-1}C^{1,\alpha}_{\beta}$ since it vanishes in a neighborhood of infinity. Consequently, by Lemma A.11, $(\delta_{\bar{g}_{B}}\delta^{*}_{\bar{g}_{B}})^{-1}\delta_{\bar{g}_{B}}\tilde{\mathbf{o}}_{k}\in r_{B}C^{3,\alpha}_{\beta,**}$, and in particular, $(\delta_{\bar{g}_{B}}\delta^{*}_{\bar{g}_{B}})^{-1}\delta_{\bar{g}_{B}}\tilde{\mathbf{o}}_{k}=-Y^{3}+o(r^{-3})$ at infinity for some homogeneous $Y^{3}\in\ker\delta_{e}\delta^{*}_{e}$ with $|Y^{3}|_{e}\propto r^{-3}$. This is the standard behavior of elliptic operators between weighted Hölder spaces as one crosses an exceptional value. We deduce that
$\pi_{\bar{g}_{B}}\tilde{\mathbf{o}}_{k}-\tilde{\mathbf{o}}_{k}=\delta_{e}^{*}Y^{3}+o(r^{-4})$ (3.8)
for some harmonic vector field $Y^{3}$ in $r^{-3}$ at infinity. The first thing to note is that $\pi_{g}$ is exactly the $L^{2}(g^{B})$-projection of $\tilde{\mathbf{o}}_{k}$ onto $\ker\delta_{g^{B}}$, so we deduce that
$\|\pi_{\bar{g}_{B}}\tilde{\mathbf{o}}_{k}-\tilde{\mathbf{o}}_{k}\|_{L^{2}(g^{B})}\leqslant C\Big((\max_{j\in J_{p_{o}}}t_{j})^{\frac{2-\beta}{4}}+\|v^{B}\|_{C^{2,\alpha}_{\beta,*}}^{2}\Big)\|\tilde{\mathbf{o}}_{k}\|_{L^{2}(g^{B})}.$
There remains to determine how large $Y^{3}$ is compared to the $L^{2}$-norm of $\tilde{\mathbf{o}}_{k}$. By definition of $\pi_{\bar{g}^{B}}$, the term $Y^{3}$ is the asymptotic term of a solution $Y=Y^{3}+o(r^{-3})$ of the solution of
$\delta_{\bar{g}^{B}}\delta_{\bar{g}^{B}}^{*}Y=\delta_{\bar{g}^{B}}(\tilde{\mathbf{o}}_{k}).$ (3.9)
We use Proposition B.2 to control it. If in the bases of Section B one has the decomposition $Y^{3}=\sum_{i}\alpha_{i}X_{a_{1}^{-},i}+\beta_{i}X_{c_{2}^{-},i}+\gamma_{i}X_{b_{0}^{-},i}$, consider $X_{1}=\sum_{i}\alpha_{i}X_{a_{1}^{+},i}+\beta_{i}X_{b_{2}^{+},i}+\gamma_{i}X_{c_{0}^{+},i}$ the dual harmonic linear vector field. Then for the solution $\mathcal{X}_{1}$ of Lemma B.1 of $\delta_{\bar{g}^{B}}\delta_{\bar{g}^{B}}^{*}\mathcal{X}_{1}=0$ asymptotic to the linear vector field $X_{1}$ at infinity, we have by (3.9) and (B.3)
$\left|\int_{M}\delta_{\bar{g}^{B}}(\tilde{\mathbf{o}}_{k})(\mathcal{X}_{1})\right|=\left|\int_{M}\delta_{\bar{g}^{B}}\delta_{\bar{g}^{B}}^{*}Y(\mathcal{X}_{1})\right|=\left|B(Y^{3},X_{1})\right|=\sum_{i}C_{\alpha}\alpha_{i}^{2}+C_{\beta}\beta_{i}^{2}+C_{\gamma}\gamma_{i}^{2}=:\|Y^{3}\|^{2}.$ (3.10)
Here, the right-hand-side is quadratic in $\|Y^{3}\|$ while the left-hand-side is linear in $X_{1}$ (hence in $Y^{3}$) and multiplied by the tensor $\delta_{\bar{g}^{B}}(\tilde{\mathbf{o}}_{k})$, which is small by (2.7).
More concretely, by (2.7) and (B.2), the left-hand side of (3.10) is controlled as follows (not the absence of squre in the denominator)
$\frac{1}{\|r_{e}^{3}Y^{3}\|_{L^{\infty}(e)}}\left|\int_{M}\delta_{\bar{g}^{B}}(\tilde{\mathbf{o}}_{k})(\mathcal{X}_{1})\right|\to 0\qquad\text{ as }t\to 0.$
Since on the finite-dimensional space of homogeneous harmonic vector fields decaying like $r^{3}$, one has $C^{-1}\|Y^{3}\|^{2}\leqslant\|r_{e}^{3}Y^{3}\|_{L^{\infty}(e)}\leqslant C\|Y^{3}\|^{2}$, we conclude that $\|r_{e}^{3}Y^{3}\|_{L^{\infty}(e)}\to 0$ as $t\to 0$.
In particular, $|\delta_{e}^{*}Y^{3}(\partial_{r_{B}},\partial_{r_{B}})|\leqslant\varepsilon(t,v^{B})r^{-4}$ where $\varepsilon(t,v^{B})\to 0$ as $(t,v^{B})\to 0$. By (3.8),
$\pi_{\bar{g}_{B}}\tilde{\mathbf{o}}_{k}(\partial_{r_{B}},\partial_{r_{B}})-\tilde{\mathbf{o}}_{k}(\partial_{r_{B}},\partial_{r_{B}})=\varepsilon(t,v^{B})r_{B}^{-4}+o(r_{B}^{-4}),$ (3.11)
Together with (3.7), this leads to the desired statement. ∎
Remark 3.4.
We cannot prevent the presence of new $r^{-4}$ terms since there may be terms $Y^{3}$ in $r^{-3}$ appearing in the vector field obtained by $(\delta_{\bar{g}_{B}}\delta_{\bar{g}_{B}}^{*})^{-1}\delta_{g^{B}}$ when gauge-fixing our tensors $\tilde{\mathbf{o}}_{k}$ thanks to $\pi_{\bar{g}_{B}}$. An example of such a tensor $Y^{3}$ is $\nabla_{e}(r^{-2})=-2r^{-3}\partial_{r}$ since $\frac{1}{2}\mathrm{Hess}_{e}(r^{-2})=\frac{3dr^{2}-r^{2}g_{\mathbb{S}^{3}}}{r^{4}}$ where $g_{e}=dr^{2}+r^{2}g_{\mathbb{S}^{3}}$.
3.2. Construction of the approximate metric and estimates
We now construct a good approximation $g^{A}=g^{A}_{t,v}$ of the Einstein metric $g$ close to $g^{D}_{t}$ in order to apply Corollary 2.18 and to contradict (3.1) when the orbifold is spherical or hyperbolic. We will perturb our trees of singularities $g^{B}$ to make them match the orbifold metric $g_{o}$ better, then estimate the errors.
Close to a singular point $p_{o}\in M_{o}$, there exist coordinates in which one has
$g_{o}=g_{e}+H_{2}+\mathcal{O}(r^{3}),$ (3.12)
where $H_{2}$ is a homogeneous quadratic tensor with $\delta_{e}H_{2}=0$ and $\bar{P}_{e}H_{2}+\lambda g_{e}=0$ if $\mathrm{E}(g_{o})+\lambda g_{o}=0$. Such coordinates are classically obtained by gauge-fixing normal coordinates.
Proposition 3.5.
There exists a symmetric $2$-tensor $h_{2}$ and $u_{o}\in\tilde{\mathbf{O}}(g^{B})$ such that:
[…]
$\left\|u_{o}^{[2]}-c\,|\Gamma|\,\Lambda\,\tilde{\mathbf{o}}_{1}\right\|_{r_{B}^{-2}C^{\alpha}_{\beta}}\to 0,\text{ as $(t,v^{B})\to 0$.}$ (3.14)
Proof.
Consider $\chi$ a cut-off function supported in a region where $\bar{g}^{B}$ has ALE coordinates, then $\bar{P}_{\bar{g}^{B}}(\chi H_{2})+\lambda\bar{g}^{B}\in r_{D}^{-2}C^{\alpha}_{\beta}$ for $\beta>0$, so there exists a couple of elements $(h^{\prime},u_{o}^{[2]})\in(C^{2,\alpha}_{\beta,*}\cap\tilde{\mathbf{O}}(g^{B})^{\perp})\times\tilde{\mathbf{O}}(g^{B})$ such that $\bar{P}_{\bar{g}^{B}}(\chi H_{2})+\lambda\bar{g}^{B}=\bar{P}_{\bar{g}^{B}}h^{\prime}+{\pi_{\bar{g}^{B}}}u_{o}^{[2]},$ and defining $h_{2}=\chi H_{2}-h^{\prime}$, we solve (3.13).
Let us now focus on the estimate (3.14). This relies on the following integration by parts: as in [Ozu22b, Proposition 5.9], one has
$\int_{N}\langle\bar{P}_{\bar{g}_{B}}(\chi H_{2}),\pi_{\bar{g}^{B}}\tilde{\mathbf{o}}_{k}\rangle_{\bar{g}_{B}}=\int_{N}\langle\chi H_{2},\bar{P}_{\bar{g}_{B}}(\pi_{\bar{g}^{B}}\tilde{\mathbf{o}}_{k})\rangle_{\bar{g}_{B}}+Q(H_{2},O_{k}^{4}),$ (3.15)
where $O_{k}^{4}$ is the homogeneous harmonic term such that $\pi_{\bar{g}^{B}}\tilde{\mathbf{o}}_{k}=O^{4}_{k}+o(r^{-4})$ at infinity, and $Q(H_{2},O_{k}^{4})=\int_{\mathbb{S}^{3}/\Gamma}(3\langle H_{2},O_{k}^{4}\rangle+O_{k}^{4}(\nabla_{e}\mathrm{tr}_{e}H_{2},\partial_{r}))dv_{\mathbb{S}^{3}/\Gamma}$, where $\bar{g}^{B}$ is asymptotic to $\mathbb{R}^{4}/\Gamma$ at infinity.
Now a combination of [DO24, Proposition 11.18] and [Ozu22b, (86)] shows that there exists a constant $c\neq 0$ such that if $H_{2}$ comes from a spherical or hyperbolic orbifold with ${\rm Ric}(g_{o})=\Lambda g_{o}$ (or equivalently $\mathrm{E}(g_{o})+\Lambda g_{o}=0$ in dimension 44), then one has
$Q(H_{2},O_{k}^{4})=-2c\Lambda\,O_{k}^{4}(\partial_{r},\partial_{r}).$
Now, from (3.11) and (3.7), we know that if $k\neq 1$, then $r^{4}O_{k}^{4}(\partial_{r},\partial_{r})\to 0$ as $r\to+\infty$, while $r^{4}O_{1}^{4}(\partial_{r},\partial_{r})\to\frac{|\Gamma|}{|\mathbb{S}^{3}|}\|\mathbf{o}_{1}\|^{2}_{L^{2}(g_{b})}$ as $r\to+\infty$. ∎
We can further expand the metrics $g_{o}$ and $\bar{g}^{B}$ in divergence-free coordinates respectively at $p_{o}$ and at infinity to define the homogeneous tensors $H_{3}$, $H_{4}$ and $H^{4}$ on Euclidean space,
$g_{o}=g_{e}+H_{2}+H_{3}+H_{4}+\mathcal{O}(r_{e}^{5}),$ (3.16)
$\bar{g}^{B}=g_{e}+H^{4}+\mathcal{O}(r_{e}^{-5}).$
these tensors are homogeneous, $|H_{k}|_{e}\propto r_{e}^{k}$ and $|H^{4}|_{e}\propto r_{e}^{-4}$, and satisfy the equations:
$\bar{P}_{e}H_{2}+\lambda g_{e}=0,$ (3.17)
$\bar{P}_{e}H_{3}=0,$
$\bar{P}_{e}H_{4}+Q_{e}(H_{2},H_{2})+\lambda H_{2}=0,\text{ and}$
$\bar{P}_{e}H^{4}=0,$
where $Q_{g}$ denotes the quadratic terms in the expansion
$h\mapsto\mathbf{\Phi}_{g}(g+h)=\bar{P}_{g}(h)+Q_{g}(h,h)+R_{g}(h),$
where $R_{g}(h)$ contains the remaining cubic and higher order terms, see [Biq13, Section 14]. The equations (3.17) simply correspond to the expansions of ${\rm Ric}(g_{o})=\Lambda g_{o}$ at $p_{o}$ and ${\rm Ric}(\bar{g}^{B})=0$ at infinity.
The following result is a simplified version of [Ozu21a, Sections 2.1.1 and 2.1.2] and [DO24, Proposition 11.10], and follows [Biq13, Section 14]. Its proof is very similar to that of Proposition 3.5.
Proposition 3.6.
There exist tensors $h_{3}$, $h_{4}$ and $h^{4}$ satisfying the following properties:
[…]
Remark 3.7.
The above tensors are introduced for two reasons:
  1. (1)
    to make the bubble metric and the orbifold metric match to a higher order in the neck region where we interpolate between them, this is why they match the asymptotic expansions of the metrics, and
  2. (2)
    not to perturb the Einstein equation too much, this is why we have the tensors solve the above equations, see the resulting controls in Corollary 3.8.
Let us number $p_{o}^{1},\dots,p_{o}^{N}\in S_{o}$ the singular points of $g_{o}$ to be desingularized and denote $(\bar{g}^{B}_{1},u^{B}_{1}),\dots,(\bar{g}^{B}_{N},u^{B}_{N})$ the Ricci-flat modulo obstructions metrics of Proposition 3.1 obtained from the associated perturbed trees of singularities. Up to renumbering the gluing scales $t_{1}=t_{p^{-1}(p_{o}^{1})},\,\dots,\,t_{N}=t_{p^{-1}(p_{o}^{N})}$ to $g_{o}$, we assume
$t_{1}\leqslant\dots\leqslant t_{N}.$
We also denote $(h_{2,1},u_{o,1}),\dots,(h_{2,N},u_{o,N})$ and $(h_{3,1},u_{o,1}^{[3]},h_{4,1},u_{o,1}^{[4]},h^{4}_{1}),\dots,(h_{3,N},u_{o,N}^{[3]},h_{4,N},u_{o,N}^{[4]},h^{4}_{N})$ the $2$-tensors constructed in (3.13) and Proposition 3.6. Consider a cut-off function $\chi_{o,j}$ supported on of $\{r_{o}<2\varepsilon_{0}\}\cap B(p_{j},1)$, equal to $1$ on $\{r_{o}<\varepsilon_{0}\}\cap B(p_{j},1)$ and with derivatives controlled as $|\nabla_{g_{o}}^{l}\chi_{o,j}|\leqslant C_{l}r_{o}^{-l}$ for $\varepsilon_{0}>0$ of Definition 2.1 and $C_{l}>0$.
We have the following estimates.
Corollary 3.8.
Define the metrics and obstructions perturbed to $4^{th}$ order
$g_{o}^{[4]}:=g_{o}+\sum_{j=1}^{N}t_{j}^{2}\,\chi_{o,j}h^{4}_{j},\quad\text{ and }$ (3.21)
$\bar{g}^{B,[4]}_{j}:=\bar{g}^{B}_{j}+t_{j}\,h_{2,j}+t_{j}^{\frac{3}{2}}\,h_{3,j}+t_{j}^{2}\,h_{4,j},\qquad\bar{w}^{B,[4]}_{j}:=u_{B}-t_{j}u_{o,j}^{[2]}-t_{j}^{\frac{3}{2}}u_{o,j}^{[3]}-t_{j}^{2}u_{o,j}^{[4]}.$ (3.22)
They satisfy the controls: for all $l\in\mathbb{N}$, there exists $C_{l}>0$ depending on the above tensors such that
$r_{B}^{2+l}\left|\nabla^{l}\left(\mathbf{\Phi}_{g^{B}_{j}}(\bar{g}^{B,[4]}_{j})+\pi_{g^{B}_{j}}\bar{w}^{B,[4]}_{j}\right)\right|_{g^{B}_{j}}\leqslant C_{l}\,\,t_{j}^{\frac{5}{2}}\,r_{B}^{5},$ (3.23)
$r_{o}^{2+l}\left|\nabla^{l}\left(\mathbf{\Phi}_{g_{o}}(g_{o}^{[4]})+\lambda g_{o}^{[4]}\right)\right|_{g_{o}}\leqslant C_{l}\left(t_{j}^{4}r_{o}^{-8}+t_{j}^{2}\,1_{\{\varepsilon_{0}<r_{o}<2\varepsilon_{0}\}}\right)\text{ on }B(p_{j},3\varepsilon_{0}).$ (3.24)
where $1_{A}$ is the indicator function of a set $A$.
Proof.
Let us drop the index $j$ for ease of notation. The expansion of $g\mapsto\mathbf{\Phi}_{\bar{g}^{B}}(g)$ close to $g=\bar{g}^{B}$ in the direction $h^{[4]}:=\bar{g}^{B,[4]}-\bar{g}^{B}=t\,h_{2}+t^{\frac{3}{2}}\,h_{3}+t^{2}h_{4}$ yields
$\mathbf{\Phi}_{\bar{g}^{B}}(\bar{g}^{B,[4]})=\mathbf{\Phi}_{\bar{g}^{B}}(\bar{g}^{B})+\bar{P}_{\bar{g}^{B}}(h^{[4]})+Q_{\bar{g}^{B}}(h^{[4]},\,h^{[4]})+R_{\bar{g}^{B}}(h^{[4]}),$
and using the definitions of the tensors in the worst terms come from $r_{B}^{2+l}|\nabla^{l}_{\bar{g}^{B}}Q_{\bar{g}^{B}}(t\,h_{2},t^{\frac{3}{2}}\,h_{3})|\leqslant t^{\frac{5}{2}}r_{B}^{5}$.
We similarly control $g_{o}^{[4]}$, where the worst terms come from: $r_{o}^{2+l}|\nabla^{l}_{g_{o}}Q_{g_{o}}(t^{2}\,h^{4}\,,\,t^{2}\,h^{4})|\leqslant t^{4}r_{o}^{-8}$ and the cut-off of $t^{2}h^{4}$ by $\chi_{o}$ which satisfies $\varepsilon_{0}^{l}|\nabla_{g_{o}}^{l}\chi_{o}|\leqslant C_{l}$ by definition. ∎
We define the approximate Einstein modulo obstruction $(g^{A},w^{A})$ iteratively as follows, using the notations of Definition 2.6.
Definition 3.9 (Approximate Einstein modulo obstruction metric).
We again assume that we ordered the scale so that $t_{1}\leqslant\dots\leqslant t_{N}$.
  1. (1)
    Define $(g^{A}_{0},w^{A}_{0})=(g_{o},0)$,
  2. (2)
    For $j\geqslant 0$, define (using the notation #\# of Definition 2.5) $(g^{A}_{j+1},w^{A}_{j+1})$ by
    $\left\{\begin{aligned} g^{A}_{j+1}&:=\left(g^{A}_{j}+t_{j+1}^{2}\,\chi_{o,j}\,h^{4}_{t_{j+1}}\right)\,\,\#_{p_{j+1},t_{j+1}}\,\,g^{B,[4]}_{j}\\ w^{A}_{j+1}&:=w^{A}_{j}\,\,\#_{p_{j+1},t_{j+1}}\,\,\left(u^{B}_{j+1}+u_{o}^{j+1}\right).\end{aligned}\right$. (3.25)
  3. (3)
    Denote $g^{A}=g^{A}_{N}$.
On $M_{o}^{16t}$, one has $g^{A}=g_{o}^{[4]}$.
We finally measure how good an approximation our metrics $g^{A}_{t,v}$ are.
Proposition 3.10.
For $0<\beta<1$, we have
$\left\|\mathbf{\Phi}_{g^{D}_{t}}(g^{A})+\pi_{g^{A}}w^{A}\right\|_{r_{D}^{-2}C^{\alpha}_{\beta}}\leqslant Ct_{N}^{\frac{5-\beta}{4}}=o(t_{N}).$ (3.26)
Remark 3.11.
The key point here are that the control is in $t_{N}$ rather than in $t_{\max}$. It does not see the relative scales deep in the trees of singularities, and this is the reason why we need to perturb these trees of Ricci-flat ALE metrics into a single Ricci-flat modulo obstruction metric in Section 3.1.
Proof.
Thanks to Corollary 3.8, there only remains to control the transition regions in order to prove (3.26). In such a region with $t_{j}^{\frac{1}{4}}<r_{D}<2t_{j}^{\frac{1}{4}}$, $g^{A}$ interpolates between the following metrics satisfying for all $\varepsilon>0$,
$\Phi^{*}_{j}g_{o}^{[4]}=g_{e}+H_{2}+H_{3}+H_{4}+t_{j}^{2}H^{4}+\mathcal{O}(t_{j}^{2}r_{o}^{-2-\varepsilon}+r_{o}^{5})\qquad\text{ and }$ (3.27)
$t_{j}\phi_{t_{j}^{-1/2}}\Psi^{*}_{j}(\bar{g}^{B,[4]})=g_{e}+t_{j}^{2}H^{4}+H_{2}+H_{3}+H_{4}+\mathcal{O}(t_{j}^{2}r_{o}^{-2}+t_{j}^{\frac{5}{2}}r_{o}^{-5}).$ (3.28)
In particular, the difference between the metrics, $\mathcal{O}(t_{j}^{2}r_{o}^{-2+\varepsilon}+t^{\frac{5}{2}}r_{o}^{-5}+t_{j}^{2}r_{o}^{-2-\varepsilon}+r_{o}^{5})=\mathcal{O}(t_{j}^{\frac{5}{4}})$ in the region $t_{j}^{\frac{1}{4}}<r_{D}<2t_{j}^{\frac{1}{4}}$ for small enough $\varepsilon>0$. Together with the analogous controls of the derivative, and the controls of the cut-off functions, this yield the stated result. ∎
3.3. Isolation of hyperbolic and spherical orbifolds
We finally show that there cannot be any other Einstein orbifold $GH$-close to a spherical or hyperbolic orbifold. The argument is by contradiction and consists in proving that $w^{A}$ cannot be small, contradicting (3.1).
Proof of Theorem 1.1.
Let $(M_{o},g_{o})$ be a spherical or hyperbolic orbifold, and assume that, a sequence of Einstein orbifolds $(M_{i},g_{i})_{i\in\mathbb{N}}$ satisfies $d_{GH}\big((M_{i},g_{i}),(M_{o},g_{o})\big)\to 0.$ Since $(M_{o},g_{o})$ is a spherical or hyperbolic orbifold, we have $\mathbf{O}(g_{o})=\{0\}$, and it is isolated as an Einstein orbifold on $M_{o}$. Hence, up to taking a subsequence, we assume $M_{i}=M\neq M_{o}$, and there is indeed a nontrivial bubble-tree degeneration.
By Theorem 2.20, for all $i$, $(g_{i},0)$ is isometric to an Einstein (modulo obstruction) perturbation $\hat{g}_{t_{i},v_{i}}$ of a naïve gluing $g^{D}_{t_{i}}$ with $v_{i}\in\tilde{\mathbf{O}}(g^{D}_{t_{i}})$ the metric and elements of the approximate kernel defined in Section 2.4. Up to taking yet another subsequence, we can order the $N$ singular points of $M_{o}$ as $p_{1},...,p_{N}$ such that
$0<t_{1,i}\leqslant t_{2,i}\leqslant...\leqslant t_{N,i},$
where $t_{k,i}$ is the scale of the gluing at $p_{k}$ in the construction of $g^{A}_{t_{i},v_{i}}$ in Definition 3.9.
Thanks to the control of Proposition 3.10 and (3.1), we deduce that for any $0<\beta<1$ we have $\|w^{A}\|_{L^{2}}\leqslant Ct_{N,i}^{\frac{5-\beta}{4}}=o(t_{N,i})$, hence denoting $\mu_{k}$, and $\tilde{\mathbf{o}}_{k}$ the elements of Section 3.1 for the last metric $\bar{g}^{B}_{t_{i},v_{i},N}$ glued in the construction of Definition 3.9, we have the following decomposition in the orthonormal basis $(\tilde{\mathbf{o}}_{k})_{k}$ of $\tilde{\mathbf{O}}(g^{D}_{t})$
$w^{A}|_{N_{N}^{16t}}$ $=u_{B_{N},i}-t_{N,i}u_{o,N}^{[2]}-t_{N,i}^{\frac{3}{2}}u_{o,N}^{[3]}-t_{N,i}^{2}u_{o,N}^{[4]}$
$=\big(\mu_{1}-t_{N,i}\lambda_{1}+o(t_{N,i})\big)\tilde{\mathbf{o}}_{1}+\sum_{k}\big(\mu_{k}+o(t_{N,i})\big)\tilde{\mathbf{o}}_{k}.$
Now, in order to have $\|w^{A}|_{N_{N}^{16t}}\|_{L^{2}}=o(t_{N,i})$, since $\|\tilde{\mathbf{o}}_{1}\|_{L^{2}(g^{D})}>0$ is bounded away from 0 independently on $(t_{i},v_{i})$, this imposes
$\mu_{1}=t_{N,i}\lambda_{1}+o(t_{N,i}),\text{ and}$ (3.29)
$\mu_{k}=o(t_{N,i}),\text{ for }k\geqslant 2$ (3.30)
The first obstruction (3.29) and (3.6) imply that
$\sum_{k\geqslant 0}\varepsilon_{k}(t_{i},v_{i}^{B})|\mu_{k}|=t_{N,i}\lambda_{1}+o(t_{N,i}).$
Since from (3.14), we have $\lambda_{1}\neq 0$ for hyperbolic and spherical orbifolds, this contradicts (3.30). ∎
Proof of Corollary 1.3.
The proofs of the gap theorems have a similar classical strategy by contradiction:
  1. (1)
    take a sequence of Einstein orbifolds $(M_{i},g_{i})$ with ${\rm Ric}(g_{i})=3g_{i}$, ${\rm Vol}(M_{i},g_{i})>v>0$ and assume one of the conditions (1)-(5) with $\varepsilon_{0}$ replaced by $\frac{1}{i}$,
  2. (2)
    take a sublimit $(M_{o},g_{o})$ by compactness, and show that it must be a spherical orbifold, and
  3. (3)
    conclude by Theorem 1.1 that for large enough $i$, all of the $(M_{i},g_{i})$ are isometric to $(M_{o},g_{o})$.
Part II Regularity of Einstein $5$-manifolds
In this part, we now use our $4$-dimensional isolation result to prove new regularity properties for $5$-manifolds with bounded Ricci curvature.
4. Uniqueness of tangent cones and the refined rectifiability
In this section, we consider the following setting: Let $(M^{5}_{i},g_{i},p_{i})\xrightarrow{pGH}(X^{5},d,p)$ be a sequence of noncollapsed pointed manifolds with uniformly bounded Ricci curvature
$|{\rm Ric}_{i}|\leqslant 4,\qquad\text{ and }\qquad{\rm Vol}(B_{1}(p_{i})>v>0.$
Our goal is to analyze the singular structure of the noncollapsed limit $(X^{5},d,p)$, in particular to establish the uniqueness of tangent cones along the top stratum $\mathcal{S}^{1}\setminus\mathcal{S}^{0}$ and to refine the rectifiability of the singular set $\mathcal{S}(X)$. We also prove regularity results for the asymptotic geometry of Ricci-flat manifolds with Euclidean volume growth.
4.1. Preliminaries
We recall the standard stratification framework for $n$-dimensional noncollapsed Ricci-bounded limit spaces and several quantitative tools that will be used throughout the paper.
Let $(M^{n}_{i},g_{i},p_{i})\xrightarrow{pGH}(X^{n},d,p)$ be a sequence of noncollapsed pointed manifolds with uniformly bounded Ricci curvature $|{\rm Ric}_{i}|\leqslant n-1,$ and ${\rm Vol}(B_{1}(p_{i}))>v>0$. We call such $X$ as a Ricci-bounded limit space.
Definition 4.1.
A metric space $Y$ is called $k$-symmetric at $y$ with respect to $\mathcal{L}^{k}$ if there exists some pointed isometry $\iota:(\mathds{R}^{k}\times C(Z),(0,z))\to(Y,y)$ for some metric cone $C(Z)$ over a compact metric space $Z$ with vertex $z$ and $\mathcal{L}^{k}=\iota(\mathds{R}^{k}\times\{z\})$.
Following [CC97], we introduce the stratification and define the $k^{th}$-stratum to be
$\mathcal{S}^{k}\equiv\{x\in X~|~\text{ no tangent cone at }x\text{ is }(k+1)\text{-symmetric}.\}$
Remark 4.2.
We will later refine the stratification in dimension $n=5$ using the uniqueness of the splitting tangent cones; see (4.7).
The regular set $\mathcal{R}(X)$ consists of points whose tangent cones are uniquely $\mathbb{R}^{n}$, and the singular set is $\mathcal{S}(X)=X\setminus\mathcal{R}(X)$. Combining the pioneering work [Col97, CC97] and the resolution of codimension four conjecture by Cheeger-Naber [CN15], we have
Theorem 4.3 ([CN15]).
The singular set $\mathcal{S}(X^{n})$ satisfies $\mathcal{S}=\mathcal{S}^{n-4}$ and $\dim(\mathcal{S})\leqslant n-4$.
In [CJN21] and [JN21], they proved a structure theorem for the singular set $\mathcal{S}$:
Theorem 4.4.
[CJN21, JN21] Let $X^{n}$ be a noncollapsed Ricci-bouned limit space. Then:
  1. (1)
    For $\mathcal{H}^{n-4}$-a.e. $x\in\mathcal{S}(X)$, the tangent cone at $x$ is unique and isometric to the cone $\mathds{R}^{n-4}\times C(\mathbb{S}^{3}/\Gamma)$.
  2. (2)
    The singular set $\mathcal{S}$ is (n4)(n-4)-rectifiable: there exists a countable collection of $\mathcal{H}^{n-4}$-measurable subsets $Z_{i}\subset Z$ and bi-Lipschitz maps $\phi_{i}:Z_{i}\to\mathds{R}^{n-4}$ such that $\mathcal{H}^{n-4}(\mathcal{S}\setminus\bigcup_{i}Z_{i})=0$.
Both the uniqueness of tangent cones and the bi-Lipschitz structure hold for the top stratum $\mathcal{S}^{n-4}$ modulo an $\mathcal{H}^{n-4}$-measure zero subset. In dimension five we will strengthen this in the next two sections, by proving uniqueness and bi-Lipschitz structure on the entire top stratum, without discarding any measure-zero subset.
Next we recall the quantitative symmetry:
Definition 4.5.
Given a metric space $X$ and $r>0,\varepsilon>0$, we define a ball $B_{r}(x)\subset X$ as $(k,\varepsilon)$-symmetric with respect to $\mathcal{L}_{x,r}$ if $d_{GH}(B_{r}(x),B_{r}(0,x^{\prime}))<\varepsilon r$ for some $k$-symmetric metric cone $\mathds{R}^{k}\times C(Z)$ with vertex $(0,x^{\prime})$, and that $\mathcal{L}_{x,r}=\iota(\mathds{R}^{k}\times\{x^{\prime}\}\cap B_{r}(0,x^{\prime}))$ where $\iota:B_{r}(x^{\prime})\to B_{r}(x)$ is the $\varepsilon r$-Gromov-Hausdorff map.
One of the fundamental tools to study the structure of noncollapsed spaces with uniform Ricci curvature lower bound is the Bishop-Gromov monotonicity: for any $(M^{n},g)$ with ${\rm Ric}_{M}\geq-\kappa g$, the volume ratio defined as follows is non-increasing in $r$:
$\mathcal{V}_{x}(r)=\mathcal{V}^{\kappa}_{x}(r)\equiv\frac{{\rm Vol}(B_{r}(x))}{{\rm Vol}_{-\kappa}(B_{r}(0))},$
where ${\rm Vol}_{-\kappa}(B_{r}(0))$ is the volume of the ball with radius $r$ in the space form with constant curvature $\kappa$. We define the volume density at $x$ as $\lim_{r\to 0}\mathcal{V}_{x}(r)$. When the volume ratio is constant, the space is isometric to the model space. In [CC96], they proved the almost rigidity:
Theorem 4.6.
For any $\varepsilon>0$ and $\delta\leqslant\delta(n,\varepsilon)$, the following holds. If $(M^{n},g)$ is a Riemannian manifold with ${\rm Ric}_{M}\geq-\delta g$, and $\mathcal{V}_{x}(2)\geq(1-\delta)\mathcal{V}_{x}(1)$, then $B_{2}(x)$ is $(0,\varepsilon)$-symmetric.
When multiple 0-symmetries occur, higher symmetry follows from the quantitative cone-splitting principle of [CN13].
Lemma 4.7.
Let $\varepsilon,\tau>0$ and $\delta\leqslant\delta(\varepsilon,\tau)$. Assume ${\rm Ric}_{M}\geq-\delta g$ and
  1. (1)
    $B_{2}(p)$ is $(k,\delta)$-symmetric with respect to $\mathcal{L}_{x,2}\subset B_{2}(p)$.
  2. (2)
    there exists some $x\in B_{1}(p)\setminus B_{\tau}(\mathcal{L}_{x,2})$ such that $B_{2}(x)$ is $(0,\delta)$-symmetric.
Then $B_{1}(p)$ is $(k+1,\varepsilon)$-symmetric.
Combining Lemma 4.7 with the codimension-four theorem 4.3 yields the $\varepsilon$-regularity theorem of [CN15].
Theorem 4.8 ([CN15]).
For all $n\in\mathbb{N},\,\,v>0$, there exists $\varepsilon_{0}(n,v)>0$ such that if ${\rm Vol}(B_{1}(p))>v>0$, $|{\rm Ric}_{M^{n}}|\leqslant n-1$ and $B_{2}(p)$ is $(n-3,\varepsilon)$-symmetric with $\varepsilon\leqslant\varepsilon_{0}$, then the harmonic radius $r_{h}(p)$ at $p$ is greater than 1. If $M^{n}$ is Einstein, then we have the curvature bound $\sup_{B_{1}(p)}|{\rm Rm}|\leqslant 1$.
According to the $\varepsilon$-regularity, if a ball has enough symmetry, then it is regular enough. Consult [And90] or [JN21] for the definition of harmonic radius.
Finally, we recall the notion of $\varepsilon$-splitting function, which is crucial to establish the bi-Lipschitz structure for the singular set.
Definition 4.9.
Let $B_{r}(p)\subset M$. We say $u:B_{r}(p)\to\mathds{R}$ is a $\varepsilon$-splitting function if the following holds:
  1. (1)
    $\Delta u=0$;
  2. (2)
    $\fint_{B_{r}}\big||\nabla u|^{2}-1\big|\leqslant\varepsilon$;
  3. (3)
    $\sup_{B_{r}(p)}|\nabla u|\leqslant 1+\varepsilon$;
  4. (4)
    $r^{2}\fint_{B_{r}(p)}|\nabla^{2}u|^{2}\leqslant\varepsilon$.
The existence of $\varepsilon$-splitting function is equivalent to the metric splitting behavior of the ball. See [CC96] and also [CJN21].
Theorem 4.10.
Let $(M^{n},g)$ be a smooth Riemannian manifold with ${\rm Ric}_{M^{n}}\geq-\delta$. For $\varepsilon>0$, if $\delta\leqslant\delta(n,\varepsilon)$, then we have the following
  1. (1)
    If $d_{GH}(B_{2}(p),B_{2}(0,x^{\prime}))\leqslant\delta$, with $(0,x^{\prime})\in\mathds{R}\times X$ for a metric space $X$, then there exists an $\varepsilon$-splitting function on $B_{1}(p)$;
  2. (2)
    If there exists a $\delta$-splitting function on $B_{2}(p)$, then $d_{GH}(B_{1}(p),B_{1}(0,x^{\prime}))\leqslant\varepsilon$ for $(0,x^{\prime})\in\mathds{R}\times X$. Moreover, $(u,v):B_{1}(p)\to\mathds{R}\times X$ gives an $\varepsilon$-Gromov-Hausdorff map for some $v:B_{1}(p)\to X$.
We also record the convergence of limiting harmonic functions. The result holds under more general assumptions. See [CJN21, JN21] and also [AH18, AHT18, Che99, Din02, GMS15, MN19, ZZ19].
Proposition 4.11.
Let $(M^{n}_{i},g_{i},p_{i},\mu_{i})\to(X,d,p,\mu)$ satisfy ${\rm Ric}_{M^{n}_{i}}\geq n-1$ and ${\rm Vol}(B_{1}(p_{i}))>v$ and $\mu_{i}=|{\rm Vol}(B_{1}(p_{i})|^{-1}{\rm Vol}$. Assume that $u_{i}:B_{r}(p_{i})\to\mathds{R}$ is harmonic with $\sup_{B_{r}(p_{i})}|\nabla u_{i}|\leqslant C$ for some $r>0$ and $C>0$. Then passing to a subsequence there exists some a harmonic function $u:B_{r}(p)\to\mathds{R}$ satisfying
  1. (1)
    $u_{i}\to u$ uniformly on compact subsets of $B_{r}$.
  2. (2)
[…]
  3. (3)
    $u_{i}\to u$ in $W^{1,p}$-sense for any $1<p<\infty$.
Using the $W^{1,2}$ convergence of splitting functions, we can also define the $\varepsilon$-splitting function on the limit space $(X,d,p)$. We say $u$ is a $\varepsilon$-splitting function on $B_{r}(p)\subset X$ if there exist $B_{r}(p_{i})\subset M_{i}\to B_{r}(p)$ and $\varepsilon_{i}$-splitting functions on $B_{r}(p_{i})$ converging uniformly to $u$ with $\varepsilon_{i}\to\varepsilon$. Therefore, Theorem 4.10 also holds for the limit space $(X,d,p)$.
4.2. Uniqueness of tangent cones
Now we focus on dimension $n=5$. Let $X^{5}$ be a noncollapsed Ricci-bounded limit space. Then the singular set $\mathcal{S}(X^{5})=\mathcal{S}^{1}$. We will prove the uniqueness of tangent cones at the points in $\mathcal{S}^{1}\setminus\mathcal{S}^{0}$ and improve the regularity estimates around those points. Similar results hold for tangent cones at infinity.
Proof of Theorem 1.9.
Suppose we have the noncollapsing convergence
$(M_{i},g_{i},p_{i})\xrightarrow{pGH}(C(Y),d_{C(Y)},y)\quad\text{ and }\quad(N_{i},\tilde{g}_{i},q_{i})\xrightarrow{pGH}(C(Z),d_{C(Z)},z).$
with both satisfying $|{\rm Ric}_{i}|\to 0$. Since we assume the cone $C(Z)$ splits off a line: $C(Z)\cong\mathds{R}\times C(Z^{\prime})$, by codimension four theorem 4.3, we conclude that $Z^{\prime}$ is indeed a smooth Einstein $3$-manifold and thus $Z^{\prime}\cong\mathbb{S}^{3}/\Gamma$ for some $\Gamma\subset O(4)$. This implies that $Z$ is a sine suspension over $\mathbb{S}^{3}/\Gamma$ and thus a spherical Einstein $4$-orbifold with isolated singularities. Also by codimension four theorem 4.3, we have $Y$ is an Einstein $4$-orbifold. Since we assume $d_{GH}(Y,Z)<\varepsilon_{0}$, by the gap theorem 1.1, we conclude that $Y\cong Z$ provided $\varepsilon_{0}$ is small enough. This completes the proof. ∎
As a consequence, we prove the uniqueness of the tangent cone that splits off a line. In fact, this is a direct consequence of Theorem  1.9 and the fact that the space of the cross sections of the tangent cones at some fixed point is connected.
Proof of Theorem 1.11 and Theorem 1.13.
We prove the local uniqueness theorem 1.11, and the proof for the uniqueness of tangent cones at infinity 1.13 proceeds verbatim. Pick any $x\in\mathcal{S}^{1}\setminus\mathcal{S}^{0}$. By definition there exists one tangent cone $C(Z_{0})$ at $x$ that splits off a line. Let $C(Z_{1})$ be a tangent cone under another rescaling sequence. Let $\mathcal{C}_{x}\equiv\{Y:C(Y)$ is a tangent cone at x}x\}. Since $\mathcal{C}_{x}$ is known to be connected, there exist finitely many compact metric spaces $\{Y_{i}\}_{i=0}^{N}$ such that
  1. (1)
    $Y_{0}\cong Z_{0}$ and $Y_{N}\cong Z_{1}$;
  2. (2)
    $d_{GH}(Y_{i},Y_{i+1})\leqslant\varepsilon_{0}/2$ for any $0\leqslant i\leqslant N-1$, where $\varepsilon_{0}$ is the gap in Theorem 1.9.
By iterating Theorem 1.9, we have $Y_{i}\cong Y_{i+1}$ for each $i$ and thus $Z_{0}\cong Z_{1}$. This proves the uniqueness of the tangent cones at $x$. The second result follows since the 0-stratum $\mathcal{S}^{0}$ is countable. ∎
Next we prove the curvature estimates around each point $x\in\mathcal{S}^{1}\setminus\mathcal{S}^{0}$.
Fix any $x\in\mathcal{S}^{1}\setminus\mathcal{S}^{0}$. By uniqueness Theorem 1.11, for each $r\leqslant r_{x}$ for some small $r_{x}$, we have $B_{r}(x)$ is $(1,\varepsilon)$-symmetric with respect to $\mathcal{L}_{r}=\mathcal{L}_{x,r}$, with $\varepsilon\downarrow 0$ as $r\downarrow 0$. Fix some $\tau>0$. By cone-splitting Lemma 4.7 and the $\varepsilon$-regularity Theorem 4.8, we conclude that if $\varepsilon\leqslant\varepsilon_{0}(v,\tau)$ is small enough, then there exists some constant $\rho\ll 1$ such that the harmonic radius at $y$ has a lower bound:
$r_{h}(y)\geq\rho r\text{ for any }y\in B_{r}(x)\setminus B_{\tau r}(\mathcal{L}_{r}).$ (4.1)
If we assume that the sequence $M_{i}$ is Einstein, then by $\varepsilon$-regularity there exists $c(r)\downarrow 0$ as $r\downarrow 0$ such that
$\sup_{B_{\rho r}(y)}|{\rm Rm}|\leqslant c(r)\cdot r^{-2}.$ (4.2)
Then we can define a wedge region $\mathcal{W}(x)$ around $x$ to be
$\mathcal{W}(x)\equiv\mathcal{W}_{\tau}(x)\equiv\bigcup_{0<r\leqslant r_{x}}\{y\in B_{r}(x)\setminus B_{\tau r}(\mathcal{L}_{r})\}.$ (4.3)
Proof of Corollary 1.12.
Pick a point $y\in\mathcal{W}(x)$, set $r\equiv 2d(x,y)$, and use (4.1) and (4.2). ∎
Remark 4.12.
We can also take union on $\tau$ to define the rounded wedged region $\hat{\mathcal{W}}(x)$ at $x$
$\hat{\mathcal{W}}(x)\equiv\bigcup_{i}B_{r_{i}}(x)\setminus B_{\tau_{i}r_{i}}(\mathcal{L}_{r_{i}}).$ (4.4)
Similar estimates hold on the rounded wedge region, replacing $d(x,y)$ by the distance from $y$ to the splitting axis $\mathcal{L}_{r}$.
We now consider the asymptotic setting. Let $(M,g,p)$ be a pointed 5-manifold with Euclidean volume growth, ${\rm Vol}(B_{R}(p))\geq vR^{5}$, and satisfy that $\sup_{B_{2R}(p)\setminus B_{R}(p)}|{\rm Ric}|=o(R^{-2})$. If one tangent cone at infinity splits off a line, then it is unique by Theorem 1.13.
Fix a point $p\in M$. Next we construct the wedge region around infinity $\mathcal{W}_{p}(\infty)$: for each $R\geq R_{p}$ for some large $R_{p}$, we have $B_{R}(p)$ is $(1,\varepsilon)$-symmetric with respect to $\mathcal{L}_{R}=\mathcal{L}_{p,R}$ with $\varepsilon\downarrow 0$ as $R\uparrow\infty$. Fix some $\tau>0$. Similarly the wedge region at infinity $\mathcal{W}(\infty)\equiv\mathcal{W}_{p,\tau}(\infty)$ is defined as
$\mathcal{W}(\infty)\equiv\bigcup_{R\geq R_{p}}\{y\in B_{R}(p)\setminus B_{\tau R}(\mathcal{L}_{R})\}.$ (4.5)
Proof of Corollary 1.14.
Fix some $p\in M$. For $y\in\mathcal{W}(\infty)$ constructed as above, we set $R=2d(p,y)$. Then the proof is completed by (4.1) and (4.2). ∎
Similarly, the rounded wedge region is defined as $\hat{\mathcal{W}}(\infty)\equiv\bigcup_{\tau>0}\mathcal{W}_{p,\tau}(\infty)$, where similar estimates hold.
Finally we provide a proof of the higher dimensional Theorem 1.23, similar to $5$-dimensional case.
Proof of Theorem 1.23.
Similar to the proof of 1.9, by assumption, every tangent cone at $x$ can be written as $\mathds{R}^{n-5}\times C(Y)$ where $Y$ is an Einstein 4-orbifold with isolated singularities. Moreover, let $\mathds{R}^{n-5}\times C(Y_{0})$ be the tangent cone at $x$ that splits off $\mathds{R}^{n-4}$. This implies that $Y_{0}$ is a sine suspension over $\mathbb{S}^{3}/\Gamma$ for some $\Gamma\subset O(4)$. Then the proof is completed by Theorem 1.1 and the fact that the space of the cross sections of tangent cones at $x$ is connected. ∎
4.3. Refined rectifiability
Let $X^{5}$ be a noncollapsed Ricci-bounded limit space. Recall that in [CJN21, JN21], it is proven that the top stratum $\mathcal{S}^{1}\setminus\mathcal{S}^{0}$ is contained in a countable union of bi-Lipschitz images from subsets of $\mathds{R}$, modulo some $\mathcal{H}^{1}$-measure zero set. In this section, building on the uniqueness Theorem, we will improve the bi-Lipschitz structure by showing that there is no need to discard any measure zero subset, thereby establishing Theorem 1.16. For related work on uniqueness and the singular set we refer to [CM16, NV17, HJ23, HJ24, HJ25, FL25b, HKM25].
For convenience, we assume that for some small $\varepsilon_{0}>0$
$|{\rm Ric}|\leqslant\varepsilon_{0}.$ (4.6)
By rescaling this assumption is always satisfied, without affecting the results in this section. First we note that by uniqueness Theorem 1.11, there is a constant $N(v)>0$ and a natural stratification of the singular set:
$\mathcal{S}=\mathcal{S}^{0}\sqcup\bigsqcup_{2\leqslant|\Gamma|\leqslant N}\mathcal{S}^{1}_{\Gamma},$ (4.7)
where $\mathcal{S}_{\Gamma}^{1}$ is the set of points whose unique tangent cone is $\mathds{R}\times\mathds{R}^{4}/\Gamma$, $\Gamma\subset O(4)$.
First we prove a crucial uniform symmetry propagation result, coming from the gap Theorem 1.1.
Proposition 4.13.
For any $v>0,\eta>0$ and $\varepsilon\leqslant\varepsilon_{0}(v,\eta)$, the following holds. If for some $x\in B_{1}(p)\cap\mathcal{S}^{1}_{\Gamma}$ we have $d_{GH}(B_{1}(x),B_{1}(0,x^{\prime}))\leqslant\varepsilon$ with $(0,x^{\prime})$ the vertex of the cone $\mathds{R}\times C(\mathbb{S}^{3}/\Gamma)$, then $d_{GH}(B_{r}(x),B_{r}(0,x^{\prime}))\leqslant\eta r$ for any $r\leqslant 1$.
Proof.
We argue by contradiction. First note that for each $v>0$, by volume convergence [Col97], there are only finitely many possible $\Gamma$ such that $\mathds{R}\times C(\mathbb{S}^{3}/\Gamma)$ could arise as a tangent cone at some point in $B_{1}(p)$. Hence in the following argument we can fix the group $\Gamma$.
We suppose that there exist $\nu>0$, $\eta_{0}>0$, sequences $x_{i}\in B_{1}(p_{i})\cap\mathcal{S}^{1}_{\Gamma}(X_{i})$ and $r_{i}\leqslant 1$ such that
$d_{GH}(B_{1}(x_{i}),B_{1}(0,x^{\prime}))\leqslant 1/i\quad\text{ while }\quad d_{GH}(B_{r_{i}}(x_{i}),B_{r_{i}}(0,x^{\prime}))>\eta_{0}r_{i}.$
If $\limsup_{i}r_{i}\geqslant r_{0}>0$, then it contradicts the fact that $d_{GH}(B_{1}(x_{i}),B_{1}(0,x^{\prime}))\leqslant 1/i$ when $i$ is large enough. Hence it suffices to consider the case $r_{i}\to 0$. Since $d_{GH}(B_{1}(x),B_{1}(0,x^{\prime}))<1/i$, then by volume convergence theorem [Col97] we have $|{\rm Vol}(B_{1}(x_{i}))-{\rm Vol}(B_{1}(0,x^{\prime}))|\to 0$ as $i\to\infty$. Since $x_{i}\in\mathcal{S}^{1}_{\Gamma}$, we have that the volume density at $x$ is equal to $|\Gamma|^{-1}$. Therefore, the volume ratio satisfies a pinched condition: for any $s\in(0,1]$
$|\mathcal{V}_{x_{i}}^{-\varepsilon_{0}}(s)-|\Gamma|^{-1}|\leqslant c(\varepsilon_{0})\varepsilon_{i},\text{ with }\varepsilon_{i}\to 0\text{ as }i\to\infty.$ (4.8)
Here $\varepsilon_{0}$ is the constant in (4.6).
Let $\delta_{0}$ be the gap in Theorem 1.1. Since each $x_{i}\in\mathcal{S}^{1}_{\Gamma}$, there exists a sequence $s_{i}\to 0$ such that
$d_{GH}(B_{s_{i}}(x_{i}),B_{s_{i}}(0,x^{\prime}))\leqslant\frac{\delta_{0}}{10}s_{i}.$
Without loss of generality, we may assume $r_{i}\geq s_{i}$ for each $i$. Consider $\alpha_{i}(t):=(\bar{B}_{t}(x_{i}),t^{-1}d_{i})$ for $t\in[s_{i},r_{i}]$. By Gromov’s precompactness theorem, we know that each $\alpha_{i}([s_{i},r_{i}])$ is contained in the compact space $(\mathcal{M},d_{GH})$, where $\mathcal{M}$ is the space of compact Ricci limit spaces of dimension $5$, diameter upper bound 1, volume lower bound $v$, and Ricci curvature lower bound 4-4. Hence there exists some uniform constant $N(v)<\infty$ such that there exists a finite set of scales $\{t_{i,j}\}_{1\leqslant j\leqslant N}$ satisfying
  1. (1)
    $t_{i,1}=s_{i}$ and $t_{i,N}=r_{i}$ for each $i$.
  2. (2)
    For each $i$ and each $1\leqslant j\leqslant N-1$ we have $d_{GH}(\alpha_{i}(t_{i,j}),\alpha_{i}(t_{i,j+1}))\leqslant\delta_{0}/10$.
Then let $i\to\infty$. By passing to subsequences, for each $1\leqslant j\leqslant N$, we have $\alpha_{i}(t_{i,j})$ converges to a metric space $W_{j}$. By (4.8) and Lemma 4.6, we conclude that $W_{j}$ is a metric cone and thus we write $W_{j}=C(Z_{j})$. Since $s_{i},r_{i}\to 0$, we have $|{\rm Ric}(\alpha_{i}(t_{i,j}))|\to 0$ as $i\to\infty$. By the codimension four theorem 4.3, each $Z_{j}$ is an Einstein orbifold. Hence, by our construction we have
  1. (1)
    $d_{GH}(Z_{1},\mathbb{S}^{4}/\Gamma)\leqslant\frac{\delta_{0}}{10}$;
  2. (2)
    $d_{GH}(Z_{N},\mathbb{S}^{4}/\Gamma)\geq\eta_{0}$;
  3. (3)
    For each $1\leqslant j\leqslant N$ we have $d_{GH}(Z_{j},Z_{j+1})\leqslant\frac{\delta_{0}}{10}$.
Since each $Z_{j}$ is an Einstein 4-orbifold, by Theorem 1.1 we have $Z_{j}\cong\mathbb{S}^{4}/\Gamma$. This contradicts (2). Hence, the proof is completed. ∎
Next we will apply the uniform symmetry result to construct a local bi-Lipschitz map around each point in $\mathcal{S}^{1}_{\Gamma}$. Another key ingredient is the neck decomposition developed in [JN21, CJN21] (see also [NV19]).
We define the constant $\bar{V}=1/|\Gamma|$, the volume density at those points in $\mathcal{S}^{1}_{\Gamma}$. Given $x\in X$, $r>0$ and $\xi>0$, we define the volume-pinched set $\mathcal{P}_{r,\xi}(x)$ as
$\mathcal{P}_{r,\xi}(x)\equiv\{y\in B_{4r}(x):\mathcal{V}_{y}(\xi r)\leq\bar{V}+\xi\}.$
Most constants are chosen compatibly with the neck decomposition in [CJN21, Section 10], generally satisfying
$0<\xi<\delta<\gamma<\varepsilon<\tau\ll 1.$
Proposition 4.14.
For any $p\in\mathcal{S}^{1}_{\Gamma}$, there exists some function $\phi$ defined on $B_{r}(p)$ for some radius $r>0$ such that $\phi$ is bi-Lipschitz on $B_{r}(x)\cap\mathcal{S}^{1}_{\Gamma}$: for some constant $C>1$ and any $x,y\in B_{r}(p)\cap\mathcal{S}^{1}_{\Gamma}$,
$C^{-1}\cdot d(x,y)\leqslant|\phi(x)-\phi(y)|\leqslant C\cdot d(x,y).$ (4.9)
Proof.
Fix $p\in\mathcal{S}^{1}_{\Gamma}$. After rescaling, by Proposition 4.13, there exists $R_{0}\gg 1$ and $\eta>0$ such that
$d_{GH}(B_{r}(x),B_{r}(0,x^{\prime}))\leq\eta r\text{ for any }x\in B_{4}(p)\cap\mathcal{S}^{1}_{\Gamma}\text{ and any }0<r\leq R_{0}$ (4.10)
with $(0,x^{\prime})$ the vertex of the cone $\mathds{R}\times\mathds{R}^{4}/\Gamma$. Moreover, we can assume $\mathcal{V}_{z}(\xi^{-1})\geq\bar{V}-\xi$ for any $z\in B_{4}(p)$ if $\eta$ is chosen sufficiently small.
We claim that if $d_{GH}(B_{r}(x),B_{r}(0,x^{\prime}))\leq\eta r$ with $\eta\leq\eta(v,\varepsilon,\gamma,\xi)$, then
${\rm Vol}\bigl(B_{\gamma r}(\mathcal{P}_{r,\xi}(x))\bigr)>\varepsilon\gamma^{4}r_{d}^{5}.$ (4.11)
Rescaling by $r_{d}$ and arguing by contradiction, we obtain a sequence of balls $B_{1}(x_{i})$ converging to $B_{1}(0,x^{\prime})\subset\mathds{R}\times\mathds{R}^{4}/\Gamma$ for which the above volume bound fails for fixed $v,\varepsilon,\gamma,\xi$. But $(-4,4)\times\{x^{\prime}\}\subset\mathcal{P}_{1,\xi}(0,x^{\prime})$ which satisfies ${\rm Vol}\bigl(B_{\gamma}(\mathcal{P}_{1,\xi}(0,x^{\prime}))\bigr)\geq c\,\gamma^{4}$ for some universal constant $c>0$. By volume convergence, this contradicts the failure of the bound whenever $\varepsilon<c/10$, proving the claim.
Hence we can apply the decomposition proposition for 1-symmetric balls by [CJN21, Proposition 10.5] (see also [JN21]), to obtain a decomposition of $B_{1}(p)$
$B_{1}(p)\subset\Big(\mathcal{C}_{0}\cup\mathcal{N}\cap B_{1}(x)\Big)\cup\bigcup_{b}B_{r_{b}}(x_{b})\cup\bigcup_{d}B_{r_{d}}(x_{d}),$
where
  1. (1)
    $\mathcal{N}=B_{2}(p)\setminus\Big(\mathcal{C}_{0}\cup\bigcup_{b}B_{r_{b}}(x_{b})\cup\bigcup_{d}B_{r_{d}}(x_{d})\Big)$ is a $(\delta,\tau)$-neck region;
  2. (2)
    each $B_{2r_{b}}(x_{b})$ satisfies $r_{h}(x_{b})>2r_{b}$, where $r_{b}$ is the harmonic radius;
  3. (3)
    each $B_{2r_{d}}(x_{d})$ satisfies ${\rm Vol}(B_{\gamma r_{d}}\mathcal{P}_{r_{d},\xi}(x_{d}))<\varepsilon^{2}\gamma^{4}r_{d}^{5}$ or $\mathcal{P}_{r_{d},\xi}(x_{d}))=\emptyset$.
The center set of the neck region is defined to be $\mathcal{C}=\mathcal{C}_{+}\cup\mathcal{C}_{0}$ where $\mathcal{C}_{+}=\bigcup_{b}\{x_{b}\}\cup\bigcup_{d}\{x_{d}\}$ with the corresponding radius. And if $x\in\mathcal{C}_{0}$, then $r_{x}=0$. Recall that the associated measure $\mu$ to the neck region $\mathcal{N}$ is defined by $\mu=\sum_{x\in\mathcal{C}_{+}}r_{x}\delta_{x}+\mathcal{H}^{1}|_{\mathcal{C}_{0}}$. By [CJN21, Theorem 2.9], for any $x\in\mathcal{C}$ and $r\geq r_{x}$ with $B_{2r}(x)\subset B_{2}(p)$, $\mu$ is Ahlfors regular: for some universal constant AA
$A^{-1}\cdot r\leq\mu(B_{r}(x))\leq A\cdot r.$ (4.12)
We claim that $\mathcal{C}_{0}\cap B_{2}(p)=\mathcal{S}^{1}_{\Gamma}\cap B_{2}(p)$. In fact, by cone-splitting 4.7 and $\epsilon$-regularity 4.8, the region $\mathcal{N}\cup\bigcup_{b}B_{r_{b}}(x_{b})$ is regular. It suffices to check that $B_{2r_{d}}(x_{d})\cap\mathcal{S}^{1}_{\Gamma}=\emptyset$. Suppose there exists some $x\in B_{2r_{d}}(x_{d})\cap\mathcal{S}^{1}_{\Gamma}$. Then by (4.10) and (4.11), we have $B_{r_{d}/2}(x)$ is $(1,\eta)$-symmetric. Therefore ${\rm Vol}(B_{\gamma r_{d}}\mathcal{P}_{r_{d},\xi}(x_{d}))>{\rm Vol}(B_{\gamma r_{d}/2}\mathcal{P}_{r_{d}/2,\xi}(x))>\varepsilon^{2}\gamma^{4}r_{d}^{5}$, contradicting the definition of the $d$-balls. This implies the claim.
Hence it suffices to find a function that is bi-Lipschitz on $\mathcal{C}_{0}$. By [CJN21, Proposition 9.3], there exists some $\delta^{\prime}$-splitting function $u$ defined on $B_{4}(p)$ such that for any $x,y\in\mathcal{C}$
$\frac{1}{2}\cdot d(x,y)^{2}\leq|u(x)-u(y)|\leq 2\cdot d(x,y).$
In particular, $u$ is 1-1 from $\mathcal{C}$ to some subset in $\mathds{R}$. We can now define the function
$\phi(x)\equiv\mu\Big(\{z\in B_{2}(p)\cap\mathcal{C}:u(z)\leq u(x)\}\Big).$
We will prove the bi-Lipschitz property for such $\phi$: for some universal constant $C>1$, and any $x,y\in\mathcal{C}_{0}$
$C^{-1}\cdot d(x,y)\leq|\phi(x)-\phi(y)|\leq C\cdot d(x,y).$
We fix two points $x,y\in\mathcal{C}_{0}$ with $d(x,y)=s$. Assume that $u(x)>u(y)$. Let $\Omega\equiv\{z\in B_{2}(p)\cap\mathcal{C}:u(y)<u(z)\leq u(x)\}$. Then $|\phi(x)-\phi(y)|=\mu(\Omega)$.
By [CJN21, Theorem 7.2], there exists some number $T=T_{x,4s}>0$ such that $Tu$ is a $\delta^{\prime\prime}$-splitting function on $B_{4s}(x)$. By Theorem 4.10, $B_{2s}(x)$ is $(1,\delta^{\prime\prime\prime})$-symmetric with respect to $\mathcal{L}_{x,2s}$ and $(Tu,v)$ is a $\delta^{\prime\prime\prime}$-GH map for some map $v:B_{2s}(p)\to\mathds{R}^{4}/\Gamma$. Hence it is easy to see that $\Omega\equiv\{z\in B_{4s}(x)\cap\mathcal{C}:Tu(y)<Tu(z)\leq Tu(x)\}$. This proves the upper bound for $\mu(\Omega)$ using (4.12) since in particular $\Omega\subset B_{4s}(x)$
$\mu(\Omega)\leq\mu(B_{4s}(x))\leq(4A)\cdot s.$ (4.13)
On the other hand, by the Reifenberg property of $\mathcal{C}$ (See [JN21, Definition 3.1 (n3)]), there exists some point $z\in B_{2s}(x)\cap\mathcal{C}$ such that $Tu|_{B_{s/10}(z)}\subset\big(Tu(y),Tu(x)\big)\subset\mathds{R}$. Moreover, by the Lipschitz property of the radius function on $\mathcal{C}$, we have $r_{z}\leq s/100$ since $r_{p}=r_{q}=0$. Therefore, we can apply (4.12) to obtain
$\mu(\Omega)\geq\mu(B_{s/10}(z))\geq(10A)^{-1}\cdot s.$ (4.14)
Combining (4.13) and (4.14) we complete the proof of the bi-Lipschitz property of $\phi$.
Proof of Theorem 1.16.
We can decompose the singular set as
$\mathcal{S}\cap B_{1}(p)=\bigcup_{i\in I}(\mathcal{S}^{1}_{\Gamma_{i}}\cap B_{1}(p))\cup(\mathcal{S}^{0}\cap B_{1}(p)).$
Since we have the uniform lower bound of ${\rm Vol}(B_{1}(x))$ for $x\in B_{1}(p)$, the index set $I$ is finite. For each $x\in\mathcal{S}^{1}_{\Gamma_{i}}\cap B_{1}(p)$, by Proposition 4.14, there exists some radius $r$ such that $B_{r}(x)\cap\mathcal{S}^{1}_{\Gamma_{i}}$ is bi-Lipschitz to a subset of $\mathds{R}$. Since $X$ is separable, there exist countably many subsets $Z_{j}\in\mathcal{S}^{1}$ such that each $Z_{j}$ is bi-Lipschitz to some subset of $\mathds{R}$ with $\mathcal{S}^{1}=\bigcup_{j\in J}Z_{j}$. The proof is completed by noting that $\mathcal{S}^{0}$ is also countable. ∎
We include another corollary of the uniqueness theorem here: if a curve has constant volume density, then all the tangent cones on the curve are the same. This will be useful to prove the structure theorem for curves of singularities 1.18 in the next section.
Proposition 4.15.
Let $\{x_{j}\}$ be a sequence in $B_{1}(p)\cap\mathcal{S}_{\Gamma}^{1}$ that converges to $y$. Suppose that the volume ratio of the tangent cones at $y$ be the same as the cone $\mathds{R}\times\mathds{R}^{4}/\Gamma$. Then $y\in\mathcal{S}^{1}_{\Gamma}$, that is, any tangent cone at $y$ is uniquely $\mathds{R}\times\mathds{R}^{4}/\Gamma$.
Proof.
It is trivial when $\Gamma=\{id\}$. Hence we assume $\Gamma\neq\{id\}$. Let $C(Y)$ be a tangent cone at $y$ with vertex $y^{*}$ from a rescaling sequence $r_{i}\downarrow 0$. By assumption we have ${\rm Vol}(B_{r}(y^{*}))={\rm Vol}(B_{r}(0,x^{\prime}))$ for any $r>0$ where $(0,x^{\prime})$ is the vertex of $\mathds{R}\times\mathds{R}^{4}/\Gamma$. By picking large $i$ we can assume $|{\rm Vol}(B_{r_{i}}(y))-{\rm Vol}(B_{r_{i}}(0,x^{\prime}))|\leqslant\varepsilon_{i}r_{i}^{5}$ where $\varepsilon_{i}\downarrow 0$ as $i\uparrow\infty$. Since the sequence $x_{j}$ converges to $y$, then for each $i$ there exists $x_{i}$ such that $|{\rm Vol}(B_{r_{i}}(x_{i}))-{\rm Vol}(B_{r_{i}}(y))|\leqslant\varepsilon_{i}r_{i}^{5}$ and $d_{GH}(B_{r_{i}}(x_{i}),B_{r_{i}}(y))\leqslant\varepsilon_{i}r_{i}$. On the other hand, since each $x_{i}\in\mathcal{S}^{1}_{\Gamma}$, there exists some $s_{i}<r_{i}$ such that $d_{GH}(B_{s_{i}}(x_{i}),B_{s_{i}}(0,x^{\prime}))<\varepsilon_{i}s_{i}$.
Consider $\alpha_{i}(t)\equiv(\bar{B}_{t}(x_{i}),t^{-1}d)$ for $t\in[s_{i},r_{i}]$. Since $|{\rm Vol}(B_{r_{i}}(x_{i}))-{\rm Vol}(B_{r_{i}}(0,x^{\prime}))|<2\varepsilon_{i}r_{i}^{5}$, By Lemma 4.6, we have $\alpha_{i}(t)$ is $\delta_{i}$-close to a cone for any $t\in[s_{i},r_{i}]$ with $\delta_{i}\downarrow 0$ as $i\uparrow\infty$. Following the proof of Proposition 4.13, by Gromov’s precompactness theorem and Theorem 4.3, we have that each $\alpha_{i}(t)$ converges to the unit ball in $C(Z_{t})$ where each $Z_{t}$ is an Einstein 4-orbifold. Moreover, we have that $\alpha_{i}(s_{i})\to B_{1}(0,x^{\prime})\subset\mathds{R}\times\mathds{R}^{4}/\Gamma$ and $\alpha_{i}(r_{i})\to B_{1}(y^{*})\subset C(Y)$. By the isolation Theorem 1.1, this implies that $\alpha_{i}(t)\to B_{1}(0,x^{\prime})$ for any $t\in[s_{i},r_{i}]$ and in particular $C(Y)\cong\mathds{R}\times\mathds{R}^{4}/\Gamma$. The proof is completed by the uniqueness Theorem 1.11. ∎
Corollary 4.16.
Let $\gamma$ be a continuous curve in $B_{1}(p)\subset X$. Assume that the volume density is constant along the curve. Then there exists some $\Gamma\subset O(4)$ such that all of the tangent cones along the curve are isometric to $\mathds{R}\times\mathds{R}^{4}/\Gamma$.
Proof.
Suppose that the curve has the Euclidean volume ratio. Then the whole curve is in the regular set by volume convergence Theorem [Col97] and all the tangent cones must be Euclidean. Therefore, we can assume $\gamma\subset\mathcal{S}$. We can decompose
$\gamma=\mathcal{S}^{0}(\gamma)\sqcup\bigsqcup_{i=1}^{N}\mathcal{S}^{1}_{\Gamma_{i}}(\gamma).$
By Proposition 4.15, each $\mathcal{S}^{1}_{\Gamma_{i}}(\gamma)$ is closed relative to $\gamma$. This implies that $\mathcal{S}^{0}(\gamma)$ is open relative to $\gamma$. Since we have the dimension estimate $\dim(\mathcal{S}^{0})\leqslant 0$, it implies that $\mathcal{S}^{0}(\gamma)=\emptyset$. Since each $\mathcal{S}^{1}_{\Gamma_{i}}(\gamma)$ is closed relative to $\gamma$, we have $\gamma=\mathcal{S}^{1}_{\Gamma_{i}}$ for some $i$. ∎
5. Orbifold structure for curves of singularities
In this section, we consider a sequence $(M_{i},g_{i},p_{i})_{i}$ of Einstein $5$-metrics with
${\rm Vol}(B_{1}(p_{i}))>v>0,\qquad\text{ and}\qquad|{\rm Ric}(g_{i})|\leqslant 4$
for uniform constants $v>0$. Then, one can take a subsequence converging in the Gromov-Hausdorff sense to a singular space $(X,d,p)$.
In the previous section we analyzed the regularity of points in $\mathcal{S}^{1}\setminus\mathcal{S}^{0}$. We now consider the case when the singularities exhaust an entire curve. In this setting we show that the curve of singularities in $\mathcal{S}^{1}_{\Gamma}$ consists of orbifold singularities lying along geodesics, and that the curvature remains uniformly bounded in a neighborhood of the curve; see Theorem 1.18.
We consider a Lipschitz curve $\gamma:[-2,2]\to\mathcal{S}^{1}_{\Gamma}$ with $\gamma(0)=x_{0}$ and $\Gamma\neq\{id\}$. In the following we let $(0,x^{\prime})$ be cone tip of $\mathds{R}\times\mathds{R}^{4}/\Gamma$. By Proposition 4.14, we can assume $\gamma$ is $(1+10^{-10})$-bi-Lipschitz by rescaling:
$\Big|d(\gamma(t),\gamma(s))-|t-s|\Big|\leqslant 10^{-10}|t-s|.$ (5.1)
We recall the definition of 5-dimensional orbifolds with curve singularities:
Definition 5.1 ($5$-orbifold with curve singularities).
We will say that a metric space $(M_{o},g_{o})$ is a 55–dimensional orbifold with singularities along curves if there exist $\varepsilon_{0}>0$ and a countable collection of embedded smooth curves $(\gamma_{k})_{k}$ in $M_{o}$, called singular curves, such that the following properties hold:
  1. (1)
    the space $\bigl(M_{o}\setminus\cup_{k}\gamma_{k},\,g_{o}\bigr)$ is a smooth manifold of dimension $5$,
  2. (2)
    for each singular curve $\gamma_{k}$ of $M_{o}$ and every point $p\in\gamma_{k}$, there exist
    • an open neighborhood $U_{k,p}\subset M_{o}$ of $p$,
    • a finite subgroup $\Gamma_{k}\subset O(4)$ acting freely on $\mathbb{S}^{3}$,
    • a diffeomorphism $\Phi_{k,p}\colon B_{e}(0,\varepsilon_{0})\subset\mathbb{R}\times\bigl(\mathbb{R}^{4}/\Gamma_{k}\bigr)\longrightarrow U_{k,p}\subset M_{o},$ such that:
      1. (a)
        $\Phi_{k,p}\bigl(\mathbb{R}\times\{0\}\cap B_{e}(0,\varepsilon_{0})\bigr)=\gamma_{k}\cap U_{k,p}$,
      2. (b)
        the pull-back metric $\Phi_{k,p}^{*}g_{o}$ lifts, on the covering $\mathbb{R}\times\mathbb{R}^{4}$, to a Riemannian metric.
We say that the orbifold has regularity $C^{k,\alpha}$, (resp. is real-analytic) if all of the metrics $\Phi_{k,p}^{*}g_{o}$ are.
The techniques in this section can be applied to higher dimensional case up to minor modifications, around a codimension four Lipschitz submanifold of singular set where all the tangent cones are $\mathds{R}^{n-4}\times\mathds{R}^{4}/\Gamma$ for the same $\Gamma$. For the gluing part 5.1, it is more technically complicated in higher dimension so for simplicity we prove it in dimension $n=5$. However, for the regularity upgrading part 5.2 we will work in all dimensions to keep track of the dimensional constants from Sobolev inequality.
5.1. Continuous metric extension
In the first step, we establish the $C^{0}$ extension of the metric near the singular curve $\gamma$. Our approach follows the strategy in [DS14] where the codimension $4$ singularities of $6$-dimensional Kähler-Einstein metrics reduces to the isolated singularity case in dimension four. In the case of isolated singularities, one typically constructs a diffeomorphism between an annulus surrounding the singular point and a corresponding annulus in the tangent cone. These diffeomorphisms are then glued together along the radial direction. The uniqueness of tangent cone then ensures that the error decays as the radial scale tends to zero.
Several key new difficulties have to be addressed in the case of a higher-dimensional singular set. At each fixed radial scale, one must first glue together the various local blocks along the directions parallel to the singular set. As the radial scale tends to zero, the number of such blocks increases, making uniform estimates essential to prevent error accumulation. Moreover, the gluing procedure must be carefully implemented to avoid shrinking the annuli in proportion to the number of blocks. We will overcome these difficulties in Lemma 5.2 and Lemma 5.4.
We first state a key uniform symmetry estimate used to control errors in the gluing procedure. The key point is that the symmetry scale $r_{0}$ is independent of the point $x$ on the curve.
Lemma 5.2.
Let $\eta>0$ be given, $x\in\mathcal{S}^{1}_{\Gamma}$ and $B_{r}(0,x^{\prime})$ metric balls centered at the tip of the cone $\mathbb{R}\times\mathbb{R}^{4}/\Gamma$. Then there exists some $r_{0}(\eta)$ such that the following estimate holds
$d_{GH}(B_{r}(x),B_{r}(0,x^{\prime}))\leqslant\eta r,\text{ for any }x\in\gamma\text{ and any }r\leqslant r_{0}.$ (5.2)
Moreover, for each $r_{0}>0$, there exists $\eta$, with $\eta\to 0$ as $r_{0}\to 0$, such that (5.2) holds.
Proof.
Let $\eta>0$ be given. Then by Proposition 4.13, there exists some $\varepsilon_{0}(\eta)$ such that:
[…]
Therefore we can define
$r_{x}\equiv\sup\big\{r:d_{GH}(B_{s}(x),B_{s}(0,x^{\prime}))<\tfrac{\varepsilon_{0}s}{2}\text{ for any }s\leqslant r\big\}>0.$
Then for any $y\in B_{\varepsilon_{0}r_{x}/2}(x)\cap\gamma$, we have
$d_{GH}(B_{r_{x}}(y),B_{r_{x}}(0,x^{\prime}))\leqslant d_{GH}(B_{r_{x}}(y),B_{r_{x}}(x))+d_{GH}(B_{r_{x}}(x),B_{r_{x}}(0,x^{\prime}))\leqslant\frac{\varepsilon_{0}r_{x}}{2}+\frac{\varepsilon_{0}r_{x}}{2}=\varepsilon_{0}r_{x}.$
Hence we have $d_{GH}(B_{r}(y),B_{r}(0,x^{\prime}))\leqslant\eta r$ for any $r\leqslant r_{x}$.
Since $\gamma$ is a closed and thus compact subset, there exist finitely many $i$ such that
$\gamma=\bigcup_{x\in\gamma}B_{\varepsilon_{0}r_{x}/2}(x)\cap\gamma=\bigcup_{i=1}^{M}B_{\varepsilon_{0}r_{x_{i}}/2}(x_{i})\cap\gamma.$
The proof is finished if we choose $r_{0}=\min\{r_{x_{1}},...,r_{x_{M}}\}$. Moreover, note that if $\eta\to 0$, then $\varepsilon_{0}(\eta)\to 0$ and thus $r_{x}\to 0$. This proves the second claim. ∎
Let $g_{0}$ denote the standard flat metric in $\mathds{R}\times\mathds{R}^{4}/\Gamma$ with vertex $(0,x^{\prime})$. We define the annuli as
$(b_{1},b_{2})\times A_{a_{1},a_{2}}\equiv\{(t,z)\in\mathds{R}\times\mathds{R}^{4}/\Gamma:b_{1}<t<b_{2},a_{1}<|z|<a_{2}\}.$
For convenience, we define $|x|_{\Gamma}\equiv\|z\|$ for any $x=(t,z)$.
Suppose $B_{200r}(x)$ be $(1,\eta)$-symmetric with respect to $\mathcal{L}_{x,200r}$ for some $\eta<1/10$, i.e.
$d_{GH}(B_{200r}(x),B_{200r}(0,x^{\prime}))\leqslant 200\eta r$
and $\mathcal{L}_{x,200r}$ is the Gromov-Hausdorff correspondence of $(\mathds{R}\times\{0\})\cap B_{200r}(0,x^{\prime})\subset\mathds{R}\times\mathds{R}^{4}/\Gamma$. Then by $\varepsilon$-regularity due to Anderson [And89], Colding [Col97], Cheeger-Naber [CN15], and the uniqueness Theorem 1.11, for any $\delta>0$, if $\eta\leqslant\eta_{0}(\delta)$, there exists some smooth embedding $\phi_{\eta}:[-20r,20r]\times{\bar{A}}_{r,100r}\to B_{200r}(x)\setminus B_{2\eta}(\mathcal{L}_{x,200r})$ satisfying
  1. (1)
    $|r^{-2}\phi_{\eta}^{*}g-g_{0}|_{C^{4}}<\delta$,
  2. (2)
    For any $y\in(-20r,20r)\times{\bar{A}}_{r,100r}$ we have
    $(1-\delta)r|y|_{\Gamma}\leqslant d(\phi_{\eta}(y),\mathcal{L}_{x,200r})\leqslant(1+\delta)r|y|_{\Gamma}.$ (5.3)
The following approximation by isometry lemma is standard and proven like [DS14, Lemma 5.8].
Lemma 5.3.
Let $\delta>0$ be given small. Then there exists some $\delta^{\prime}$, with $\delta^{\prime}\to 0$ as $\delta\to 0$, such that for any smooth map $\phi:[11,19]\times\bar{A}_{1,100}\to\mathds{R}\times\mathds{R}^{4}/\Gamma$ with $|\phi^{*}g_{0}-g_{0}|_{C^{4}}\leqslant\delta$, then there exists some isometry $P_{i}\in\operatorname{Isom}(\mathds{R})\times\operatorname{Isom}(\mathds{R}^{4}/\Gamma)$ such that $|P_{i}\circ\phi-\mathrm{id}|_{C^{3}}<\delta^{\prime}$ in a smaller region.
Proof.
The proof follows from contradiction arguments. ∎
Now we glue the local diffeomorphisms together along the direction parallel to the singular curve $\gamma$. The subtle point is that the domains $(-1,1)\times\bar{A}_{10r,90r}$ are not scale invariant; however, all estimates remain uniform in $r$, independent of the parameter $t$. Crucially, $\operatorname{Isom}(\mathds{R})\times\operatorname{Isom}(\mathds{R}^{4}/\Gamma)$ is much smaller than $\operatorname{Isom}(\mathds{R}^{5})$. The isometry $P$ from Lemma 5.3 can only be a translation or reflection in $\mathds{R}$, combined with a rotation respecting $\Gamma$-action on $\mathds{R}^{4}$.
Lemma 5.4.
For any $r$ small, there exists a smooth embedding $H_{r}:(-1,1)\times\bar{A}_{10r,90r}\to B_{2}(x_{0})$ such that
$(1-\delta)|y|_{\Gamma}\leqslant d(H_{r}(y),\gamma)\leqslant(1+\delta)|y|_{\Gamma},\text{ and }|r^{-2}H_{r}^{*}g-g_{0}|_{C^{2}}\leqslant\delta(r).$ (5.4)
And $\delta(r)\to 0$ as $r\to 0$.
Proof.
Let $r\in(0,10^{-2})$ be given small. Let $\eta=\eta(r)>0$ be as in Lemma 5.2 and $\phi_{\eta,x}$ be the embedding defined in previous discussion for each $x\in\gamma$.
Consider the point set $\{x_{i}\}$ where $x_{i}=\gamma(30ir)$ for integers $i\in[-N(r),N(r)]$ with $N(r)\leqslant C/r$ and their associated maps $\phi_{i}\equiv\phi_{\eta,x_{i}}$. Define $A_{i}\equiv\phi_{i}((-20r,20r)\times{\bar{A}}_{r,100r})\subset B_{2}(x_{0}).$
If we choose $r$ small and thus $\eta$ small enough, then by (5.1) and $\varepsilon$-regularity Theorem 4.8, for $x_{1}=\gamma(30r)\in B_{100r}(x_{0})$ we have
$d_{H}(\mathcal{L}_{x_{0},200r}\cap B_{100r}(x_{0}),\mathcal{L}_{x_{1},200r}\cap B_{100r}(x_{0}))\leqslant 10^{-10}r.$
Similar inequality holds for general $x_{i}$ and $x_{i+1}$. Hence by (5.3) we can conclude that
  1. (1)
    $\bigcup_{i}A_{i}\subset\big\{x\in X:0.9r<d(x,\gamma)<101r\big\}.$
  2. (2)
    For any |ij|>1|i-j|>1, we have $A_{i}\cap A_{j}=\emptyset$.
  3. (3)
    Let $D_{i}\equiv A_{i}\cap A_{i+1}$. Then we have $(11r,19r)\times\bar{A}_{2r,99r}\subset\phi_{i}^{-1}(D_{i})$ and $(-19r,-11r)\times\bar{A}_{2r,99r}\subset\phi^{-1}_{i+1}(D_{i})$ for any such $i$.
In the following we assume $r=1$ for simplicity. First we glue $A_{i}$ and $A_{i+1}$ together.
Consider the map $\Psi_{i}\equiv\phi_{i+1}^{-1}\circ\phi_{i}:[11,19]\times\bar{A}_{2,99}\to[-20,20]\times\bar{A}_{1,100}$. Then $\Psi_{i}$ satisfies
$(1-3\delta)|x|_{\Gamma}\leqslant|\Psi_{i}(x)|_{\Gamma}\leqslant(1+3\delta)|x|_{\Gamma}\quad\text{ and }\quad|\Psi_{i}^{*}g_{0}-g_{0}|_{C^{4}}\leqslant C_{1}\delta,$
with constant $\delta$ as in (5.3). By Lemma 5.3, there exists some isometry $P\in\operatorname{Isom}(\mathds{R})\times\operatorname{Isom}(\mathds{R}^{4}/\Gamma)$ such that $|P\circ\Psi_{i}-id|_{C^{3}}\leqslant C_{2}(\delta)$ on $(12,18)\times A_{3,98}$. Then up to some reflection in $t$-direction, we can pick $P_{i}=(T_{+30},R_{i})$, where $T_{+30}(t,z)=(t+30,z)$ is a translation along $t$-direction and that $R_{i}\in\operatorname{Isom}(\mathds{R}^{4}/\Gamma)$ is a rotation, such that $|P_{i}\circ\Psi_{i}-id|_{C^{3}}\leqslant C_{2}(\delta)$.
Now we choose a smooth cut-off function $\chi\in C^{\infty}(im(\phi_{i}))$ using the map $\phi_{i}$ satisfying
  1. (1)
    $0\leqslant\chi\leqslant 1$ in $im(\phi_{i})=A_{i}$.
  2. (2)
    $\chi(\phi_{i}(t,z))=1$ for $z\in\bar{A}_{1+\theta(t),100-\theta(t)}$ and $t\in[-20,14]$.
  3. (3)
    $\chi(\phi_{i}(t,z))=0$ for $t\geqslant 16$.
  4. (4)
    $\chi(\phi_{i}(t,z))=0$ for $|z|\notin\bar{A}_{2,99}$ for $t\geqslant 9$.
Here $\theta(t)$ is a non-decreasing function with $\theta(t)\equiv 0$ for $t\leqslant 5$ and $\theta(t)\equiv 2$ for $t\geqslant 9$. We can trivially extend $\chi$ by assuming $\chi\equiv 0$ so that $\chi$ is defined smoothly in $A_{i}\cup A_{i+1}$ satisfying $|\chi|_{C_{g}^{4}}\leqslant C_{3}$.
Consider the pinched annulus $\Omega_{i}$ defined as a smooth connected region $\Omega_{i}\equiv\Omega_{i,1}\cup\Omega_{i,2}\subset A_{i}\cup A_{i+1}$ where
$\Omega_{i,1}$ $\equiv\{\phi_{i}(t,z):-20\leqslant t\leqslant 20,\text{ and }z\in\bar{A}_{1+2\theta(t),100-2\theta(t)}\};$
$\Omega_{i,2}$ $\equiv\{\phi_{i+1}(t,z):-20\leqslant t\leqslant 20,\text{ and }z\in\bar{A}_{1+\theta(-t),100-\theta(-t)}\}.$
We construct a map defined on $\Omega_{i}\subset A_{i}\cup A_{i+1}$ by
$h_{i}\equiv(1-\chi)\cdot P_{i}\circ\phi^{-1}_{i+1}+\chi\cdot\phi_{i}^{-1}.$ (5.5)
Note that $h_{i}$ satisfies that
  1. (1)
    $h_{i}(x)=P_{i}\circ\phi_{i+1}^{-1}(x)$ if $x\in\Omega_{i}\setminus A_{i}$,
  2. (2)
    $h_{i}(x)=\phi_{i}^{-1}(x)$ if $x\in\Omega_{i}\setminus A_{i+1}$,
  3. (3)
    $(1-C_{4}(\delta))|h_{i}(x)|_{\Gamma}\leqslant d(x,\gamma)\leqslant(1+C_{4}(\delta))|h_{i}(x)|_{\Gamma}$,
  4. (4)
    $|h_{i}^{*}g_{0}-g|_{C^{2}_{g}}\leqslant C_{5}(\delta)$.
Recall that $P_{i}=(T_{+30},R_{i})$ with $R_{i}$ a rotation of $\mathds{R}^{4}/\Gamma$. Hence $h_{i}$ is a diffeomorphism onto $h_{i}(\Omega_{i})$ where
  1. (1)
    $\Big((-20,5)\cup(25,50)\Big)\times\bar{A}_{1,100}\subset h_{i}(\Omega_{i})$.
  2. (2)
    $(-20,50)\times\bar{A}_{10,90}\subset h_{i}(\Omega_{i})$.
We can then glue $\Omega_{i}$ and $A_{i-1}$. Since the left end of $\Omega_{i}$ is the same as $A_{i}$, and the pinched region $D_{i}$ is disjoint with $A_{i-1}$, we can glue these two using the same arguments with the uniform estimate: there exists a smooth embedding $H_{i-1}:\tilde{\Omega}_{i-1}\to\mathds{R}\times\mathds{R}^{4}/\Gamma$ such that
  1. (1)
    $\Big((-50,-25)\cup(25,50)\Big)\times\bar{A}_{1,100}\subset H_{i-1}(\tilde{\Omega}_{i})$,
  2. (2)
    $(-50,50)\times\bar{A}_{10,90}\subset H_{i-1}(\tilde{\Omega}_{i})$,
  3. (3)
    $(1-C_{4}(\delta))|H_{i-1}(x)|_{\Gamma}\leqslant d(x,\gamma)\leqslant(1+C_{4}(\delta))|H_{i-1}(x)|_{\Gamma}$,
  4. (4)
    $|H_{i-1}^{*}g_{0}-g|_{C^{2}_{g}}\leqslant C_{5}(\delta)$.
The same holds for the extension to the right end. Note that at each gluing, we only modify the overlapping region but keep the non-overlapping ends unchanged. Therefore, after at most finitely many times gluing, the estimates hold uniformly and we obtain a diffeomorphism satisfying the required properties. ∎
Hence we can glue all $t$-directional blocks together with uniform estimates. Then we can glue these annuli along the $\Gamma$-direction.
Proposition 5.5.
For any small $r$, there exists a diffeomorphism $F:B^{*}_{r}(0)\subset\big(\mathds{R}\times\mathds{R}^{4}/\Gamma\setminus\mathds{R}\times\{0\}\big)\to B_{2}(x_{0})\setminus\gamma$ such that $F^{*}g$ extends continuously to a metric tensor over the universal cover $\hat{B}_{r}(0)\subset\mathds{R}\times\mathds{R}^{4}$.
Proof.
For any small $r$, by Lemma 5.4, there exists a smooth diffeomorphism $H_{r}$ satisfying properties (5.4) with error $\delta(r)\downarrow 0$ as $r\downarrow 0$. Then we apply similar gluing arguments as in Lemma 5.4 to the map $V_{0}\equiv H_{r_{0}}$ with some fixed small $r_{0}$, and the map $H_{2^{-1}r_{0}}$, to obtain a new diffeomorphism $V_{1}:(-1,1)\times\bar{A}_{5r_{0},90r_{0}}\to B_{2}(x_{0})\setminus\gamma$ with estimates
$(1-\delta^{\prime})|y|_{\Gamma}\leqslant d(V_{1}(y),\gamma)\leqslant(1+\delta^{\prime})|y|_{\Gamma},\text{ and }|V_{1}^{*}g-g_{0}|_{C^{2}}\leqslant\delta^{\prime}.$
Then by induction we can glue the maps $V_{i}$ and $H_{2^{-{i}}r_{0}}$ together to obtain $V_{i+1}$ satisfying the same estimates above with $\delta^{\prime}(i)\to 0$ as $i\to\infty$. Therefore we get a map $F:B^{*}_{r_{0}}(0)\to B_{2}(x_{0})$, which satisfies
$|F^{*}g-g_{0}|_{C^{0}(-1,1)\times\bar{A}_{2^{-i}r_{0},90r_{0}}}\leqslant\delta^{\prime}(i).$
Since $\delta^{\prime}(i)\to 0$ as $i\to\infty$, $F^{*}g$ extends continuously in $\hat{B}_{r_{0}}(0)$. ∎
5.2. $L^{\infty}$ curvature bound
In this section we prove that the curvature is bounded around the singular curve $\gamma:[-2,2]\to\mathcal{S}^{1}_{\Gamma}$ with $\gamma(0)=x_{0}$. By Lemma 5.2 and rescaling, we can simply assume that there exists some small $\eta>0$ such that
$d_{GH}(B_{r}(x),B_{r}(0,x^{\prime}))\leqslant\eta r,\text{ for any }x\in\gamma([-1,1])\text{ and any }r\leqslant 1.$ (5.6)
The main theorem of this section is the following uniform $L^{\infty}$ curvature bound:
Theorem 5.6.
If $\eta$ is small enough, then for any integer $k\geq 0$, there exists some constant $C_{k}(\eta)$ such that
$\sup_{B_{1}(x_{0})}|\nabla^{k}{\rm Rm}|\leqslant C_{k}(\eta).$ (5.7)
By Proposition 5.5, we have a $C^{0}$ coordinate chart for the ball $B_{1}(x_{0})$ and the metric is $C^{0}$-close to the standard metric in $\mathds{R}\times\mathds{R}^{4}/\Gamma$. Since we will focus on the curvature estimate in this section, without loss of generality we study the universal cover of the ball $B_{1}(x_{0})$. The sobolev inequalities hold in $B_{1}(x_{0})$ thanks to [Cro80]. To keep track of the dimensional constants, we will work in any dimension $n\geq 5$ in this section.
We define $B\equiv B_{1}^{x}(0^{n-4})\times B_{1}^{y}(0^{4})$ and the ball minus a segment $B^{*}\equiv B\setminus(B_{1}^{x}(0^{n-4})\times\{0\})$. Also, we let $\nabla_{x},\nabla_{y}$ denote the gradient along $x,y$ direction respectively.
In the following we consider the equation which holds in $B^{*}$:
$\Delta u\geqslant-fu.$ (5.8)
It will be applied to $f=|{\rm Rm}|$ and both $u=|\nabla^{k}{\rm Rm}|$ or $u=|{\rm Rm}|^{1-\alpha}$ for $k\in\mathbb{N}$ and some dimensional $\alpha$.
We will adapt Moser’s iteration argument to prove that $|{\rm Rm}|$ is bounded. To achieve this, we first prove the following Caccioppoli type inequality, motivated by [Sib85].
Lemma 5.7.
Let $u\in C^{1}(B^{*})$ and $u\geqslant 0$ satisfying the equation (5.8) in the sense of distributions. Suppose $u\in L^{2q}$ for some $q\geqslant 1+\varepsilon>1$ and $|f(x,y)|\leqslant\delta|y|^{-2}$ for some $\delta\leqslant\delta_{0}(\varepsilon,q)$. Then for any $\eta\in C^{\infty}_{c}(B)$ we have
$\int_{B}\eta^{2}|\nabla u^{q}|^{2}\leqslant C(\varepsilon)q\int_{B}|\nabla\eta|^{2}u^{2q}.$ (5.9)
Proof.
By integration by parts, the following holds for any function $\chi\in C_{c}^{\infty}(B^{*})$ with compact support in $B^{*}$
$\int_{B}\nabla\chi\cdot\nabla u\leqslant\int_{B}\chi fu.$ (5.10)
We define the following test function for some large constant $K>0$
$F(t)=\frac{t^{q}}{(1+\frac{t}{K})^{q/2}}$ (5.11)
whose motivation is that if $u\in L^{2q}$, then $F(u)\in L^{4}$ since we temper the large values of the function.
Note that the function $F$ is $C^{\infty}$ for $t>0$. We define $G(t)\equiv F(t)F^{\prime}(t)$. Then the following properties hold by direct computation:
$F(t)$ $\leqslant K^{q/2}t^{q/2};$ (5.12)
$tG(t)$ $\leqslant qF(t)^{2};$ (5.13)
$G^{\prime}(t)$ $\geqslant c(\varepsilon)F^{\prime}(t)^{2}.$ (5.14)
Now we pick $\eta,\bar{\eta}_{\varepsilon}\in C_{c}^{\infty}(B)$ and further let $\bar{\eta}_{\varepsilon}$ vanish in $B_{1}^{x}\times B_{\varepsilon}^{y}$. Choose the cutoff function $\chi\equiv(\eta\bar{\eta}_{\varepsilon})^{2}G(u)$ and by (5.10) and (5.13) we have
$\int_{B}\nabla\Big((\eta\bar{\eta}_{\varepsilon})^{2}G(u)\Big)\cdot\nabla u\leqslant\int_{B}(\eta\bar{\eta}_{\varepsilon})^{2}G(u)fu\leqslant q\int_{B}(\eta\bar{\eta}_{\varepsilon})^{2}|f||F(u)|^{2}.$ (5.15)
Note that $\nabla F(u)=F^{\prime}(u)\nabla u$ and $\nabla G(u)=G^{\prime}(u)\nabla u$. Hence by (5.14) we have $\nabla G(u)\cdot\nabla u=|\nabla u|^{2}G^{\prime}(u)\geqslant c(\varepsilon)F^{\prime}(u)^{2}|\nabla u|^{2}=c(\varepsilon)|\nabla F(u)|^{2}$. Hence
$\int_{B}\nabla\Big((\eta\bar{\eta}_{\varepsilon})^{2}G(u)\Big)\cdot\nabla u$ $=\int_{B}(\eta\bar{\eta}_{\varepsilon})^{2}\nabla G(u)\cdot\nabla u+\int_{B}G(u)\nabla(\eta\bar{\eta}_{\varepsilon})^{2}\cdot\nabla u$ (5.16)
$\geqslant c(\varepsilon)\int_{B}(\eta\bar{\eta}_{\varepsilon})^{2}|\nabla F(u)|^{2}-\int_{B}|2\eta\bar{\eta}_{\varepsilon}F^{\prime}(u)\nabla u|\cdot|F(u)\nabla(\eta\bar{\eta}_{\varepsilon})|$ (5.17)
$=c(\varepsilon)\int_{B}(\eta\bar{\eta}_{\varepsilon})^{2}|\nabla F(u)|^{2}-\int_{B}|2\eta\bar{\eta}_{\varepsilon}\nabla F(u)|\cdot|F(u)\nabla(\eta\bar{\eta}_{\varepsilon})|$ (5.18)
Combining these and (5.15) we obtain
$c(\varepsilon)\int_{B}(\eta\bar{\eta}_{\varepsilon})^{2}|\nabla F(u)|^{2}\leqslant\int_{B}|2\eta\bar{\eta}_{\varepsilon}\nabla F(u)|\cdot|\nabla(\eta\bar{\eta}_{\varepsilon})F(u)|+q\int_{B}(\eta\bar{\eta}_{\varepsilon})^{2}|f|F(u)^{2}\equiv I_{1}+I_{2}.$ (5.19)
By Cauchy-Schwarz inequality, for some $c^{\prime}(\varepsilon)$ we have
$I_{1}\leqslant\frac{c(\varepsilon)}{2}\int_{B}(\eta\bar{\eta}_{\varepsilon})^{2}|\nabla F(u)|^{2}+c^{\prime}(\varepsilon)\int_{B}|\nabla(\eta\bar{\eta}_{\varepsilon})|^{2}F(u)^{2}.$ (5.20)
To estimate $I_{2}$, we apply Hardy inequality to get
$I_{2}\leqslant\delta q\int_{B_{1}^{x}}\int_{B_{1}^{y}}(\eta\bar{\eta}_{\varepsilon})^{2}F(u)^{2}|y|^{-2}dydx\leqslant 4\delta q\int_{B_{1}^{x}}\int_{B_{1}^{y}}|\nabla_{y}(\eta\bar{\eta}_{\varepsilon}\cdot F(u))|^{2}(x,y)dydx.$ (5.21)
Note that $|\nabla_{y}F(u)|^{2}\leqslant|\nabla F(u)|^{2}$. Hence by choosing $\delta\leqslant\delta_{0}(\varepsilon,q)$ small, we can absorb all terms involving $|\nabla F|$ into the left hand side of (5.19) and thus obtain
$\int_{B}(\eta\bar{\eta}_{\varepsilon})^{2}|\nabla F(u)|^{2}\leqslant C(\varepsilon)q\int_{B}|\nabla(\eta\bar{\eta}_{\varepsilon})|^{2}F^{2}(u).$ (5.22)
Here is the only place where we require $\delta$ to be small depending on $q$.
Note that we can choose $\bar{\eta}_{\varepsilon}$ such that $\eta\cdot\bar{\eta}_{\varepsilon}\to\eta$ and $\int_{B\cap\mathrm{supp}(\eta)}|\nabla\bar{\eta}_{\varepsilon}|^{4}\to 0$ as $\varepsilon\to 0$ since $B\setminus\mathrm{supp}(\bar{\eta}_{\varepsilon})\subset B_{1}^{x}\times B_{2\varepsilon}^{y}\to B_{1}^{x}\times\{0\}\subset\mathds{R}^{n-4}\times\{0\}$. Combining $F(u)\leqslant K^{q/2}u^{q/2}$ and $u\in L^{2q}$, we have $F(u)\in L^{4}$ and thus
$\int_{B}|\nabla\bar{\eta}_{\varepsilon}|^{2}\eta^{2}F^{2}(u)\leqslant C(K)\Bigg(\int_{B}|\nabla\bar{\eta}_{\varepsilon}|^{4}\Bigg)^{1/2}\cdot\Bigg(\int_{B}F(u)^{4}\Bigg)^{1/2}\to 0,\text{ as }\varepsilon\to 0.$ (5.23)
Combining all, we can now conclude that
$\int_{B}\eta^{2}|\nabla F(u)|^{2}\leqslant C(\varepsilon)q\int_{B}|\nabla\eta|^{2}F^{2}(u).$ (5.24)
Finally we let $K\to\infty$. Note that $F(u)\to u^{q}$. The proof is finished by Fatou’s Lemma. ∎
As a consequence, we get higher integrability of $u$ whenever $\delta$ is small enough.
Lemma 5.8.
Let $u\in C^{1}(B^{*})$ and $u\geqslant 0$ satisfying the equation (5.8) in the sense of distributions. Suppose $u\in L^{2q}$ for some $q\geqslant 1+\varepsilon>1$. For any $N>1$, if $|f(x,y)|\leqslant\delta|y|^{-2}$ for some $\delta\leqslant\delta_{0}(n,\varepsilon,N)$, we have
[…]
Proof.
We will prove it using Sobolev inequality and iteration argument. We assume $q<N$. Otherwise, there is nothing to prove.
Let $v_{0}=u^{q}$. Define $v_{k}=v_{k-1}^{\frac{n}{n-2}}=u^{q\cdot(\frac{n}{n-2})^{k}}$. We further define the radii $r_{k}=\frac{1}{2}+\frac{1}{2^{k+1}}$. Then we can define the cutoff functions $\eta_{k}\equiv 1$ in $B_{r_{k+1}}$ and vanishing outside $B_{r_{k}}$.
By Sobolev inequality in $B_{1}\subset\mathds{R}^{n}$ and (5.9), with $2^{*}\equiv\frac{2n}{n-2}$ we have
[…]
This implies that $v_{0}\in L^{2^{*}}(B_{r_{1}})$, equivalently $u\in L^{\frac{2qn}{n-2}}(B_{r_{1}})$. If $\frac{qn}{n-2}\geqslant N$, we finish the proof. Otherwise, we can apply (5.9) and Sobolev inequality again to $\eta_{1}v_{1}$ to obtain
[…]
Next we iterate this step for finitely many times until $N\equiv q\cdot(\frac{n}{n-2})^{M}>\frac{n}{2}$. Then by induction
[…]
Next we will use Moser iteration argument to obtain $L^{\infty}$-estimate for the following equation for some $q_{0}=q_{0}(n)>1$ in $B_{1}$
$\Delta u\geqslant-Cu^{1+q_{0}}.$ (5.28)
Note that in Lemma 5.7, we obtain a Caccioppoli type inequality when $\delta$ is small but depends on $q$. Here we can revise the proof to obtain improved estimates with small $\delta$ independent of $q$, which is crucial to run the iteration argument towards $q=\infty$. The idea is as follows. By Lemma 5.8, whenever $\delta(n,\varepsilon)$ is chosen small, we can obtain $L^{N}$-norm for $u^{q}$ with $N>n/2$, and thus for $f=u^{q_{0}}$ in the case of (5.28). Then in the proof of Lemma 5.7 we replace the estimate $|f|\leqslant\delta|y|^{-2}$ by $L^{N}$-norm of $f=u^{q_{0}}$. Note that the $L^{N}$-estimate of $f$ does not depend on $\delta$. Hence we can get similar Caccioppoli type inequality for all $q$. Also, since we assume the bound $|f(x,y)|=|u^{q_{0}}(x,y)|\leqslant\delta|y|^{-2}$, we have $u\in L^{2q}$ for some $q(q_{0})>1$ particularly.
Lemma 5.9.
Let $u\in C^{1}(B_{1}^{*})$ and $u\geqslant 0$ satisfying the equation (5.28) in the sense of distributions. Suppose $u\in L^{2+2\varepsilon}$ for some $\varepsilon>0$, and $|u^{q_{0}}(x,y)|\leqslant\delta|y|^{-2}$ with $q_{0}$ in (5.28) and some $\delta\leqslant\mathbf{\delta_{0}(n,q_{0})}$. Then for any $q>1$, we have $u\in L^{2q}$. Moreover, there exists some $\alpha(n)>0$ such that for any $\eta\in C^{\infty}_{c}(B_{1/2})$,
[…]
Proof.
We can let $f=Cu^{q_{0}}$ in this case. Then by Lemma 5.8, we have $u\in L^{2q}$ for any $q>1$. Moreover, if we choose $\delta\leqslant\delta_{0}(n,\varepsilon)$, we know $f=Cu^{q_{0}}\in L^{2n}(B_{1/2})$ and in fact $\|f\|_{L^{2n}(B_{1/2})}\leqslant C(n,\varepsilon)\|u^{1+\varepsilon}\|_{L^{2}(B_{1})}$.
Then we can go back to the proof of Lemma 5.7. Now we know $u\in L^{2q}$ and we want to prove (5.29). Following the same steps as in the proof of Lemma 5.7, note that the only place where we use the estimate for $f$ is the estimate of $I_{2}$, see (5.21). Now using the $L^{2n}$-norm of $f$, we can compute $I_{2}$ over $B_{1/2}$ in (5.19) as follows:
$I_{2}=q\int_{B_{1/2}}(\eta\bar{\eta}_{\varepsilon})^{2}|f|F(u)^{2}$ $\leqslant q\Big(\int_{B_{1/2}}|f|^{2n}\Big)^{\frac{1}{2n}}\cdot\Big(\int_{B_{1/2}}|(\eta\bar{\eta}_{\varepsilon})F(u)|^{\frac{4n}{2n-1}}\Big)^{1-\frac{1}{2n}}.$ (5.30)
Since $2^{*}\equiv\frac{2n}{n-2}>\frac{4n}{2n-1}>2$, then by interpolation inequality and Sobolev inequality we have
[…]
where $\alpha$ is a positive constant depending on $n$ and $\theta>0$ to be chosen. This implies
[…]
Finally we can run the iteration argument to conclude the $L^{\infty}$-estimate.
Theorem 5.10.
Let $u\in C^{1}(B_{1}^{*})$ and $u\geqslant 0$ satisfying the equation (5.28) in the sense of distributions. Suppose $|u^{q_{0}}(x,y)|\leqslant\delta|y|^{-2}$ for some $\delta\leqslant\mathbf{\delta_{0}(n,q_{0})}$. Then
[…]
Proof.
Since $|u^{q_{0}}(x,y)|\leqslant\delta|y|^{-2}$, we have $\|u\|_{L^{2+2\varepsilon}(B_{1})}\leqslant C(n,\varepsilon,\delta,q_{0})$ for some $\varepsilon\in(0,q_{0}-1)$. Let $q_{*}=1+\varepsilon$ for simplicity.
We consider the similar setting as in the previous Lemma 5.9. Namely, let $v_{0}=u^{q_{*}}$ and $v_{k}=v_{k-1}^{\tau}=(u^{q_{*}})^{\tau^{k}}$ with $\tau=\frac{n}{n-2}$. We further define the radii $r_{k}=\frac{1}{4}+\frac{1}{4^{k+1}}$. Then we can define the cutoff functions $\eta_{k}\equiv 1$ in $B_{r_{k+1}}$ and vanishing outside $B_{r_{k}}$.
According to Lemma 5.9, since $v_{k}=u^{q_{*}\cdot\tau^{k}}$ and $|\nabla\eta_{k}|^{2}\leqslant C10^{k}$, we have
$\int_{B_{1/2}}\eta_{k}^{2}|\nabla v_{k}|^{2}\leqslant C\cdot(q_{*}\tau^{k})^{\alpha}\int_{B_{1/2}}(|\nabla\eta_{k}|^{2}+\eta^{2})v_{k}^{2}\leqslant C_{1}(10\tau^{\alpha})^{k}\int_{B_{r_{k}}}v_{k}^{2}.$ (5.37)
[…]
Then by Sobolev inequality and the inequality above, we have
$\Big(\int_{B_{r_{k+1}}}v_{k}^{2\tau}\Big)^{1/\tau}\leqslant\Big(\int_{B_{1/2}}|\eta_{k}v_{k}|^{2\tau}\Big)^{1/\tau}\leqslant C\int_{B_{1/2}}|\nabla(\eta_{k}v_{k})|^{2}\leqslant C\int_{B_{1/2}}|\nabla\eta_{k}|^{2}v_{k}^{2}+\eta_{k}^{2}|\nabla v_{k}|^{2}\leqslant C_{2}(10\tau^{\alpha}+1)^{k}\int_{B_{r_{k}}}v_{k}^{2},$ (5.38)
[…]
Therefore, we can now estimate $u^{q_{*}}||_{L^{2\tau^{k+1}(B_{r_{k+1}})}}$ as follows
[…]
$\leqslant\Big(C_{2}(10\tau^{\alpha}+1)^{k}\int_{B_{r_{k}}}v_{k}^{2}\Big)^{\frac{1}{2\tau^{k}}}$ (5.40)
[…]
Then by induction we can conclude that
[…]
Let $k\to\infty$, then $\tau^{k}\to\infty$ and $\sum_{j=1}^{\infty}\frac{j}{\tau^{j}}<C(n)$. Therefore,
[…]
This finishes the proof if we choose $\varepsilon=\frac{1}{2}(q_{0}-1)$. ∎
Finally we prove the curvature bound Theorem 5.6.
Proof of Theorem 5.6.
Recall that for Einstein metrics we have the following equations: there exist constants $C_{k}(n)>0$ and $\alpha(n)>0$ such that for any $k\in\mathbb{N}$
  • $\Delta|\nabla^{k}{\rm Rm}|\geqslant-C_{k}|{\rm Rm}|\,|\nabla^{k}{\rm Rm}|$,
  • $\Delta|{\rm Rm}|^{1-\alpha}\geqslant-C_{0}|{\rm Rm}|^{2-\alpha}$.
These inequalities are classical consequences of the fact that on an Einstein metric, the curvature tensor is harmonic (seen as a $\Lambda^{2}$ valued $2$-form) and a Weitzenböck formula; see [BKN89] for example.
Let $u=|{\rm Rm}|^{1-\alpha}$ and then $u$ satisfies the equation (5.28) for $q_{0}=(1-\alpha)^{-1}$. By Theorem 1.12, we have $u^{q_{0}}(x,y)=|{\rm Rm}|(x,y)\leqslant\delta|y|^{-2}$. Therefore, the proof of the bound for $|{\rm Rm}|$ is now completed by Theorem 5.10. The estimate for higher order derivatives will be standard; see for example [DS14]. ∎
5.3. Analytic orbifold coordinates
Recall that in the previous section 5.1, we construct the $C^{0}$ coordinates around curves of singularities by gluing techniques. In this section, we will improve the regularity of the coordinate charts using the curvature estimate $|\nabla^{k}{\rm Rm}|\leqslant C$ following [DS14].
First let us recall our setting: Let $(X,p)$ be a 5-dimensional noncollapsed limit space. Let $\gamma:[-2,2]\subset\mathds{R}\to X$ be a 1-Lipschitz curve with $\gamma(0)=p$ parametrized by arc length. Suppose each $\gamma(t)$ have the same tangent cone $\mathds{R}\times\mathds{R}^{4}/\Gamma$ with $\Gamma\neq id$. Then by Proposition 5.5 we know $B_{1}(p)\setminus\gamma$ is a smooth manifold with metric $g$.
Let $\mathcal{C}^{\gamma}_{j}\equiv\{x\in X:d(x,\gamma)=j^{-1}\}$ be the cylindrical hypersurface surrounding $\gamma$. The regularity of such cylinders is not good enough for higher order derivative estimates. So we need to construct a better set of cylinders. We define $\mathcal{C}_{0}\equiv(-1,1)\times\mathbb{S}^{3}\subset\mathds{R}\times\mathds{R}^{4}$ to be the standard cylinder in $\mathds{R}^{5}$.
According to Lemma 5.4, we have the following immediate corollary.
Lemma 5.11.
There exists a sequence of smooth embeddings $f_{j}:\mathcal{C}_{0}=(-1,1)\times\mathbb{S}^{3}\to X$ such that the following holds:
  1. (1)
    $d_{GH}(\mathcal{C}_{f_{j}},\mathcal{C}^{\gamma}_{j})\cap B_{1}(p))\leqslant\varepsilon_{j}$, where $\mathcal{C}_{f_{j}}\equiv f_{j}(\mathcal{C}_{0})$;
  2. (2)
    $\big|j^{2}f_{j}^{*}g-h_{0}\big|_{C^{4}_{h_{0}}}\leqslant\varepsilon_{j}$;
  3. (3)
    $\big|j^{-1}A_{\mathcal{C}_{j}}-id\big|_{C^{3}_{h_{0}}}\leqslant\varepsilon_{j}$.
Then we can construct the analytic coordinate around the curve.
Theorem 5.12.
There exists an analytic diffeomorphism $F:B^{*}\to B_{1}(p)\setminus\gamma$ such that $F^{*}g$ extends to an analytic metric on the whole $B$.
Proof.
First we define the maps $F_{j}:\mathcal{C}_{0}\times[j^{-1},1]$ as follows:
$F_{j}(z,r)=\exp_{f_{j}(z)}\Big((r-j^{-1})N(f_{j}(z))\Big).$
where $z=(t,\theta)\in\mathcal{C}_{0}=(-1,1)\times\mathbb{S}^{3}$ and $N(f_{j}(z))$ is the unit outward normal vector of $\mathcal{C}_{f_{j}}$ at $f_{j}(z)$.
We want to show that each $F_{j}$ is a diffeomorphism from $\mathcal{C}_{0}\times[j^{-1},1]$ to its image. To prove this, let $J(r)$ a Jacobi field along the geodesic $\alpha_{z}(r)\equiv F_{j}(z,r)$. Since we have $|{\rm Rm}|\leqslant C$ in $B_{1}(p)\setminus\gamma$ by Theorem 5.6, by Rauch comparison theorem there exists some constant $C>0$ and
$C^{-1}|J(j^{-1})|_{g}\leqslant|J(r)|_{g}\leqslant Cj|J(j^{-1})|_{g}.$
Hence $F_{j}$ has no critical point in $\mathcal{C}_{0}\times[j^{-1},1]$.
We can write the metric $F_{j}^{*}g=dr^{2}+r^{2}h_{j}(z,r)$. Let $h_{0}$ be the standard metric on $\mathcal{C}_{0}=(-1,1)\times\mathbb{S}^{3}$ and we will prove $h_{j}\to h_{0}$ as $j\to\infty$ in $C^{1,\alpha}$.
Given $z\in(-1,1)\times\mathbb{S}^{3}$ and a unit tangential vector $\xi$ at $z$. Consider the Jacobi field $J(r)$ along the geodesic $\alpha_{z}(r)=F_{j}(z,r)$ satisfying
  1. (1)
    $J(j^{-1})=df_{j}(\xi)$;
  2. (2)
    $J^{\prime}(j^{-1})=A_{\mathcal{C}_{j}}(J(j^{-1}))$.
This implies that $J(r)=dF_{j}|_{(z,r)}(\xi)$. Moreover, by Lemma 5.11 we have
  1. (1)
    $\big||J(j^{-1})|_{g}-j^{-1}\big|\leqslant j^{-1}\varepsilon_{j}$;
  2. (2)
    $\big||J^{\prime}(j^{-1})|_{g}-j|J(j^{-1})|_{g}\big|\leqslant 2\varepsilon_{j}$.
Let $\{e_{1}(r),...,e_{5}(r)=\alpha^{\prime}(r)\}$ be the parallel orthonormal frame along $\alpha_{z}(r)$ and write $J(r)=\sum_{k=1}^{5}J_{k}(r)e_{k}(r)$. Then by the Jacobi field equation we have
$J_{k}^{\prime\prime}(r)+\sum_{l}R_{k5l5}|_{\alpha_{z}(r)}J_{l}(r)=0,$
where $R_{k5l5}=\langle{\rm Rm}(e_{k},e_{5})e_{l},e_{5}\rangle$. Then
$\begin{split}J_{k}(r)&=J_{k}(j^{-1})+J^{\prime}_{k}(j^{-1})\cdot(r-j^{-1})-\int_{j^{-1}}^{r}\int_{j^{-1}}^{s}R_{k5l5}|_{\alpha_{z}(t)}J_{k}(t)dtds.\end{split}$
By Theorem 5.6, we have $|{\rm Rm}|\leqslant C_{1}$. This implies that for any $r\in[j^{-1},1]$ we have
$\Big||J(r)|_{g}-r\Big|\leqslant C_{2}(j^{-1}+\varepsilon_{j}r+r^{3}).$
Hence, at the metric level,
$\Big|h_{j}-h_{0}\Big|_{C^{0}_{h_{0}}}\leqslant C_{2}(j^{-1}r^{-1}+\varepsilon_{j}+r^{2}).$
By differentiating the Jacobi equation and the bound $|\nabla^{k}{\rm Rm}|\leqslant C(k)$ we obtain the high order derivative estimates
$\begin{split}r^{-1}\Big|\nabla^{g_{0}}(h_{j}-h_{0})\Big|_{C^{0}_{h_{0}}}\leqslant C_{3}(j^{-1}r^{-2}+\varepsilon_{j}r^{-1}+r);\\ r^{-2}\Big|\nabla^{g_{0}}\nabla^{g_{0}}(h_{j}-h_{0})\Big|_{C^{0}_{h_{0}}}\leqslant C_{4}(j^{-1}r^{-2}+\varepsilon_{j}r^{-1}+1).\end{split}$
For any fixed $r\ll 1$ , we let $j\to\infty$. Then the metric $F^{*}_{j}$ converges in $C^{3}$ to some map on $r,1]\times\mathcal{C}_{0}$. Then we let $r\downarrow 0$ and thus we obtain a limit map $F:(0,1]\times\mathcal{C}_{0}$ such that the metric $F^{*}g$ satisfies
$r^{-2}\Big|F^{*}g-g_{0}\Big|_{C^{0}_{g_{0}}}+r^{-1}\Big|F^{*}g-g_{0}\Big|_{C^{1}_{g_{0}}}+\Big|F^{*}g-g_{0}\Big|_{C^{2}_{g_{0}}}\leqslant C_{5}.$
This implies that $F^{*}g$ extends to a $C^{1,1}$ metric on the whole ball $B_{1}$. It is classical that if an Einstein metric is $C^{1,\alpha}$ in some coordinates, then they can be upgraded to a better set of coordinates where the metric is real-analytic. This can be proven thanks to harmonic coordinates, or coordinates in Bianchi gauge or divergence-free gauge with respect to a background flat metric. ∎
5.4. Structure for curves of singularities
We now prove Theorem 1.18 for curves of singularities.
Proof of Theorem 1.18.
Let $\gamma:I\to\mathcal{S}(X)$ be a singular curve and $\mathcal{V}(t)$ be the volume density at the point $\gamma(t)$ for $t\in I$. Since $\mathcal{V}$ is lower semi-continuous with values in a finite set, then there exists a nowhere dense subset $\mathcal{I}\subset I$ such that
  1. (1)
    $I=\mathcal{I}\cup\bigcup_{j=1}^{\infty}(s_{j},s_{j+1})$;
  2. (2)
    $\mathcal{V}$ is constant on each interval $(s_{j},s_{j+1})$.
By Corollary 4.16, for each $j$ there exists some $\Gamma_{j}$ such that the tangent cones on each segment $\gamma((s_{j},s_{j+1}))$ are all isometric to $\mathds{R}\times\mathds{R}^{4}/\Gamma_{j}$. Combining Theorem 5.6 and 5.12, we can prove (3). To prove (2), note that the metric is invariant under the nontrivial group $\Gamma_{j}$ acting by isometry and fixing a curve. It is a classical consequence that this curve must be a geodesic. ∎
5.5. Geodesics on limits of Einstein $5$-manifolds
We conclude by classifying all geodesics in the limit space $X$. We first recall the Hölder continuity estimate along geodesic in [CoN12] and also [Den25].
Theorem 5.13.
Let $\sigma:[0,l]\to X$ be a geodesic in the limit space $X$. Then there exists $\alpha,C>0$ such that given any $\delta>0$ with $\delta l<s<t<l-\delta l$ and $Y_{\sigma(s)},Y_{\sigma(t)}$ tangent cones from the same sequence of rescalings, then we have
$d_{GH}(B_{1}(Y_{\sigma(s)}),B_{1}(Y_{\sigma(t)}))<\frac{C}{\delta l}|s-t|^{\alpha}.$
As an immediate consequence, if one point of the geodesic lies in the regular set $\mathcal{R}$, then the entire interior of the geodesic lies in $\mathcal{R}$ [CoN12]. In particular the regular set $\mathcal{R}$ is convex. Using Theorem 5.13 together with our isolation Theorem 1.9, we can now classify all geodesics in $X$:
Proof of Theorem 1.21.
It suffices to consider the case where the entire $\sigma((0,l))\in\mathcal{S}$. By the stratification Theorem 1.15, there must be some $\Gamma$ and $t_{0}\in(0,l)$ such that $\sigma(t_{0})\in\mathcal{S}^{1}_{\Gamma}$ since $\mathcal{S}^{0}$ is only countable. Then by Theorem 5.13 and isolation Theorem 1.9, there exists some $\theta_{0}>0$ such that the tangent cone at $\sigma(t)$ is isometric to $\mathds{R}\times\mathds{R}^{4}/\Gamma$ for any $t\in(t_{0}-\theta_{0},t_{0}+\theta_{0})$. By connectedness this implies that any interior point has the unique tangent cone $\mathds{R}\times\mathds{R}^{4}/\Gamma$. Therefore we can use Theorem 1.16 to prove that the interior of this geodesic is removable in the sense of Theorem 1.16. ∎
Theorem 1.21 therefore provides a complete classification of the interior of all geodesics. We say that a geodesic is of $\Gamma$-type if the tangent cones at all interior points are $\mathds{R}\times\mathds{R}^{4}/\Gamma$ (we allow $\Gamma=\{id\}$). We now turn to the behavior of the end points. We prove that the tangent cones at end points can only have worse singularities. Similar results can be formulated for any curves thanks to Theorem 1.16.
Proposition 5.14.
Let $\sigma:[0,l)\to X$ be a geodesic of $\Gamma$-type in $X$. Let $C(Z)$ be a tangent cone at $\sigma(0)$. Then $Z$ has at least one singularity of the type $\mathds{R}^{4}/\Gamma^{\prime}$ with $|\Gamma^{\prime}|\geq|\Gamma|$.
Proof.
Let $C(Z)$ be the tangent cone at $\sigma(0)$ from the rescaling sequence $s_{i}$. By passing to a subsequence, we have $\sigma(s_{i})\to z$ for some point $z\in Z$. Since $\sigma$ is of $\Gamma$-type, by Bishop-Gromov monotonicity and Colding’s volume convergence theorem, the volume density at $z\in Z\subset C(Z)$ cannot be larger than that of $\mathds{R}^{4}/\Gamma$. This completes the proof. ∎
Part III Appendix: Analysis in weighted spaces
Appendix A Function spaces
For a tensor $s$, a point $x$, $\alpha>0$ and a Riemannian manifold $(M,g)$. The Hölder seminorm is defined as
$s]_{C^{\alpha}(g)}(x):=\sup_{\{y\in T_{x}M,|y|<\textup{inj}_{g}(x)\}}\Big|\frac{s(x)-s(\exp^{g}_{x}(y))}{|y|^{\alpha}}\Big|_{g}.$
For orbifolds, we consider a norm which is bounded for tensors decaying at the singular points.
Definition A.1 (Weighted Hölder norms on an orbifold).
Let $\beta\in\mathbb{R}$, $k\in\mathbb{N}$, $0<\alpha<1$ and $(M_{o},g_{o})$ an orbifold and $S_{o}$ a subset of its singularities. Then, for any tensor $s$ on $M_{o}$, we define
$\|s\|_{C^{k,\alpha}_{\beta}(g_{o})}$ $:=\sup_{M_{o}}r_{o}^{-\beta}\Big(\sum_{i=0}^{k}r_{o}^{i}|\nabla_{g_{o}}^{i}s|_{g_{o}}+r_{o}^{k+\alpha}[\nabla_{g_{o}}^{k}s]_{C^{\alpha}(g_{o})}\Big).$
For ALE manifolds, we will consider a norm which is bounded for tensors decaying at infinity.
Definition A.2 (Weighted Hölder norms on an ALE orbifold).
Let $\beta\in\mathbb{R}$, $k\in\mathbb{N}$, $0<\alpha<1$ and $(N,g_{b})$ be an ALE manifold and $S_{b}$ a subset of its singularities. Then, for all tensor $s$ on $N$, we define
$\|s\|_{C^{k,\alpha}_{\beta}(g_{b})}:=\sup_{N}r_{b}^{\beta}\Big(\sum_{i=0}^{k}r_{b}^{i}|\nabla_{g_{b}}^{i}s|_{g_{b}}+r_{b}^{k+\alpha}[\nabla_{g_{b}}^{k}s]_{C^{\alpha}({g_{b}})}\Big).$
On $M$, using the partition of unity $1=\chi_{M_{o}^{t}}+\sum_{j}\chi_{N_{j}^{t}}$ from Definition 2.11, we define a global norm.
Definition A.3 (Weighted Hölder norm on a naïve desingularization).
Let $\beta\in\mathbb{R}$, $k\in\mathbb{N}$ and $0<\alpha<1$. We define for $s\in TM^{\otimes l_{+}}\otimes T^{*}M^{\otimes l_{-}}$ a tensor $(l_{+},l_{-})\in\mathbb{N}^{2}$, with $l:=l_{+}-l_{-}$ the associated conformal weight,
$\|s\|_{C^{k,\alpha}_{\beta}(g^{D})}:=\|\chi_{M_{o}^{t}}s\|_{C^{k,\alpha}_{\beta}(g_{o})}+\sum_{j}T_{j}^{\frac{l}{2}}\|\chi_{N_{j}^{t}}s\|_{C^{k,\alpha}_{\beta}(g_{b_{j}})}.$
With the notations of Definition 2.1, denote for each singular point $k$, $A_{k}(t,\varepsilon_{0}):=(\Phi_{k})_{*}A_{{e}}(\varepsilon_{0}^{-1}\sqrt{t_{k}},\varepsilon_{0})$ and $B_{k}(\varepsilon_{0}):=(\Phi_{k})_{*}B_{{e}}(0,\varepsilon_{0})$, as well as cut-off functions $\chi_{A_{k}(t,\varepsilon_{0})}$ and $\chi_{B_{k}(\varepsilon_{0})}$ respectively supported in $A_{k}(t,\varepsilon)$ and $B_{k}(\varepsilon_{0})$, and equal to $1$ on $A_{k}(t,2\varepsilon_{0})$ and $B_{k}(\varepsilon_{0}/2)$.
Definition A.4 ($C^{k,\alpha}_{\beta,*}$-norm on $2$-tensors).
Let $h$ be a $2$-tensor on $(M,g^{D})$, $(M_{o},g_{o})$ or $(N,g_{b})$. We define
$\|h\|_{C^{k,\alpha}_{\beta,*}}:=\inf_{h_{*},H_{k}}\|h_{*}\|_{C^{k,\alpha}_{\beta}}+\sum_{k}|H_{k}|_{{e}},$
where the infimum is taken on the $(h_{*},(H_{k})_{k})$ satisfying $h=h_{*}+\sum_{k}\chi_{A_{k}(t,\varepsilon_{0})}H_{k}$ for $(M,g^{D})$ or $h=h_{*}+\sum_{k}\chi_{B_{k}(\varepsilon_{0})}H_{k}$ for $(M_{o},g_{o})$ or $(N,g_{b})$, where each $H_{k}$ is a constant traceless symmetric $2$-tensors on $\mathbb{R}^{4}/\penalty 50\Gamma_{k}$.
Definition A.5 ($rC^{k,\alpha}_{\beta,*}$-norm of a vector fields).
Let $X$ a vector field on $(M,g^{D})$ (respectively $(M_{o},g_{o})$ or $(N,g_{b})$). We define its $rC^{k,\alpha}_{\beta,*}$-norm, where $r$ is the function $r_{D}$ (respectively $r_{o}$ or $r_{b}$) by
$\|X\|_{rC^{k,\alpha}_{\beta,*}}:=\inf_{X_{*},X_{k}}\|X_{*}\|_{rC^{k,\alpha}_{\beta}}+\sum_{k}\|X_{k}\|_{rC^{0}_{0}(e)},$
where the infimum is taken among the couples $(X_{*},X_{k})$ satisfying $X=X_{*}+\sum_{k}\chi_{\mathcal{A}_{k}(t,\varepsilon)}X_{k}$ (respectively $X=X_{*}+\sum_{k}\chi_{B_{o}(\varepsilon)}X_{k}$ or $X=X_{*}+\chi_{B_{b}(\varepsilon)}X_{k}$).
A last function space is adapted to the asymptotics of $2$-tensors and vector fields at ALE ends.
Definition A.6 ($C^{2,\alpha}_{\beta,**}$ norm on a tree of ALE orbifold).
[…]
$\|h\|_{C^{2,\alpha}_{\beta,**}}:=\|r^{2}H^{2}\|_{L^{\infty}(e)}+\|r^{4}H^{4}\|_{L^{\infty}(e)}+\big\|(1+r_{b})^{4}(h-\chi(H^{2}+H^{4}))\big\|_{C^{2,\alpha}_{\beta,*}}.$
Remark A.7.
If for a metric $g$ such that $\|g-g^{B}\|_{C^{2,\alpha}_{\varepsilon}}\leqslant 1$, one has $\delta_{g}h=0$ and $h\in C^{2,\alpha}_{\beta,**}$, then there is no term $H^{2}$ in $r^{-2}$ in the decomposition of $h$ since these terms are not divergence-free by [Ozu20, Lemma 2.50]. There is no term in $r^{-3}$ because of the action of $\Gamma\neq\{\mathrm{id}\}$.
Definition A.8 ($r_{b}C^{3,\alpha}_{\beta,**}$-norm on a tree of ALE orbifold).
Let $(N,g^{B})$ a naïve gluing of a tree of Ricci-flat ALE orbifolds. Let $X$ be a vector field on $N$, and assume that $X=Y^{3}+\mathcal{O}(r_{b}^{-3-\beta})$ for $\beta>0$ and $Y^{3}$ a homogeneous element of the kernel of $\delta_{e}\delta_{e}^{*}$ with $|Y^{3}|\sim r_{e}^{-3}$. We define its $r_{b}C^{3,\alpha}_{\beta,**}$-norm by
$\|X\|_{r_{b}C^{3,\alpha}_{\beta,**}}:=\sup\|r_{e}^{3}Y^{3}\|_{L^{\infty}(e)}+\big\|(1+r_{b})^{4}(X-\chi(\varepsilon r_{b})Y^{3})\big\|_{r_{b}C^{3,\alpha}_{\beta,*}}.$
Remark A.9.
The norms $C^{2,\alpha}_{\beta,**}$ and $rC^{3,\alpha}_{\beta,**}$ only differ from the norms $C^{2,\alpha}_{\beta,*}$ and $rC^{3,\alpha}_{\beta,*}$ in the way they weigh a neighborhood of infinity. They are the same in the rest of the manifold.
The reason why we consider these two function spaces is because we have the following mapping properties for the linearization $\bar{P}$ of $\mathbf{\Phi}$. The following statement corrects [Ozu22b, Lemma 5.5].
Lemma A.10.
[Corrected statement of [Ozu22b, Lemma 5.5]] Let $(N,g^{B})$ be a tree of ALE orbifolds. Then, there exists $C>0$ such that we have, for any $h\perp\tilde{\mathbf{O}}(g^{B})$,
$\|h\|_{C^{2,\alpha}_{\beta,**}}\leqslant C\big\|(1+r_{b})^{4}\bar{P}_{g^{B}}h\big\|_{r_{D}^{-2}C^{\alpha}_{\beta}}.$ (A.1)
We also have a similar statement for $\delta\delta^{*}$.
Lemma A.11 ([Ozu22b, Lemma 5.6]).
Let $(N,g^{B})$ be a tree of ALE orbifold. Then, there exists $C>0$ such that for any vector field $X$ on $N$, we have
$\|X\|_{r_{B}C^{3,\alpha}_{\beta,**}}\leqslant C\big\|(1+r_{B})^{4}\delta_{g^{B}}\delta_{g^{B}}^{*}X\big\|_{r_{B}^{-1}C^{1,\alpha}_{\beta}}.$ (A.2)
Remark A.12.
In [Ozu22b], the $H^{2}$ term was omitted in the definition of the $C^{2,\alpha}_{\beta,**}$ norm, see [Ozu22b, Definition 5.3]. We a priori need to include these $H^{2}$ terms even though they will instantly be killed by our gauge condition, see Remark A.7. In particular in their current forms, [Ozu22b, Lemma 5.5 and Corollary 5.7] should either include a $H^{2}$ term, or impose that $h$ is in gauge.
This does not affect the results of [Ozu22b]: [Ozu22b, Corollary 5.7] is only applied to gluings of hyperkähler ALE metrics where there is no obstruction, and the $H^{2}$ term do not appear since they are not divergence-free. Concretely, in [Ozu22b], Lemma 5.5 and Corollary 5.7 are mentioned in Lemma 5.15, but only applied to the obstructions $\mathbf{o}_{j}$ which are in gauge, and they are mentioned in Propositions 6.15 and 6.16 where they are applied to hyperkähler metrics where there are no obstructions and the tensors are in gauge.
Appendix B Duality for vector fields
We will have to carefully understand the duality between elements of the kernel of $\delta_{e}\delta_{e}^{*}$, where we identify $1$-forms and vector fields through $X\sim g(X,\cdot)$. From [Ozu22b, Section 3.1], which essentially follows [CT94, Section 2], the homogeneous $1$-forms in $\ker\delta_{e}\delta_{e}^{*}$ are of the form
  1. (1)
    $r^{a^{\pm}_{j}}\psi$ with $a^{\pm}_{j}:=\pm(1+j)$, $j\in\mathbb{N}^{*}$, where $\psi$ is an eigenvector the Hodge Laplacian,
  2. (2)
    $r^{b^{\pm}_{j}}d_{\mathbb{S}^{3}/\penalty 50\Gamma}\phi\;+\;b^{\pm}_{j}r^{b^{\pm}_{j}-1}\phi dr$, $b_{j}^{\pm}=-1\pm(1+j)$ or
  3. (3)
    $2r^{c^{\pm}_{j}+2}d_{\mathbb{S}^{3}/\penalty 50\Gamma}\phi\;+\;c^{\mp}_{j}r^{c^{\pm}_{j}+1}\phi dr$, with $c_{j}^{\pm}=-1\pm(1+j)$, $j\in\mathbb{N}$ and $\phi$ an eigenfunction of the Laplacian.
Consider for each of the above tensors, $(\phi_{i}^{j})_{i}$ and $(\psi_{i}^{j})_{i}$ orthonormal bases of eigentensors of the eigenspaces above. Denote $X_{a_{j}^{\pm},i}$, $X_{b_{j}^{\pm},i}$ and $X_{c_{j}^{\pm},i}$ the associated vector fields in $\ker\delta_{e}\delta_{e}^{*}$.
Among these tensors, those whose norm is in $r$ correspond to: $a_{1}^{+}$, $b_{2}^{+}$ and $c^{+}_{0}$. Those with a norm in $r^{-3}$ correspond to and those whose norm grows like $r^{-3}$ are: $a_{1}^{-}$, $b_{0}^{-}$ and $c^{-}_{2}$.
An application of [Ozu22b, Lemma 3.4] yields the following result.
Lemma B.1.
Let $(N,g^{B})$ be a tree of ALE orbifold. Then, for any linear combination $X_{1}:=\alpha X_{a_{1}^{+}}+\beta X_{b_{2}^{+}}+\gamma X_{c_{0}^{+}}$ for $\alpha,\beta,\gamma\in\mathbb{R}$, there exists a unique vector field $\mathcal{X}_{1}$ on $N$ such that: for a cut-off $\chi$ supported in a neighborhood of infinity of $N$ where $g^{B}$ has ALE coordinates,
$\left\{\begin{aligned} &\delta_{g^{B}}\delta_{g^{B}}^{*}\mathcal{X}_{1}=0,\\ &\mathcal{X}_{1}-\chi X_{1}\in r_{B}C^{3,\alpha}_{\delta},\text{ for }0<\delta<4\end{aligned}\right$. (B.1)
We additionally have:
$\|\mathcal{X}_{1}-\chi X_{1}\|_{r_{B}C^{3,\alpha}_{\delta}}\leqslant C\|r^{-1}X_{1}\|_{L^{\infty}(g_{e})}$ (B.2)
There is a duality between tensors in $\ker\delta_{e}\delta^{*}_{e}$ whose norms are growing in $r$ and in $r^{-3}$. Namely, when integrating by parts $\int_{M}\delta\delta^{*}X(Y)-\delta\delta^{*}Y(X)$ on an ALE manifold $M$, one gets a boundary term
$B(X,Y):=\lim_{s\to\infty}\int_{\{r=s\}}\left(\delta^{*}X(Y,\partial_{r})-\delta^{*}Y(X,\partial_{r})\right).$
It turns out that if $Y=Y^{3}+o(r^{-3})$ and $X=X_{1}+o(r)$ for $|Y^{3}|\propto r^{-3}$ and $|X_{1}|\propto r$, then $B(X,Y)=B(X_{1},Y^{3})$ defines an important coupling. Explicit computations show that for three positive constants $C_{a}$, $C_{b}$ and $C_{c}$, one has: for any $\alpha_{i},\beta_{i},\gamma_{i},\alpha^{\prime}_{k},\beta^{\prime}_{k},\gamma^{\prime}_{k}\in\mathbb{R}$,
$\sum_{i,k}B\left(\alpha_{i}X_{a_{1}^{+},i}+\beta_{i}X_{b_{2}^{+},i}+\gamma_{i}X_{c_{0}^{+},i}\,,\,\,\alpha^{\prime}_{k}X_{a_{1}^{-},k}+\beta^{\prime}_{k}X_{c_{2}^{-},k}+\gamma^{\prime}_{k}X_{b_{0}^{-},k}\right)=\sum_{i}C_{a}\alpha_{i}\alpha^{\prime}_{i}+C_{b}\beta_{i}\beta^{\prime}_{i}+C_{c}\gamma_{i}\gamma^{\prime}_{i}.$ (B.3)
In other words, the bilinear form $B$ essentially diagonalizes in the above basis. This is a tool to precisely estimate how large the terms in $r^{-3}$ are in the expansion of vector fields in $r_{B}C^{3,\alpha}_{\delta,**}$.
Proposition B.2.
Let $Y=Y^{3}+o(r^{-3})\in r_{B}C^{3,\alpha}_{\delta,**}$ for $\delta>0$, with a decomposition $Y^{3}=\alpha^{\prime}_{k}X_{a_{1}^{-},k}+\beta^{\prime}_{k}X_{c_{2}^{-},k}+\gamma^{\prime}_{k}X_{b_{0}^{-},k}$, consider $X_{1}=\alpha_{i}X_{a_{1}^{+},i}+\beta_{i}X_{b_{2}^{+},i}+\gamma_{i}X_{c_{0}^{+},i}$ and the associated $\mathcal{X}_{1}$ from Lemma B.1. Then, one has
$\int_{M}\delta\delta^{*}Y(\mathcal{X}_{1})=-\sum_{i}C_{a}\alpha_{i}\alpha^{\prime}_{i}+C_{b}\beta_{i}\beta^{\prime}_{i}+C_{c}\gamma_{i}\gamma^{\prime}_{i}.$
Proof.
This is a direct consequence of the following integration by parts: using (B.3) in the last line,
$-\int_{M}\delta\delta^{*}Y(\mathcal{X}_{1})$ $=\int_{M}\delta\delta^{*}\mathcal{X}_{1}(Y)-\delta\delta^{*}Y(\mathcal{X}_{1})=B(Y^{3},X_{1})$
$=\sum_{i}C_{a}\alpha_{i}\alpha^{\prime}_{i}+C_{b}\beta_{i}\beta^{\prime}_{i}+C_{c}\gamma_{i}\gamma^{\prime}_{i}.$